### v1

# On the number of modes of Gaussian kernel density estimators

Abstract
We consider the Gaussian kernel density estimator with bandwidth $\beta^{-\frac{1}{2}}$ of $n$ iid Gaussian samples. Using the Kac-Rice formula and an Edgeworth expansion, we prove that the expected number of modes on the real line scales as $\Theta(\sqrt{\beta\log\beta})$ as $\beta,n\to\infty$ provided $n^{c}\lesssim\beta\lesssim n^{2-c}$ for some constant $c>0$. An impetus behind this investigation is to determine the number of clusters to which Transformers are drawn in a metastable state.
Keywords. Kernel density estimator, Kac-Rice formula, Edgeworth expansion, self-attention, mean-shift.
ams classification.62G07, 60G60, 60F05, 68T07.
1 Introduction
1.1 Setup and main result
For $\beta>0$ and $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$, the Gaussian kernel density estimator (KDE) with bandwidth $h=\beta^{-1/2}$ is defined as
$\widehat{P}_{n}(t):=\frac{1}{n}\sum_{i=1}^{n}\mathsf{K}_{h}\ast\delta_{X_{i}}(t)=\frac{\sqrt{\beta}}{n\sqrt{2\pi}}\sum_{i=1}^{n}e^{-\frac{\beta}{2}(t-X_{i})^{2}},\hskip 28.45274ptt\in\mathbb{R}.$ (1.1)
Here, “Gaussian” refers to the choice of kernel $\mathsf{K}_{h}$.
In this paper we are interested in determining the expected number of modes (local maxima) of $\widehat{P}_{n}$ over $\mathbb{R}$. While this is a classical question, addressed in even more general settings than (1.1)—such as non-Gaussian kernels, compactly supported samples, and higher dimensions [MMF92, Mam95, KM97]—a definite answer has not been given in the literature. Indeed, the best-known results fall into one of two settings: either considering samples drawn from a compactly supported density (instead of $N(0,1)$ as done here), or counting the modes within a fixed compact interval. In the special case of the Gaussian KDE (1.1), one has in the latter setting for instance
Theorem 1.1 ([Mam95, Thm. 1]).
Let $\widehat{P}_{n}$ be the Gaussian KDE defined in (1.1), with bandwidth $h>0$, of $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$. Asymptotically as $n\to\infty$, the expected number of modes of $\widehat{P}_{n}$ in a fixed interval $[a,b]$ is $1+o(1)$ if $1\ll\beta\ll n^{2/3}$, and $\widetilde{\Theta}\mathopen{}\mathclose{{}\left(\sqrt{\beta}}\right)$ if $n^{2/3}\lesssim\beta\ll n^{2}/\log^{6}n$.
In [MMF92, Mam95, KM97], the authors additionally conduct more refined casework on the bandwidth to provide more precise estimates, such as pinpointing the leading constants. In fact, [MMF92] does count modes in $\mathbb{R}$, but the underlying distribution of the samples $X_{i}$ is supported on a closed interval (thus excluding $N(0,1)$), so there are no modes outside the interval anyway.
Our main result provides the answer to the case of counting modes of (1.1) over $\mathbb{R}$, and reads as follows. In particular, it generalizes Theorem 1.1.
Theorem 1.2.
Let $\widehat{P}_{n}$ be the Gaussian KDE defined in (1.1), with bandwidth $\beta^{-1/2}$, of $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$. Suppose $n^{c}\lesssim\beta\lesssim n^{2-c}$ for arbitrarily small $c>0$, and 4.2, then asymptotically as $n,\beta\to\infty$,
  1. 1.
    In expectation over $X_{i}$, the number of modes of $\widehat{P}_{n}$ is $\Theta\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$.
  2. 2.
    Almost all modes lie in two intervals of length $\Theta\mathopen{}\mathclose{{}\left(\sqrt{\log\beta}}\right)$—namely, the expected number of modes $t\in\mathbb{R}$, such that $t^{2}\not\in\mathopen{}\mathclose{{}\left[2\log n-3\log\beta,2\log n-\log\beta}\right]$, is $o\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$.
Figure 1: A realization of the random function (1.1) for $n=10^{4}$, with $\beta=100$ (left) and $\beta=300$ (right).
4.2 is related to the decay of the tails of the modulus of continuity of $\widehat{P}_{n}^{\prime\prime}(\cdot)$, and is needed to apply the Kac-Rice formula (Theorem 2.1), which states that the expected number of modes of $\widehat{P}_{n}$ is some conditional expectation with respect to the joint law of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{\prime\prime}(t))$. We postpone a further discussion to Section 4.1.
Remark 1.3.
  • To better appreciate the range of values for $\beta$ in this theorem as well as subsequent ones, we use minimax theory as a benchmark; see, e.g., [Tsy09]. The reparametrization $h=\sqrt{\beta}$ is motivated by the connection to the Transformer model described in Section 1.2. Using an optimal bias-variance tradeoff [Tsy09, Chapter 1], we see that the optimal scaling of the bandwidth parameter $h$ depends on the smoothness of the underlying density of interest: if the underlying density has $s$ bounded (fractional) derivatives, then the optimal choice of $h$ is given by $h\asymp n^{-\frac{1}{2s+1}}$. This gives $\beta\asymp n^{\frac{2}{2s+1}}$. For $s\in(0,\infty)$, we get $\beta\in[n^{c},n^{2-c}]$ for some $c>0$. In particular, the transition of the number of modes from 1 to $\sqrt{\beta}$ in Theorem 1.1 is achieved for $\beta\approx n^{2/3}$, which is the optimal choice for Lipschitz densities. The message of our main Theorem 1.2 below is that this scaling in $\sqrt{\beta}$ is the prevailing one for the whole range $\beta\in[n^{c},n^{2-c}]$ if one does not restrict counting modes in a bounded interval $[a,b]$.
  • Point 2. in Theorem 1.2 shows that most of the modes are at distance at least $C\log n$ from the origin provided $\beta>n^{\frac{2-C}{3}}$ for $C>0$ small. This corresponds to a choice of a bandwidth adapted to smoothness $s<1$. This result is in agreement with and completes the picture drawn by Theorem 1.1.
Remark 1.4.
Through refined computations, one can determine the modes in the regime $1\ll\beta\ll n^{2}$ and also pinpoint the leading constant. For the sake of clarity, we stick to the regime where
$2\log n-\log\beta\asymp\log\beta\asymp\log n,$
and comment on how to do expand the regime in appropriate places.
Remark 1.5.
We further motivate Point 2. in Theorem 1.2 by considering a qualitative picture of the distribution of the modes displayed in Figure 4.
  • Near the origin, we find most of the samples $X_{i}$ and they are densely packed in the shape of a Gaussian. The corresponding Gaussian summands in 1.1 cancel to create one mode, as shown already in Theorem 1.1.
  • In the two intervals of length $\Theta\mathopen{}\mathclose{{}\left(\sqrt{\log\beta}}\right)$, the samples $X_{i}$ are separated enough that the corresponding Gaussian summands do not cancel, but rather form $\Omega\mathopen{}\mathclose{{}\left(\sqrt{\log\beta}}\right)$ isolated bumps, as discussed in more generality in [DG85, Section 9.3].
  • Further away at the tails, the phenomena of isolated bumps occur, but there are so few samples $X_{i}$ that the number of modes created is a negligible fraction.
We revisit this discussion and Figure 4 in Remark 3.3.
To prove Theorem 1.2, we truncate the real line to the interval
$T:=\mathopen{}\mathclose{{}\left[-\sqrt{2\log n-\log\beta-\omega(\beta)},\sqrt{2\log n-\log\beta-\omega(\beta)}}\right]$ (1.2)
where $\omega$ is a fixed, slow growing function such that
$1\ll\omega(\beta)\ll\log\log\beta,$
and so $T$ is well-defined for large enough $\beta$. Motivated by Theorem 1.2, we also define the interval
$T^{\prime}:=\mathopen{}\mathclose{{}\left[-\sqrt{2\log n-3\log\beta},\sqrt{2\log n-3\log\beta}}\right]$ (1.3)
if $\beta\leqslant n^{2/3}$ and define $T^{\prime}=\varnothing$ if $\beta>n^{2/3}$.
We see how Theorem 1.2 implies Theorem 1.1 with $h=\beta^{-1/2}$. From the former, we see that almost all the modes lie in $T\setminus T^{\prime}$. If $h\gg n^{-1/3}$ so that $\sqrt{2\log n-3\log\beta}\gg 1$, then $T\setminus T^{\prime}$ is disjoint from $[a,b]$, so there are few modes in $[a,b]$; if $h\ll n^{-1/3}$ so $\sqrt{2\log n-3\log\beta}\ll 1$, the fixed interval $[a,b]$ contains a near constant fraction of length of $T\setminus T^{\prime}$ and thus a near constant fraction of the modes.
Figure 2: (Left) Plot of the average number of modes as a function of $\beta$ for $n=10^{3}$ (top) and $n=10^{4}$ (bottom). (Right) Log-log plot for $n=10^{3}$ (top) and $n=10^{4}$ (bottom); the predicted linear regression line (red) corroborates a power-law of the form average # of $\text{average \# of modes}\approx 0.179\cdot\beta^{0.504}$, in line with Theorem 1.2.
1.2 Motivation
The question of estimating the number of modes as a function of the bandwidth has a plethora of applications in statistical inference and multimodality tests—see [MMF92, Mam95, KM97] and the references therein. Another application which has stimulated some of the recent progress on the topic is data clustering. The latter can be achieved nonparametrically using a KDE, whose modes, and hence clusters, can be detected using the mean-shift algorithm [FH75, Che95, CM02, CP00, CPW03, CP07, RL14, CP15], which can essentially be seen as iterative local averaging. The main idea in mean-shift clustering is to perform a mean-shift iteration starting from each data point and then define each mode as a cluster, with all points converging to the same mode grouped into the same cluster. The analysis of this algorithm has led to upper bounds on the number of modes of (1.1) [CPW03].
Figure 3: Metastability of self-attention dynamics at temperature $\beta=81$ initialized with $n$ iid uniform points on the circle, with $n=200$ (top) and $n=1000$ (bottom). The number of clusters appears of the correct order $\sim\sqrt{\beta}$. (Code available at github.com/borjanG/2023-transformers-rotf.)
We were instead brought to this problem from another perspective, motivated by the study of self-attention dynamics [SABP22, GLPR23, GLPR24, GRRB24]—a toy model for Transformers, the deep neural network architecture that has driven the success of large language models [VSP+17]. These dynamics form a mean-field interacting particle system
$\frac{\,\mathrm{d}}{\,\mathrm{d}\tau}x_{i}(\tau)=\sum_{j=1}^{n}\frac{e^{\beta\langle x_{i}(\tau),x_{j}(\tau)\rangle}}{\displaystyle\sum_{k=1}^{n}e^{\beta\langle x_{i}(\tau),x_{j}(\tau)\rangle}}\mathsf{P}_{x_{i}(\tau)}^{\perp}(x_{j}(\tau)),$
evolving on the unit sphere $\mathbb{S}^{d-1}$ because of $\mathsf{P}_{x}^{\perp}:=I_{d}-xx^{\top}$. Here, $\tau\geqslant 0$ plays the role of layers, the $n$ particles $x_{i}(\tau)$ represent tokens evolving through a dynamical system. This system is characterized by a temperature parameter $\beta\geqslant 0$ that governs the space localization of particle interations. One sees that all particles move in time by following the field $\nabla\log(\mathsf{K}_{\beta^{-1/2}}\ast\mu_{\tau})$; here, $\mu_{\tau}$ is the empirical measure of the particles $x_{1}(\tau),\ldots,x_{n}(\tau)$ at time $\tau$.
It is shown that for almost every initial configuration $x_{1}(0),\ldots,x_{n}(0)$, and for $\beta\geqslant 0$ in dimension $\geqslant 3$, or $\beta\leqslant 1\vee\beta\gtrsim n^{2}$ in dimension $2$, all particles converge to a single cluster in infinite time [GLPR23]. Rather than converging quickly, the authors in [GKPR24] prove that the dynamics instead manifest metastability: particles quickly approach a few clusters, remain in the vicinity of these clusters for a very long period, and eventually coalesce into a single cluster in infinite time. Concurrently, and using different methods, the authors in [BPA24] show a similar result: starting from a perturbation of the uniform distribution, beyond a certain time, the empirical measure of the $n$ particles approaches an empirical measure of $O(\sqrt{\beta})$-equidistributed points on the circle in the mean-field limit, and stays near it for long time. This is done by a study of the linearized system and leveraging nonlinear stability results from [Gre00].
Our interest lies in counting the number of clusters during the first metastable phase in dimension $d=2$. At time $\tau=0$, particles are initialized at $n$ iid points from the uniform distribution on the circle. In turn, the stationary points of $\mathsf{K}_{\beta^{-1/2}}\ast\mu_{0}$ partition the circle into intervals, with points clustering within their respective interval. This highlights the importance of counting the number of stationary points.
Here, we focus on a simplified setting by working on the real line instead of the circle (or higher-dimensional spheres), but we believe the analysis could be extended to these cases pending technical adaptations. Notwithstanding, Theorem 1.2 reflects what is seen in simulations (Figure 3).
1.3 Sketch of the proof
The spirit of the proof of results such as Theorem 1.1 and others presented in [MMF92, Mam95, KM97] is similar to ours—one applies the Kac-Rice formula (Theorem 2.1) to a Gaussian approximation of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{{}^{\prime\prime}}(t))$ and argues its validity. However, the main limitation of these works is that modes are counted in a fixed and finite interval $[a,b]$ (and $[0,1]^{d}$ in the higher dimensional cases). Extending these techniques to the whole real line demands for different, significantly stronger, approximation results using Edgeworth expansions.
We sketch the key ideas that allow us to count modes over $\mathbb{R}$. We use the Kac-Rice formula to compute the expected number of modes of $\widehat{P}_{n}$ in the symmetric interval $T$ and $T^{\prime}$ defined in 1.2 and 1.3. All asymptotics are as $n,\beta\to\infty$.
Proposition 1.6.
If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for arbitrarily small $c>0$, then under 4.2,
  1. 1.
    In expectation over $X_{i}$, the number of modes of $\widehat{P}_{n}$ in $T$ is $\Theta\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$.
  2. 2.
    In expectation over $X_{i}$, the number of modes of $\widehat{P}_{n}$ in $T^{\prime}$ is $O\mathopen{}\mathclose{{}\left(e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}}\right)$.
The Kac-Rice computation appears tractable only when the joint distribution of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{\prime\prime}(t))$ is Gaussian, which it is not. To overcome this obstacle, we apply the Kac-Rice formula over a Gaussian approximation of the joint distribution in Section 2. For the specific underlying density and KDE in 1.1, we are able to justify in Section 3 the approximation for all $t$ in the growing interval $T$ instead of a fixed interval. This is why Theorem 1.1 only counts modes in a fixed interval.
To show the validity of the Gaussian approximation, we use the Edgeworth expansion of the joint distribution of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{\prime\prime}(t))$ around the Gaussian distribution with matching first two moments. We bound the error due to the third order term of the expansion directly, and deal with the higher order terms by appealing to the error bounds of densities in the Edgeworth approximation similar to [BR10, Theorem 19.2]. We note that [KM97] employ the same theorem to justify the Gaussian process approximation over a fixed interval.
In doing so, we will see that the normal approximation is invalid outside of $T$ (see Remark 3.3), but crucially $T$ is sufficiently large to cover almost all modes, as observed empirically in Remark 1.5 and Figure 4 and given below.
Proposition 1.7.
If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for arbitrarily small $c>0$, then the expectation over $X_{i}$ of the number of modes of $\widehat{P}_{n}$ that lie outside of $T$ is $O\mathopen{}\mathclose{{}\left(\sqrt{\beta\exp(\omega(\beta))}}\right)$.
We prove this in Section 4.2 with an argument from scale-space theory: we bound the number of modes outside $T$ by the number of samples $X_{i}$ outside $T$, which we then bound naively. This is precisely the argument used by [CPW03, Theorem 2] to show Gaussian mixtures over $\mathbb{R}$ with $n$ components must have at most $n$ modes. This argument crucially relies on the kernel density estimator being Gaussian (see Remark 4.5). Now, Theorem 1.2 follows from Propositions 1.6 and 1.7 upon recalling that $1\ll\omega(\beta)\ll\log\log\beta$.
1.4 Notation
We adopt standard notation from asymptotic analysis: we write $f(x)\ll g(x)$ or $f(x)=o(g(x))$ if $f(x)/g(x)\to 0$ as $x\to\infty$; $f(x)\lesssim g(x)$ or $f(x)=O(g(x))$ if there exists a finite, positive constant $C$ such that $f(x)\leqslant Cg(x)$; and we write $f(x)\asymp g(x)$ or $f(x)=\Theta(g(x))$ if $f(x)\lesssim g(x)$ and $g(x)\lesssim f(x)$.
2 Kac-Rice for the normal approximation
2.1 The Kac-Rice formula
We say that $\Psi:\mathbb{R}\to\mathbb{R}$ has an upcrossing of level $u$ at $t\in\mathbb{R}$ if $\Psi(t)=u$ and $\Psi^{\prime}(t)>0$. The Kac-Rice formula allows us to compute the expected number of up-crossings when $F$ is a random field (i.e., a stochastic process).
Theorem 2.1 (Kac-Rice, [AW09, pp. 62], [AT09, Section 11.1]).
Consider a random $\Psi:\mathbb{R}\to\mathbb{R}$, some fixed $u\in\mathbb{R}$ and a compact $T\subset\mathbb{R}$. Suppose
  1. 1.
    $\Psi$ is a.s. in $\mathscr{C}^{1}(\mathbb{R})$, and $\Psi,\Psi^{\prime}$ both have finite variance over $T$;
  2. 2.
    The law of $\Psi(t)$ admits a density $p_{t}^{[1]}(x)$ which is continuous for $t\in T$ and $x$ in a neighborhood of $u$;
  3. 3.
    The joint law of $(\Psi(t),\Psi^{\prime}(t))$ admits a density $p_{t}(x,y)$ which is continuous for $t\in T$, $x$ in a neighborhood of $u$, and $y\in\mathbb{R}$;
  4. 4.
    $\mathbb{P}(\upomega(\eta)>\varepsilon)=O(\eta)$ as $\eta\to 0^{+}$ for any $\varepsilon>0$, where $\upomega(\cdot)$ denotes the modulus of continuity111defined, for $f:\mathbb{R}\to\mathbb{R}$, as $\upomega(\eta)=\sup_{t,s\colon|t-s|\leq\eta}|f(t)-f(s)|$. of $\Psi^{\prime}(\cdot)$.
Define the number of up-crossings in $T$ of $\Psi$ at level $u\in\mathbb{R}$ as
$U_{u}(\Psi,T):=\mathopen{}\mathclose{{}\left|\{t\in T:\Psi(t)=u,\Psi^{\prime}(t)>0\}}\right|.$
Then, with expectation taken over the randomness of $\Psi$,
$\mathbb{E}U_{u}(\Psi,T)=\int_{T}\int_{0}^{\infty}yp_{t}(u,y)\,\mathrm{d}y\,\mathrm{d}t.$ (2.1)
The Kac-Rice formula extends to any dimension $d\geqslant 1$, and also on manifolds other than $\mathbb{R}^{d}$—see [AT09, Section 11.1]. It is the classical tool for computing the expected number of critical points of random fields, with many recent applications including spin glasses [AAČ13, FMM21] and landscapes of loss functions arising in machine learning [MAB20]. While the method applies to general densities, the conditional expectation appears infeasible to compute or estimate beyond the Gaussian case.
For the KDE $\widehat{P}_{n}$ defined in 1.1, define the random function $F_{n}:\mathbb{R}\to\mathbb{R}$ by
$F_{n}(t)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathopen{}\mathclose{{}\left(t-X_{i}}\right)e^{-\frac{\beta}{2}\mathopen{}\mathclose{{}\left(t-X_{i}}\right)^{2}}=-\sqrt{\frac{2\pi n}{\beta^{3}}}\widehat{P}_{n}^{\prime}(t).$ (2.2)
Then $t\in\mathbb{R}$ is an upcrossing of $F_{n}$ at level 00 if and only if $F_{n}(t)=0$ and $F^{\prime}_{n}(t)>0$. This is equivalent to $\widehat{P}^{\prime}_{n}(t)=0$ and $\widehat{P}_{n}^{\prime\prime}(t)<0$, i.e. $t$ is a mode of $\widehat{P}_{n}$. Thus, the number of modes of $\widehat{P}_{n}$ in $T$ is given by $U_{0}(F_{n},T)$. For $T,T^{\prime}$ defined in (1.2)–(1.3), Propositions 1.6 and 1.7 yield
$\mathbb{E}U_{0}\mathopen{}\mathclose{{}\left(F_{n},T}\right)$ $\asymp\sqrt{\beta\log\beta},$ (2.3)
$\mathbb{E}U_{0}\mathopen{}\mathclose{{}\left(F_{n},T^{\prime}}\right)$ $\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta},$
$\mathbb{E}U_{0}\mathopen{}\mathclose{{}\left(F_{n},\mathbb{R}\setminus T}\right)$ $\lesssim e^{\frac{\omega(\beta)}{2}}\sqrt{\beta}.$
2.2 Computing the Gaussian approximation
Without loss of generality, fix $t\in T$ with $t\geqslant 0$. We can rewrite $F_{n}(t)$ from 2.2 and compute its derivative: for independent copies $(G_{i},G_{i}^{\prime})$ of
$\begin{bmatrix}G(t)\\ G^{\prime}(t)\end{bmatrix}=e^{-\frac{\beta}{2}(t-X)^{2}}\begin{bmatrix}t-X\\ 1-\beta(t-X)^{2}\end{bmatrix},$ (2.4)
where $X\sim N(0,1)$, we have
$\begin{bmatrix}F_{n}(t)\\ F_{n}^{\prime}(t)\end{bmatrix}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\begin{bmatrix}G_{i}(t)\\ G_{i}^{\prime}(t)\end{bmatrix}\sim p_{t}.$
We prove that $p_{t}$ is a well-defined density in Proposition 4.1, and defer the following computation to Sections A.1 and A.1.
Fact 2.2.
The mean and covariance matrix of the random vector $(F_{n}(t),F_{n}^{\prime}(t))$ are given respectively by
$\mu_{t}$ $:=\sqrt{n}\begin{bmatrix}\mathbb{E}G(t)\\ \mathbb{E}G^{\prime}(t)\end{bmatrix}\asymp n^{\frac{1}{2}}\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}t\\ -t^{2}\end{bmatrix}$ (2.5)
$\Sigma_{t}$ $:=\begin{bmatrix}\operatorname{Var}G(t)&\operatorname{Cov}(G(t),G^{\prime}(t))\\ \operatorname{Cov}(G(t),G^{\prime}(t))&\operatorname{Var}G^{\prime}(t)\end{bmatrix}\asymp\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}1&-t\\ -t&\beta\end{bmatrix}.$
We proceed to centering and rescaling the density $p_{t}$. Let $Y_{i}(t)$, $i\in[n]$, be independent copies of
$Y(t)=\Sigma_{t}^{-\frac{1}{2}}\begin{bmatrix}G(t)-\mathbb{E}G(t)\\ G^{\prime}(t)-\mathbb{E}G^{\prime}(t)\end{bmatrix}$ (2.6)
Let $q_{t}$ denote the density of $\sqrt{n}\sum_{i=1}^{n}Y_{i}(t)$. By construction $q_{t}$ has mean 00 and covariance $I_{2}$. Moreover, by the change-of-variables formula, it holds
$p_{t}(x,y)=\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}q_{t}\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(x,y)-\mu_{t}]}\right).$ (2.7)
Now, let $\varphi:\mathbb{R}^{2}\to\mathbb{R}$ be the density of $N(0,I_{2})$, i.e.,
$\varphi(x):=\frac{1}{\sqrt{2\pi}}e^{-\frac{\|x\|^{2}}{2}}.$
We aim to approximate the Kac-Rice integral 2.1 as follows:
$\int_{T}\int_{0}^{\infty}yp_{t}(0,y)\,\mathrm{d}y\approx\int_{T}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t.$ (2.8)
The validity of this approximation is deferred to Section 3. In the remainder of this section, we solely focus on computing the right hand side integral.
Lemma 2.3.
There exists some $A_{t}\asymp\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}$ such that
$\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)$ $\asymp\exp\mathopen{}\mathclose{{}\left(-A_{t}-\Theta\mathopen{}\mathclose{{}\left(\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}}\right)y^{2}}\right),$ (2.9)
$\int_{0}^{\infty}y\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y$ $\asymp\beta^{-\frac{1}{2}}e^{-\frac{t^{2}}{2}}e^{-A_{t}}.$
Proof of Lemma 2.3.
Since $\det\Sigma_{t}\asymp\beta^{-2}e^{-t^{2}}$, we have
$\Omega:=\Sigma_{t}^{-1}\asymp\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}\begin{bmatrix}\beta&t\\ t&1\end{bmatrix}.$
Now as $t\in T$ and so $t^{2}\ll\beta$, we find
$\mathopen{}\mathclose{{}\left\|\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right\|^{2}$ $=\mathopen{}\mathclose{{}\left\langle(-\mu_{t,1},y-\mu_{t,2}),\Sigma_{t}^{-1}(-\mu_{t,1},y-\mu_{t,2})}\right\rangle$
$\asymp\Omega_{11}\mu_{t,1}^{2}-2\Omega_{12}\mu_{t,1}(y-\mu_{t,2})+\Omega_{22}(y-\mu_{t,2})^{2}$
$\asymp\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}\mu_{t,1}^{2}\mathopen{}\mathclose{{}\left[\beta-2t\mathopen{}\mathclose{{}\left(\frac{y}{\mu_{t,1}}+t}\right)+\mathopen{}\mathclose{{}\left(\frac{y}{\mu_{t,1}}+t}\right)^{2}}\right]$
$\asymp\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}\mu_{t,1}^{2}(\beta-t^{2})+\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}y^{2}$
$\asymp\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}+\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}y^{2}.$
This shows the first statement in 2.9. For the second, we have
$\int_{0}^{\infty}y\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y$ $\asymp e^{-A_{t}}\int_{0}^{\infty}ye^{-\Theta\mathopen{}\mathclose{{}\left(\beta^{1/2}e^{t^{2}/{2}}}\right)y^{2}}\,\mathrm{d}y$
$\asymp\beta^{-\frac{1}{2}}e^{-\frac{t^{2}}{2}}e^{-A_{t}}$
by Gaussian integral computations (see A.1). ∎
2.3 The Kac-Rice integral over $\varphi$
Figure 4: $n=10^{5}$ is fixed throughout. (Left) Empirical distribution of the modes of $\widehat{P}_{n}$ over $T$ for $\beta=100$ (top) and $\beta=300$ (bottom). (Right) The function $t\mapsto\sqrt{\beta}\exp(-A_{t})$ for $\beta=100$ (top) and $\beta=300$ (bottom), which, due to the Kac-Rice formula, is an approximation for the distribution of the number of modes of $\widehat{P}_{n}$ in $T$. Shaded in grey is the interval $T$. (Code available at github.com/KimiSun18/2024-gauss-kde-attention.)
We compute 2.1 under the approximation 2.8. By 2.9 and 2.5, we have that
$\int_{S}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t\asymp\sqrt{\beta}\int_{S}e^{-A_{t}}\,\mathrm{d}t$ (2.10)
for any measurable $S\subset\mathbb{R}$. Assuming validity of the Gaussian approximation (see Section 3), it follows from the Kac-Rice formula that the density of modes at $t\in\mathbb{R}$ is proportional to $\sqrt{\beta}e^{-A_{t}}$. We plot this density in Figure 4 with the same choice of $n$ and $\beta$ as in the empirical distribution. We see that they match on the highlighted interval $T$, but not outside of $T$ where the Gaussian approximation—see Remark 3.3.
We compute 2.10 explicitly for $S=T$ and $S=T^{\prime}$.
Lemma 2.4.
If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for some $c>0$, then
$\int_{T}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$ $\asymp\sqrt{\beta\log\beta},$
$\int_{T^{\prime}}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$ $\lesssim\sqrt{\beta}.$
Proof of Lemma 2.4.
Recall $A_{t}$ from Lemma 2.3. By 2.10, it suffices to show that
$\int_{T}e^{-A_{t}}\,\mathrm{d}t\asymp\sqrt{\log\beta}\quad\text{and}\quad\int_{T^{\prime}}e^{-A_{t}}\,\mathrm{d}t\lesssim 1.$ (2.11)
As $A_{t}>0$, the integral is at most the length of $T$, which is $O(\sqrt{\log\beta})$ by 1.2. Let $t_{s}:=\sqrt{2\log n-s\log\beta}$. As the integrand is positive, for constants $C,C^{\prime}>0$
$\int_{T}e^{-A_{t}}\,\mathrm{d}t$ $\geqslant\int_{t_{2}}^{t_{5/2}}\exp\mathopen{}\mathclose{{}\left(-C\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}}\right)\,\mathrm{d}t$
$\geqslant\mathopen{}\mathclose{{}\left(t_{5/2}-t_{2}}\right)\exp\mathopen{}\mathclose{{}\left(-C\beta^{-\frac{3}{2}}nt_{5/2}^{2}e^{-\frac{t_{2}^{2}}{2}}}\right)$
$\gtrsim\sqrt{\log\beta}\exp\mathopen{}\mathclose{{}\left(-C^{\prime}\beta^{-\frac{1}{2}}\log n}\right)$
$\gtrsim\sqrt{\log\beta}$
as $n,\beta\to\infty$ with $\log n\asymp\log\beta$. Now if $t\in T^{\prime}=[-t_{3},t_{3}]$, we have
$e^{-\frac{t^{2}}{2}}\geqslant e^{-\frac{t_{3}^{2}}{2}}=\beta^{\frac{3}{2}}n^{-1}.$
Hence
$\int_{T^{\prime}}e^{-A_{t}}\,\mathrm{d}t\lesssim\int_{0}^{t_{3}}\exp\mathopen{}\mathclose{{}\left(-C\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}}\right)\,\mathrm{d}t\leqslant\int_{0}^{t_{3}}e^{-Ct^{2}}\,\mathrm{d}t\lesssim 1.\qed$
3 Leveraging the Edgeworth expansion
In this section, we show the approximation of $p_{t}$ by $\varphi$ is valid in $T$ by showing
$\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{}\left|q_{t}-\varphi}\right|\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t\ll\sqrt{\beta\log\beta}.$ (3.1)
One natural idea is to use some asymptotic series to expand $p_{t}$ around $\varphi$, e.g. the Edgeworth expansion, to argue that $|p_{t}-\varphi|\ll\varphi$ in the sense of the integral over $y$. We discuss the two major obstacles we have to overcome in order to implement this approach.
  • Firstly, all known results on validity of asymptotic series such as Edgeworth expansions treat densities $q_{t}$ and $\varphi$ that are independent of $n$. As $\beta$ grows in $n$, we will need to re-derive these results and carefully track the dependence on $\beta$. This will give extra constraints on $(t,\beta,n)$ for the validity of such an asymptotic series. Fortunately, this will be satisfied precisely when $t\in T$.
  • Secondly, even without the $n$-dependence of $\beta$, expanding $q_{t}$ to the order $\varphi$ plus error in for example [BR10, Theorem 19.2] gives error roughly
[…]
    but then the integral over $y$ in 3.1 fails to converge. Therefore, we will need to go to the next term $\psi$ in the Edgeworth series. We control its integral over $y$ and $t$ in 3.1 manually in Section 3.1. Then, we control the higher order terms in Section 3.2 using the approach motivated above, so our error now converges upon integrating over $y$, i.e. roughly
    $\mathopen{}\mathclose{{}\left|p_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right|(\mathbf{x})\lesssim\frac{1}{1+\|\mathbf{x}\|^{3}}\hskip 28.45274pt\text{ for all }\mathbf{x}\in\mathbb{R}^{2}.$
3.1 Bounding the third order error
Recall $Y$ from 2.6. For a multi-index $\alpha\in\mathbb{Z}_{\geqslant 0}^{2}$, let $\kappa^{\alpha}_{t}$ be the cumulant of $Y$ indexed by $\alpha$, which depends on $t$. Let $H^{\alpha}_{t}$ denote the standard Hermite polynomials of with index $\alpha$. The next term in the Edgeworth series is $n^{-1/2}\psi$ with
$\psi(\mathbf{x}):=\frac{\varphi(\mathbf{x})}{6}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}(\mathbf{x})\quad\text{where}\quad H^{\alpha}:=(-1)^{|\alpha|}\frac{\partial^{\alpha}\varphi}{\varphi}$ (3.2)
If $|\alpha|=3$, $\kappa_{t}^{\alpha}$ are the third moments of $Y$, which we bound in Section A.1.
Fact 3.1.
Let $\eta_{3}$ be the largest third cumulant of $Y$. Then
$\eta_{3}:=\max\mathopen{}\mathclose{{}\left\{\kappa_{t}^{(k,3-k)}:k\in\{0,1,2,3\}}\right\}\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}.$
We bound $H^{(k,3-k)}\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-1/2}[(0,y)-\mu_{t}]}\right)$ by a polynomial in $\tilde{y}=\Theta(\beta^{1/2}e^{{t^{2}}/{4}}y)$. Now, by a similar method as Lemma 2.3, we obtain the following bound. It says that when we integrate the Edgeworth series $p_{t}=\varphi+n^{-1/2}\psi+\dots$ over $y$ and $t$, the contribution $\varphi$ dominates $n^{-1/2}\psi$, hinting at the validity of the approximation.
Lemma 3.2.
Recall $T$ from 1.2 and $A_{t}$ from Lemma 2.3. Then
$\int_{T}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left(n\det\Sigma_{t}}\right)^{-\frac{1}{2}}\psi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}.$
Proof of Lemma 3.2.
By the proof of Lemma 2.3, 3.2, and 3.1
$\int_{T}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left(n\det\Sigma_{t}}\right)^{-\frac{1}{2}}\psi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$
$=\frac{1}{6}\sum_{k=0}^{3}\int_{T}\mathopen{}\mathclose{{}\left(n\det\Sigma_{t}}\right)^{-\frac{1}{2}}\kappa_{t}^{(k,3-k)}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left[\varphi H^{(k,3-k)}}\right]\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$
$\asymp\int_{T}\mathopen{}\mathclose{{}\left(n\det\Sigma_{t}}\right)^{-\frac{1}{2}}\eta_{3}e^{-A_{t}}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left(\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}y}\right)^{O(1)}e^{-\Theta\mathopen{}\mathclose{{}\left(\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}}\right)y^{2}}\,\mathrm{d}y\,\mathrm{d}t$
[…]
$\lesssim n^{-\frac{1}{2}}\beta^{\frac{3}{4}}\sup_{t\in T}\mathopen{}\mathclose{{}\left(e^{\frac{t^{2}}{4}}}\right)\int_{T}e^{-A_{t}}\,\mathrm{d}t$
$\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}$
where the last step follows 2.11 and the definition 1.2 of $T$. ∎
Remark 3.3.
One can see at this is actually an asymptotic equality by checking the Gaussian integrals in the proof above are of their typical order (i.e. no cancellation of leading terms). Hence, the decay is only a factor of $e^{-\omega(\beta)/4}$. For $t\not\in T$, even $t=\sqrt{2\log n-0.99\log\beta}$, the last inequality in Lemma 3.2 fails and we get a bound of polynomially larger that $\sqrt{\beta}$. Then, as the third order error is asymptotically larger than the contribution of the Gaussian approximation, so the normal approximation is no longer valid. This can be seen by comparing the plots in Figure 4.
3.2 Bounding higher order errors
We follow the classical proof about the validity of the Edgeworth expansion as an asymptotic series to show bound the higher order pointwise error of density function as follows.
Lemma 3.4.
Assume that $p_{t}$ is bounded almost everywhere for any fixed $t\in T$. If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for some $c>0$, then for any $t\in T$ and $\mathbf{x}\in\mathbb{R}^{2}$, we have that
$\mathopen{}\mathclose{{}\left(1+\|\mathbf{x}\|^{3}}\right)\mathopen{}\mathclose{{}\left|q_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right|(\mathbf{x})\lesssim e^{-\frac{\omega(\beta)}{4}}.$ (3.3)
We defer the discussion of the assumption to 4.2 and defer the proof to Section A.3. Here, we comment on the differences with [BR10, Theorem 19.2] for $s=3$, which we follow in spirit but modify to allow $n$ dependence in $q_{t}$ in the form of $\beta$. There, it is shown that the left hand side is $o(n^{-1/2})$ provided third moments $\eta_{3}=O(1)$. With the bound 3.1 on $\eta_{3}$ in our case which is not constant, the error from the asymptotic series decays not in powers of $n^{-1/2}$ but powers of $O(n^{-1/2}e^{{t^{2}}/{4}}\beta^{1/4})=O(\exp(-\omega(\beta)/2))$.
Corollary 3.5.
If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for $c>0$, then asymptotically in $n,\beta\to\infty$
$\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{}\left|q_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right|\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}.$
Proof of Corollary 3.5.
By Lemma 3.4 and computations in the proof of Lemma 2.3
$\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{}\left|q_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right|\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$
$\lesssim e^{-\frac{\omega(\beta)}{4}}\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{}\left(1+\mathopen{}\mathclose{{}\left\|\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right\|^{3}}\right)^{-1}\,\mathrm{d}y\,\mathrm{d}t$
$\lesssim e^{-\frac{\omega(\beta)}{4}}\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{}\left(1+\Theta\mathopen{}\mathclose{{}\left(\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}+\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}y^{2}}\right)^{\frac{3}{2}}}\right)^{-1}\,\mathrm{d}y\,\mathrm{d}t$
$\lesssim e^{-\frac{\omega(\beta)}{4}}\int_{T}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}\mathopen{}\mathclose{{}\left(\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}}\right)^{-2}\int_{0}^{\infty}\frac{\tilde{y}}{1+\tilde{y}^{3}}\,\mathrm{d}\tilde{y}$
$\asymp e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}$
for $\tilde{y}\asymp\beta^{1/4}e^{{t^{2}}/{4}}y$, where we note that $T$ has length $\Theta(\sqrt{\log\beta})$ by 1.2 and
$\int_{0}^{\infty}\frac{\tilde{y}}{1+\tilde{y}^{3}}\,\mathrm{d}\tilde{y}=\frac{2\pi}{3\sqrt{3}}=O(1).\qed$
4 Proof of Theorem 1.2
We prove Propositions 1.6 and 1.7 by checking 2.3, thereby proving Theorem 1.2.
4.1 Proof of Proposition 1.6
To prove Proposition 1.6 we seek to apply Theorem 2.1 to $\mathbb{E}U_{0}(F_{n},T)$. This in turn requires checking all the assumptions of Theorem 2.1. We have
Proposition 4.1.
Let $\beta,n$ be as in Theorem 1.2, and fix $t\in T$. Let $\upmu_{t}$ denote the law of $(F_{n}(t),F_{n}^{\prime}(t))$ defined in (2.2). Then $\upmu_{t}$ admits a density $p_{t}\in\mathscr{C}^{0}(\mathbb{R}^{2})$ satisfying $p_{t}({\bf{x}})\to 0$ as $\|\bf{x}\|\to\infty$. Moreover, conditions 1, 2 in Theorem 2.1 also hold for $\Psi(t)=F_{n}(t)$.
We defer the proof to Section A.4. We work under assumption
Assumption 4.2.
Consider the setting of Proposition 4.1. We assume that condition 4 of Theorem 2.1 holds for $\Psi(t)=F_{n}(t)$.
To check conditions on moduli of continuity such as 4 in Theorem 2.1 in the Gaussian setting, one usually resorts to using results such as the Borell-TIS inequality [AT09, Theorem 2.1.1]. Checking the validity of this assumption in the present, non-Gaussian, setting does not appear straightforward.
With Proposition 4.1 and 4.2, we deduce
Lemma 4.3.
With the notation as in Theorem 2.1,
$\mathbb{E}U_{0}\mathopen{}\mathclose{{}\left(F_{n},T}\right)=\int_{T}\int_{0}^{\infty}yp_{t}(0,y)\,\mathrm{d}y\,\mathrm{d}t.$ (4.1)
Proof of Proposition 1.6.
Combining Lemmas 2.4, 3.2 and 3.5 gives
$\mathbb{E}U_{0}\mathopen{}\mathclose{{}\left(F_{n},T}\right)$ $=\int_{T}\int_{0}^{\infty}yp_{t}(0,y)\,\mathrm{d}y\,\mathrm{d}t$
$=\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}yq_{t}\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$
$=\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$
$\quad+\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(n\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\psi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$
$\quad+\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{}\left[q_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right]\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$
$\asymp\sqrt{\beta\log\beta}$
for $t\in T$, as the first summand is $\Theta\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$ while the last two are $o\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$. Replacing $T$ by $T^{\prime}=[-\sqrt{2\log n-3\log\beta},\sqrt{2\log n-3\log\beta}]$, the first summand is $o\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$ by Lemma 2.4. By positivity of the integrand, we bound the last two integrals over $T^{\prime}$ by those over $T$, which are themselves $o\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$. Now, all three summands are $o\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$, as desired in Proposition 1.6. ∎
Figure 5: An estimate of the density $p_{t}=p_{t}(x,y)$ of $(F_{n}(t),F_{n}^{\prime}(t))$ for $t=0,1,2,3$ (clockwise from top left), where $\beta=81$ and $n=6500$, so that $\sqrt{2\log n-\log\beta}\approx 3$. (Code available at github.com/KimiSun18/2024-gauss-kde-attention.)
4.2 Proof of Proposition 1.7
In this section, we prove Proposition 1.7.
Lemma 4.4.
For any $a>0$ and $X_{1},\dots,X_{n}\in\mathbb{R}$, the number of modes of $\widehat{P}_{n}$ in $(a,\infty)$ is at most $|I|$ where $I=\{i\in[n]:X_{i}\geqslant a\}$.
Proof of Lemma 4.4.
Note that $\widehat{P}_{n}(t)=\sum_{i=1}^{n}g_{i}(t)$ where for $i\in[n]$ we define
$g_{i}(t):=\sqrt{\frac{\beta}{2\pi n^{2}}}\mathsf{K}_{\beta^{-1/2}}\mathopen{}\mathclose{{}\left(t-X_{i}}\right).$ (4.2)
For $i\not\in I$, $g_{i}$ is monotonically decreasing on $[X_{i},\infty)\supset(a,\infty)$, so $\sum_{i\not\in I}g_{i}(t)$ has no modes in $(a,\infty)$. To this Gaussian mixture, we add in $g_{i}(t)$ for $i\in I$ one-by-one. By [CPW03, Theorem 2], each time the number of modes in $(a,\infty)$ increases by at most one. In $|I|$-many steps, there are at most $|I|$ such modes. ∎
Remark 4.5.
This argument crucially relies on the KDE being Gaussian. As discussed in [CPW03], the Gaussian kernel is the only kernel where for any fixed samples the number of modes of the KDE is non-increasing in the bandwidth $h$. For other kernels, we do suspect the analog of Lemma 4.4 to hold, but a different argument is needed. In particular, [MMF92, Mam95, KM97] avoids this problem by counting modes on compact sets.
Proof of Proposition 1.7.
By Lemma 4.4, symmetry of $T$ in 1.2 around $t=0$, linearity of expectations, and the tail bound $\mathbb{P}(|X|\geqslant a)\leqslant 2e^{-a^{2}/2}$ for $X\sim N(0,1)$
$\mathbb{E}U_{0}\mathopen{}\mathclose{{}\left(F_{n},\mathbb{R}\setminus T}\right)$ $\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left|\{i:X_{i}\not\in T\}}\right|$ (4.3)
$=n\mathbb{P}\mathopen{}\mathclose{{}\left(X\not\in T}\right)$
$\leqslant 2n\exp\mathopen{}\mathclose{{}\left(-\frac{2\log n-\log\beta-\omega(\beta)}{2}}\right)$
$=2\sqrt{\beta\exp(\omega(\beta))}$
$\ll\sqrt{\beta\log\beta}$
by the definition of $\omega(\beta)$, proving Proposition 1.7. ∎
This concludes the proof of Theorem 1.2.
5 Concluding remarks
We showed that the expected number of modes of a Gaussian KDE with bandwidth $\beta^{-\frac{1}{2}}$ of $n\geqslant 1$ samples drawn iid from $N(0,1)$ is of order $\Theta(\sqrt{\beta\log\beta})$ for $n^{c}\lesssim\beta\lesssim n^{2-c}$, where $c>0$ is arbitrarily small. We also provide a precise picture of where the modes are located.
The question in the higher-dimensional case, as well as on the unit sphere $\mathbb{S}^{d-1}$ with uniformly distributed samples, remains open.
Acknowledgments
The authors would like to thank Enno Mammen for useful discussion and sharing important references. We also thank Dan Mikulincer for discussions on Gaussian approximation using Edgeworth expansions, Valeria Banica for comments on the method of stationary phase, and Alexander Zimin for providing Figure 2.
Funding. P.R. was supported by NSF grants DMS-2022448, CCF-2106377, and a gift from Apple. Y.S. was supported by the MIT UROP and MISTI France Programs.
Appendix A Additional proofs
A.1 Proof of 2.2
In this section we compute the first two moments of $(G,G^{\prime})$ to prove 2.2. Note that if $n^{c}\lesssim\beta\lesssim n^{2-c}$ for some $c>0$, and for $t\in T$, then we have $\exp\Theta(t^{2}/\beta)\to 1$. This implies that exponentials in the moments are asymptotically $e^{-{t^{2}}/{2}}$. We frequently use the following fact about Gaussian integrals both in exact and asymptotic forms.
Fact A.1.
Let $\Gamma$ denote the Gamma function. For any $\alpha>0$ and integer $n\geqslant 0$,
$\int_{0}^{\infty}z^{n}e^{-\alpha u^{2}}\,\mathrm{d}u=\frac{1}{2}\Gamma\mathopen{}\mathclose{{}\left(\frac{n+1}{2}}\right)\alpha^{-\frac{n+1}{2}}.$
We first compute $\mu_{t}$. Completing the square gives
$\frac{\beta}{2}z^{2}+\frac{1}{2}(z-t)^{2}=\frac{\beta+1}{2}u^{2}+\frac{\beta t^{2}}{2(\beta+1)}\quad\text{where}\quad u=z-\frac{t}{\beta+1}.$
Hence, using A.1 we compute
$\mathbb{E}G(t)=\int_{-\infty}^{\infty}ze^{-\frac{\beta}{2}z^{2}}\,\mathrm{d}\gamma_{t,1}(z)$ $=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}ze^{-\frac{\beta}{2}z^{2}-\frac{1}{2}(z-t)^{2}}\,\mathrm{d}z$
$=\frac{e^{-\frac{\beta}{2(\beta+1)}t^{2}}}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathopen{}\mathclose{{}\left(u+\frac{t}{\beta+1}}\right)e^{-\frac{\beta+1}{2}u^{2}}\,\mathrm{d}u$
$=\frac{e^{-\frac{\beta}{2(\beta+1)}t^{2}}}{\sqrt{2\pi}}\mathopen{}\mathclose{{}\left(\frac{t}{\beta+1}}\right)\frac{\sqrt{\pi}}{\mathopen{}\mathclose{{}\left(\frac{\beta+1}{2}}\right)^{\frac{1}{2}}}$
$=\frac{e^{-\frac{\beta}{2(\beta+1)}t^{2}}t}{(\beta+1)^{\frac{3}{2}}},$
as well as
$\mathbb{E}G^{\prime}(t)$ $=\int_{-\infty}^{\infty}(1-\beta z^{2})z^{-\frac{\beta}{2}z^{2}}\,\mathrm{d}\gamma_{t,1}(z)$
$=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathopen{}\mathclose{{}\left(1-\beta z^{2}}\right)e^{-\frac{\beta}{2}z^{2}-\frac{1}{2}(z-t)^{2}}\,\mathrm{d}z$
$=\frac{e^{-\frac{\beta}{2(\beta+1)}t^{2}}}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathopen{}\mathclose{{}\left[1-\beta\mathopen{}\mathclose{{}\left(u+\frac{t}{\beta+1}}\right)^{2}}\right]e^{-\frac{\beta+1}{2}u^{2}}\,\mathrm{d}u$
$=\frac{e^{-\frac{\beta}{2(\beta+1)}t^{2}}}{\sqrt{2\pi}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(1-\frac{\beta t^{2}}{(\beta+1)^{2}}}\right)\frac{\sqrt{\pi}}{\mathopen{}\mathclose{{}\left(\frac{\beta+1}{2}}\right)^{\frac{1}{2}}}-\frac{\beta\sqrt{\pi}}{2\mathopen{}\mathclose{{}\left(\frac{\beta+1}{2}}\right)^{\frac{3}{2}}}}\right]$
$=\frac{e^{-\frac{\beta}{2(\beta+1)}t^{2}}}{(\beta+1)^{\frac{5}{2}}}\mathopen{}\mathclose{{}\left((\beta+1)^{2}-\beta t^{2}-\beta(1+\beta)}\right)$
$=\frac{e^{-\frac{\beta}{2(\beta+1)}t^{2}}}{(\beta+1)^{\frac{5}{2}}}\mathopen{}\mathclose{{}\left(1+\beta-\beta t^{2}}\right).$
From these computations, and the remark after A.1, we readily obtain the asymptotics of $\mu_{t}$ as in 2.2 upon multiplying by $\sqrt{n}$.
We now compute $\Sigma_{t}$. Completing the square gives
$z^{2}+\frac{1}{2}(z-t)^{2}=\frac{2\beta+1}{2}u^{2}+\frac{\beta t^{2}}{2\beta+1}\quad\text{where}\quad u=z-\frac{t}{2\beta+1}.$
Hence using A.1 we compute
$\mathbb{E}G^{2}(t)$ $=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}z^{2}e^{-\beta z^{2}-\frac{1}{2}(z-t)^{2}}\,\mathrm{d}z$
$=\frac{e^{-\frac{\beta}{(2\beta+1)}t^{2}}}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathopen{}\mathclose{{}\left(u+\frac{t}{2\beta+1}}\right)^{2}e^{-\frac{1+2\beta}{2}u^{2}}\,\mathrm{d}u$
$=\frac{e^{-\frac{\beta}{(2\beta+1)}t^{2}}}{\sqrt{2\pi}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{t}{2\beta+1}}\right)^{2}\frac{\sqrt{\pi}}{\mathopen{}\mathclose{{}\left(\frac{2\beta+1}{2}}\right)^{\frac{1}{2}}}+\frac{\sqrt{\pi}}{2\mathopen{}\mathclose{{}\left(\frac{2\beta+1}{2}}\right)^{\frac{3}{2}}}}\right]$
$=\frac{e^{-\frac{\beta}{(2\beta+1)}t^{2}}}{(2\beta+1)^{\frac{5}{2}}}\mathopen{}\mathclose{{}\left(t^{2}+2\beta+1}\right),$
as well as
$\mathbb{E}\mathopen{}\mathclose{{}\left[G(t)G^{\prime}(t)}\right]=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}z(1-\beta z^{2})e^{-\beta z^{2}-\frac{1}{2}(z-t)^{2}}\,\mathrm{d}z$
$=\frac{e^{-\frac{\beta}{(2\beta+1)}t^{2}}}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathopen{}\mathclose{{}\left(u+\frac{t}{2\beta+1}-\beta\mathopen{}\mathclose{{}\left(u+\frac{t}{2\beta+1}}\right)^{3}}\right)e^{-\frac{1+2\beta}{2}u^{2}}\,\mathrm{d}u$
$=\frac{e^{-\frac{\beta}{2\beta+1}t^{2}}}{\sqrt{2\pi}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{t}{2\beta+1}-\beta\mathopen{}\mathclose{{}\left(\frac{t}{2\beta+1}}\right)^{3}}\right)\frac{\sqrt{\pi}}{\mathopen{}\mathclose{{}\left(\frac{2\beta+1}{2}}\right)^{\frac{1}{2}}}-\mathopen{}\mathclose{{}\left(\frac{3t\beta}{2\beta+1}}\right)\frac{\sqrt{\pi}}{2\mathopen{}\mathclose{{}\left(\frac{2\beta+1}{2}}\right)^{\frac{3}{2}}}}\right]$
$=\frac{e^{-\frac{\beta}{2\beta+1}t^{2}}}{(2\beta+1)^{\frac{7}{2}}}\mathopen{}\mathclose{{}\left[t(2\beta+1)^{2}-\beta t^{3}-3t\beta(2\beta+1)}\right]$
$=\frac{e^{-\frac{\beta}{2\beta+1}t^{2}}}{(2\beta+1)^{\frac{7}{2}}}\mathopen{}\mathclose{{}\left(-2\beta^{2}t+\beta t-\beta t^{3}+t}\right),$
and, finally,
$\mathbb{E}G^{\prime 2}(t)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}(1-\beta z^{2})^{2}e^{-\beta z^{2}-\frac{1}{2}(z-t)^{2}}\,\mathrm{d}z$
$=\frac{e^{-\frac{\beta}{2\beta+1}t^{2}}}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathopen{}\mathclose{{}\left(1-\beta\mathopen{}\mathclose{{}\left(u+\frac{t}{2\beta+1}}\right)^{2}}\right)^{2}e^{-\frac{1+2\beta}{2}u^{2}}\,\mathrm{d}u$
$=\frac{e^{-\frac{\beta}{2\beta+1}t^{2}}}{\sqrt{2\pi}}\Bigg{[}\mathopen{}\mathclose{{}\left(1-\frac{\beta t^{2}}{(2\beta+1)^{2}}}\right)^{2}\frac{\sqrt{\pi}}{\mathopen{}\mathclose{{}\left(\frac{2\beta+1}{2}}\right)^{\frac{1}{2}}}$
$\hskip 78.24507pt+\mathopen{}\mathclose{{}\left(\frac{6\beta^{2}t^{2}}{(2\beta+1)^{2}}-2\beta}\right)\frac{\sqrt{\pi}}{2\mathopen{}\mathclose{{}\left(\frac{2\beta+1}{2}}\right)^{\frac{3}{2}}}+\beta^{2}\cdot\frac{3\sqrt{\pi}}{4\mathopen{}\mathclose{{}\left(\frac{2\beta+1}{2}}\right)^{\frac{5}{2}}}\Bigg{]}$
$=\frac{e^{-\frac{\beta}{2\beta+1}t^{2}}}{(2\beta+1)^{\frac{9}{2}}}\Big{[}((2\beta+1)^{2}-\beta t^{2})^{2}+(2\beta+1)6\beta^{2}t^{2}-2\beta(2\beta+1)^{3}+3\beta^{2}(2\beta+1)^{2}\Big{]}$
$=\frac{e^{-\frac{\beta}{2\beta+1}t^{2}}}{(2\beta+1)^{\frac{9}{2}}}\mathopen{}\mathclose{{}\left(12\beta^{4}+4\beta^{3}(t^{2}+5)+\beta^{2}(t^{4}-2t^{2}+15)-2\beta(t^{2}-3)+1}\right).$
We check that entries of $\Sigma_{t}$ are asymptotically the corresponding second moments. Indeed, surpressing the dependence on $t$, we have that
  • $\mathbb{E}G^{2}\asymp\beta^{-\frac{5}{2}}(t^{2}+\beta)e^{\frac{t^{2}}{2}}\gg\beta^{-3}t^{2}E^{2}_{t}\asymp(\mathbb{E}G)^{2},$
  • $\mathbb{E}GG^{\prime}\asymp-\beta^{-\frac{7}{2}}e^{\frac{t^{2}}{2}}(\beta^{2}t+\beta t^{3})\gg\beta^{-4}E^{2}_{t}t(\beta-\beta t^{2})\asymp(\mathbb{E}G)(\mathbb{E}G^{\prime}),$
  • $\mathbb{E}G^{\prime 2}\asymp\beta^{-\frac{9}{2}}e^{\frac{t^{2}}{2}}(\beta^{4}+\beta^{2}t^{4})\gg\beta^{-5}E^{2}_{t}\beta^{2}(1+t^{4})\asymp(\mathbb{E}G^{\prime})^{2},$
where we bound $e^{\frac{t^{2}}{2}}\leqslant 1$. From these computations, and the remark after A.1, we readily obtain the asymptotics of $\Sigma_{t}$ as indicated in 2.2.∎
A.2 Proof of 3.1
In this section, we prove 3.1 on third moments of $Y=\Sigma_{t}^{-1/2}\mathopen{}\mathclose{{}\left(G-\mathbb{E}G,G^{\prime}-\mathbb{E}G^{\prime}}\right)$. To upper bound, we do not need to track the leading coefficients to ensure that they do not vanish when we combine applications of A.1. Recalling $\Sigma_{t}^{-1}$ from the proof of Lemma 2.3, we upper bound asymptotically via Hölder’s inequality:
$\eta_{3}$ $=\max_{k}\mathopen{}\mathclose{{}\left|Y^{(k,3-k)}}\right|$
$\leqslant\mathbb{E}\|Y\|^{3}$
$\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left\|\mathopen{}\mathclose{{}\left(G-\mathbb{E}G,G^{\prime}-\mathbb{E}G^{\prime}}\right)^{\intercal}\Sigma_{t}^{-1}\mathopen{}\mathclose{{}\left(G-\mathbb{E}G,G^{\prime}-\mathbb{E}G^{\prime}}\right)}\right\|^{\frac{3}{2}}$
$\lesssim\beta^{\frac{3}{4}}e^{\frac{3t^{2}}{4}}\mathbb{E}\mathopen{}\mathclose{{}\left|\beta(G-\mathbb{E}G)^{2}+2t(G-\mathbb{E}G)(G^{\prime}-\mathbb{E}G^{\prime})+(G^{\prime}-\mathbb{E}G^{\prime})^{2}}\right|^{\frac{3}{2}}$
$\lesssim\beta^{\frac{3}{4}}e^{\frac{3t^{2}}{4}}\mathopen{}\mathclose{{}\left(\beta^{\frac{3}{2}}\mathbb{E}|G|^{3}+\mathbb{E}|G^{\prime}|^{3}}\right)$
$\lesssim\beta^{\frac{3}{4}}e^{\frac{3t^{2}}{4}}\int_{-\infty}^{\infty}\mathopen{}\mathclose{{}\left(\beta^{\frac{3}{2}}|z|^{3}+|1-\beta z^{2}|^{3}}\right)e^{-\frac{3\beta}{2}z^{2}-\frac{1}{2}(z-t)^{2}}\,\mathrm{d}z$
$\lesssim\beta^{\frac{3}{4}}e^{\frac{t^{2}}{4}}\int_{0}^{\infty}h\mathopen{}\mathclose{{}\left(u+\frac{t}{3\beta+1}}\right)e^{-\frac{3\beta+1}{2}u^{2}}\,\mathrm{d}u,$
where we factor out $e^{-\frac{3\beta}{2(3\beta+1)}t^{2}}\asymp e^{-\frac{t^{2}}{2}}$ of the integral by substituting
$h(z)=\beta^{\frac{3}{2}}z^{3}+(1+\beta z^{2})^{3}\quad\text{and}\quad u=z-\frac{t}{3\beta+1}.$
By linearity of integration and A.1, each monomial $u^{n}$ integrates to $O(\beta^{-(n+1)/2})$. By monotonicity of $h$ on $\mathbb{R}_{\geqslant 0}$ and since $t\in T$, for some constant $C>0$,
$\eta_{3}\lesssim\beta^{\frac{3}{4}}e^{\frac{t^{2}}{4}}\int_{0}^{\infty}h\mathopen{}\mathclose{{}\left(u+\frac{t}{3\beta+1}}\right)e^{-\frac{3\beta+1}{2}u^{2}}\,\mathrm{d}u$ $\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}h\mathopen{}\mathclose{{}\left(C\beta^{-\frac{1}{2}}+\frac{t}{3\beta+1}}\right)$
$\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}h\mathopen{}\mathclose{{}\left(2C\beta^{-\frac{1}{2}}}\right)$
$\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}$
upon noting $w\mapsto h(w/\sqrt{\beta})$ has constant coefficients. This proves 3.1.∎
A.3 Proof of Lemma 3.4
Fix $n,\beta$ sufficiently large. For a multi-index $\alpha$, define
$h(\mathbf{x})$ $=\mathbf{x}^{\alpha}\mathopen{}\mathclose{{}\left(q_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right)(\mathbf{x}),$
$\mathscr{F}h(\mathbf{z})$ $=\partial^{\alpha}\mathopen{}\mathclose{{}\left(\mathscr{F}(q_{t})-\frac{\varphi}{6\sqrt{n}}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}}\right)(\mathbf{z}),$
where
$(\mathscr{F}h)(\mathbf{z})=\int_{\mathbb{R}^{2}}e^{-\mathrm{i}\langle\mathbf{x},\mathbf{z}\rangle}h(\mathbf{x})\,\mathrm{d}\mathbf{x}$
denotes the Fourier transform of $h$. By Fourier inversion, it suffices to show that for any multi-index $\alpha$ with order $|\alpha|\leqslant 3$,
$\mathopen{}\mathclose{{}\left|h(\mathbf{x})}\right|=\mathopen{}\mathclose{{}\left|\frac{1}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}e^{-\mathrm{i}\langle\mathbf{z},\mathbf{x}\rangle}\mathscr{F}{h}(\mathbf{z})\,\mathrm{d}\mathbf{z}}\right|\lesssim\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\mathscr{F}{h}(\mathbf{z})}\right|\,\mathrm{d}\mathbf{z}$ (A.1)
is $O(\exp(-\omega(\beta)/2))$. We apply [BR10, Theorem 9.12]—which is not asymptotic and has explicit constants—so we may use it even though $q_{t}$ depends on $n$. Upon verifying the conditions on $Y$ via 2.2 and 3.1, we have that
$\mathopen{}\mathclose{{}\left|\mathscr{F}{h}(\mathbf{z})}\right|\lesssim n^{-\frac{1}{2}}\eta_{3}^{\frac{1}{2}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{4}}$
provided $\|\mathbf{z}\|\leqslant a\sqrt{n}$ for some $a\asymp\eta_{3}^{-1/2}$. By 3.1, we have that
$\int_{\|\mathbf{z}\|\leqslant a\sqrt{n}}\mathopen{}\mathclose{{}\left|\mathscr{F}{h}(\mathbf{z})}\right|\,\mathrm{d}\mathbf{z}\lesssim n^{-\frac{1}{2}}\eta_{3}^{\frac{1}{2}}\int_{\mathbb{R}^{2}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{4}}\,\mathrm{d}\mathbf{z}\lesssim e^{-\frac{\omega(\beta)}{2}}.$ (A.2)
Recall that $q_{t}$ is the density of $n^{-1/2}\sum_{i=1}^{n}Y_{i}$, and let $f$ denote the density of $Y$. Now, we proceed as the proof of [BR10, Theorem 19.2]. As $p_{t}$ is bounded, so is $f$, and so [BR10, Theorem 19.1] gives $\mathscr{F}f\in L^{1}(\mathbb{R}^{2})$ and
$\delta:=\sup_{\|\mathbf{z}\|>a}\mathopen{}\mathclose{{}\left|\mathscr{F}{f}(\mathbf{z})}\right|<1.$
By properties of the Fourier transform and the product rule,
$\int_{\|\mathbf{z}\|>a\sqrt{n}}\mathopen{}\mathclose{{}\left|\partial^{\alpha}\mathscr{F}{q_{t}}(\mathbf{z})}\right|\,\mathrm{d}\mathbf{z}$ $\lesssim\eta_{|\alpha|}n^{\frac{|\alpha|}{2}}\delta^{n-|\alpha|-1}\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\mathscr{F}{f}\mathopen{}\mathclose{{}\left(\frac{\mathbf{z}}{\sqrt{n}}}\right)}\right|\,\mathrm{d}\mathbf{z}$ (A.3)
$\lesssim\mathopen{}\mathclose{{}\left(n\beta e^{\frac{t^{2}}{2}}}\right)^{O(1)}\delta^{n-O(1)}$
$\lesssim e^{-\frac{\omega(\beta)}{4}}$
for sufficiently large $n$. Finally, we bound similar to Lemma 3.2:
$\int_{\|\mathbf{z}\|>a\sqrt{n}}$ $\mathopen{}\mathclose{{}\left|\partial^{\alpha}\frac{\varphi}{6\sqrt{n}}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}}\right|(\mathbf{z})\,\mathrm{d}\mathbf{z}$ (A.4)
$\lesssim n^{-\frac{1}{2}}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\partial^{\alpha}H^{(k,3-k)}\varphi}\right|(\mathbf{z})\,\mathrm{d}\mathbf{z}$
$\lesssim n^{-\frac{1}{2}}\eta_{3}\int_{\mathbb{R}^{2}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{2}}\,\mathrm{d}\mathbf{z}$
$\lesssim e^{-\frac{\omega(\beta)}{4}}.$
Combining A.2, A.3 and A.4 proves A.1 and hence Lemma 3.4.∎
A.4 Proof of Proposition 4.1
Point 1 in Theorem 2.1 can readily be seen to hold because of the explicit form of both of the fields. We focus on showing Point 3, the proof of which can be repeated essentially verbatim to deduce Point 2.
[…]
$\begin{bmatrix}G(t)\\ G^{\prime}(t)\end{bmatrix}=\begin{bmatrix}g(Z)\\ g^{\prime}(Z)\end{bmatrix}$
[…]
We have
[…]
$=\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\int_{\mathbb{R}}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|^{n}\,\mathrm{d}\xi.$
Recalling that $n\to\infty$ in our regime, we can suppose $n\geqslant 4$, and for the above integral to be finite, it suffices to show that
$\mathopen{}\mathclose{{}\left|\int_{\mathbb{R}}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|\lesssim\frac{1}{\sqrt{\xi_{1}+\xi_{2}}}\hskip 28.45274pt\text{ as }\xi_{1},\xi_{2}\to\infty.$
Observe that the critical points of $g$ are ±$\pm\sqrt{\frac{1}{\beta}}$, whereas those of $g^{\prime}$ are 00 and ±$\pm\sqrt{\frac{3}{\beta}}$. Since $\beta\to\infty$ as well, pick $\varepsilon>0$ sufficiently small and such that $\varepsilon>\sqrt{\frac{3}{\beta}}$. We first see that
$\mathopen{}\mathclose{{}\left|\int_{|x|>\varepsilon}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
$=\mathopen{}\mathclose{{}\left|\int_{|x|>\varepsilon}\frac{1}{\mathrm{i}}\frac{1}{\xi_{1}g^{\prime}(x)+\xi_{2}g^{\prime\prime}(x)}\frac{\,\mathrm{d}}{\,\mathrm{d}x}\mathopen{}\mathclose{{}\left(e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}}\right)e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
$\lesssim\frac{1}{\xi_{1}+\xi_{2}},$
where we used integration by parts to obtain the last inequality—this is in fact an elementary version of the method of non-stationary phase. For the integral over $\{|x|\leqslant\varepsilon\}$, we look to use the method of stationary phase as $\xi_{1},\xi_{2}\to\infty$, by distinguishing the three regimes $\xi_{1}\gg\xi_{2}$, $\xi_{2}\gg\xi_{1}$, and $\xi_{1}\sim\xi_{2}$. When $\xi_{1}\gg\xi_{2}$, we have
$\mathopen{}\mathclose{{}\left|\int_{|x|\leqslant\varepsilon}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
$=\mathopen{}\mathclose{{}\left|\int_{|x|\leqslant\varepsilon}e^{-\mathrm{i}(\xi_{1}+\xi_{2})\mathopen{}\mathclose{{}\left(\frac{\xi_{1}}{\xi_{1}+\xi_{2}}g(x)+\frac{\xi_{2}}{\xi_{1}+\xi_{2}}g^{\prime}(x)}\right)}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
$=\mathopen{}\mathclose{{}\left|\int_{|x|\leqslant\varepsilon}e^{-\mathrm{i}(\xi_{1}+\xi_{2})(g(x)+O(\xi_{1}\varepsilon))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
$\lesssim\frac{1}{\sqrt{\xi_{1}+\xi_{2}}}+o\mathopen{}\mathclose{{}\left(\frac{1}{\sqrt{\xi_{1}+\xi_{2}}}}\right)$
by the method of stationary phase applied to the phase $g$, since $g^{\prime\prime}$ is non-degenerate at the critical points ±$\pm\sqrt{\frac{1}{\beta}}$. Similarly when $\xi_{2}\gg\xi_{1}$, we have
$\mathopen{}\mathclose{{}\left|\int_{|x|\leqslant\varepsilon}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$ $=\mathopen{}\mathclose{{}\left|\int_{|x|\leqslant\varepsilon}e^{-\mathrm{i}(\xi_{1}+\xi_{2})(g^{\prime}(x)+O(\xi_{2}\varepsilon))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
$\lesssim\frac{1}{\sqrt{\xi_{1}+\xi_{2}}}+o\mathopen{}\mathclose{{}\left(\frac{1}{\sqrt{\xi_{1}+\xi_{2}}}}\right)$
by the method of stationary phase applied to the phase $g^{\prime}$, since $g^{\prime\prime\prime}$ is non-degenerate at the critical points 00 and ±$\pm\sqrt{\frac{3}{\beta}}$. The same argument then carries through when $\xi_{1}\sim\xi_{2}$, using the phase $g+g^{\prime}$. This yields the desired conclusion.
To deduce that $(t,\mathbf{x})\mapsto p_{t}(\mathbf{x})$ is continuous on $T\times\mathbb{R}^{2}$, we note that
$p_{t}(\mathbf{x})=\frac{1}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}e^{\mathrm{i}\langle\mathbf{x},\mathbf{z}\rangle}\int_{\mathbb{R}^{n}}\exp\mathopen{}\mathclose{{}\left(-\mathrm{i}\mathopen{}\mathclose{{}\left\langle\mathbf{z},\begin{bmatrix}\sum_{j=1}^{n}g(\xi_{j})\\ \sum_{j=1}^{n}g^{\prime}(\xi_{j})\end{bmatrix}}\right\rangle}\right)\gamma_{t}(\xi_{1})\cdots\gamma_{t}(\xi_{n})\,\mathrm{d}\xi\,\mathrm{d}\mathbf{z}.$
We can conclude by the Lebesgue dominated convergence theorem.∎

### v2

@@
 1.1 Setup and main result
-For $\beta>0$ and $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$, the Gaussian kernel density estimator (KDE) with bandwidth $h=\beta^{-1/2}$ is defined as
-$\widehat{P}_{n}(t):=\frac{1}{n}\sum_{i=1}^{n}\mathsf{K}_{h}\ast\delta_{X_{i}}(t)=\frac{\sqrt{\beta}}{n\sqrt{2\pi}}\sum_{i=1}^{n}e^{-\frac{\beta}{2}(t-X_{i})^{2}},\hskip 28.45274ptt\in\mathbb{R}.$ (1.1)
+For $\beta>0$ and $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$, the Gaussian kernel density estimator (KDE) with bandwidth $h=\beta^{-\frac{1}{2}}$ is defined as
+$\widehat{P}_{n}(t)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\mathsf{K}_{h}\ast\delta_{X_{i}}(t)=\frac{\sqrt{\beta}}{n\sqrt{2\pi}}\sum_{i=1}^{n}e^{-\frac{\beta}{2}(t-X_{i})^{2}},\hskip 28.45274ptt\in\mathbb{R}.$ (1.1)
 Here, “Gaussian” refers to the choice of kernel $\mathsf{K}_{h}$.
@@
 Let $\widehat{P}_{n}$ be the Gaussian KDE defined in (1.1), with bandwidth $h>0$, of $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$.
-Asymptotically as $n\to\infty$, the expected number of modes of $\widehat{P}_{n}$ in a fixed interval $[a,b]$ is $1+o(1)$ if $1\ll\beta\ll n^{2/3}$, and $\widetilde{\Theta}\mathopen{}\mathclose{{}\left(\sqrt{\beta}}\right)$ if $n^{2/3}\lesssim\beta\ll n^{2}/\log^{6}n$.
+Asymptotically as $n\to\infty$, the expected number of modes of $\widehat{P}_{n}$ in a fixed interval $[a,b]$ is $1+o(1)$ if $1\ll\beta\ll n^{\frac{2}{3}}$, and $\widetilde{\Theta}\mathopen{}\mathclose{{}\left(\sqrt{\beta}}\right)$ if $n^{\frac{2}{3}}\lesssim\beta\ll n^{2}/\log^{6}n$.
 In [MMF92, Mam95, KM97], the authors additionally conduct more refined casework on the bandwidth to provide more precise estimates, such as pinpointing the leading constants.
@@
 Theorem 1.2.
-Let $\widehat{P}_{n}$ be the Gaussian KDE defined in (1.1), with bandwidth $\beta^{-1/2}$, of $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$.
-Suppose $n^{c}\lesssim\beta\lesssim n^{2-c}$ for arbitrarily small $c>0$, and 4.2, then asymptotically as $n,\beta\to\infty$,
+Let $\widehat{P}_{n}$ be the Gaussian KDE defined in (1.1), with bandwidth $\beta^{-\frac{1}{2}}$, of $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$.
+Suppose $n^{c}\lesssim\beta\lesssim n^{2-c}$ for arbitrarily small $c>0$.Then asymptotically as $n,\beta\to\infty$,
   1. 1.
@@
 Figure 1: A realization of the random function (1.1) for $n=10^{4}$, with $\beta=100$ (left) and $\beta=300$ (right).
-4.2 is related to the decay of the tails of the modulus of continuity of $\widehat{P}_{n}^{\prime\prime}(\cdot)$, and is needed to apply the Kac-Rice formula (Theorem 2.1), which states that the expected number of modes of $\widehat{P}_{n}$ is some conditional expectation with respect to the joint law of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{\prime\prime}(t))$.
-We postpone a further discussion to Section 4.1.
+Several comments are in order.
 Remark 1.3.
   • To better appreciate the range of values for $\beta$ in this theorem as well as subsequent ones, we use minimax theory as a benchmark; see, e.g., [Tsy09].
-The reparametrization $h=\sqrt{\beta}$ is motivated by the connection to the Transformer model described in Section 1.2.
+The reparametrization $h=\beta^{-\frac{1}{2}}$ is motivated by the connection to the Transformer model described in Section 1.2.
 Using an optimal bias-variance tradeoff [Tsy09, Chapter 1], we see that the optimal scaling of the bandwidth parameter $h$ depends on the smoothness of the underlying density of interest: if the underlying density has $s$ bounded (fractional) derivatives, then the optimal choice of $h$ is given by $h\asymp n^{-\frac{1}{2s+1}}$.
 This gives $\beta\asymp n^{\frac{2}{2s+1}}$.
-For $s\in(0,\infty)$, we get $\beta\in[n^{c},n^{2-c}]$ for some $c>0$.
-In particular, the transition of the number of modes from 1 to $\sqrt{\beta}$ in Theorem 1.1 is achieved for $\beta\approx n^{2/3}$, which is the optimal choice for Lipschitz densities.
+For $s>0$, we get $\beta\in[n^{c},n^{2-c}]$ for some $c>0$.
+In particular, the transition of the number of modes from 1 to $\sqrt{\beta}$ in Theorem 1.1 is achieved for $\beta\approx n^{\frac{2}{3}}$, which is the optimal choice for Lipschitz densities.
 The message of our main Theorem 1.2 below is that this scaling in $\sqrt{\beta}$ is the prevailing one for the whole range $\beta\in[n^{c},n^{2-c}]$ if one does not restrict counting modes in a bounded interval $[a,b]$.
@@
 To prove Theorem 1.2, we truncate the real line to the interval
-$T:=\mathopen{}\mathclose{{}\left[-\sqrt{2\log n-\log\beta-\omega(\beta)},\sqrt{2\log n-\log\beta-\omega(\beta)}}\right]$ (1.2)
+$T\coloneqq\mathopen{}\mathclose{{}\left[-\sqrt{2\log n-\log\beta-\omega(\beta)},\sqrt{2\log n-\log\beta-\omega(\beta)}}\right]$ (1.2)
 where $\omega$ is a fixed, slow growing function such that
@@
 and so $T$ is well-defined for large enough $\beta$. Motivated by Theorem 1.2, we also define the interval
-$T^{\prime}:=\mathopen{}\mathclose{{}\left[-\sqrt{2\log n-3\log\beta},\sqrt{2\log n-3\log\beta}}\right]$ (1.3)
-if $\beta\leqslant n^{2/3}$ and define $T^{\prime}=\varnothing$ if $\beta>n^{2/3}$.
-We see how Theorem 1.2 implies Theorem 1.1 with $h=\beta^{-1/2}$.
+$T^{\prime}\coloneqq\mathopen{}\mathclose{{}\left[-\sqrt{2\log n-3\log\beta},\sqrt{2\log n-3\log\beta}}\right]$ (1.3)
+if $\beta\leqslant n^{\frac{2}{3}}$ and define $T^{\prime}=\varnothing$ if $\beta>n^{\frac{2}{3}}$.
+We see how Theorem 1.2 implies Theorem 1.1 with $h=\beta^{-\frac{1}{2}}$.
 From the former, we see that almost all the modes lie in $T\setminus T^{\prime}$.
-If $h\gg n^{-1/3}$ so that $\sqrt{2\log n-3\log\beta}\gg 1$, then $T\setminus T^{\prime}$ is disjoint from $[a,b]$, so there are few modes in $[a,b]$; if $h\ll n^{-1/3}$ so $\sqrt{2\log n-3\log\beta}\ll 1$, the fixed interval $[a,b]$ contains a near constant fraction of length of $T\setminus T^{\prime}$ and thus a near constant fraction of the modes.
+If $h\gg n^{-\frac{1}{3}}$ so that $\sqrt{2\log n-3\log\beta}\gg 1$, then $T\setminus T^{\prime}$ is disjoint from $[a,b]$, so there are few modes in $[a,b]$; if $h\ll n^{-\frac{1}{3}}$ so $\sqrt{2\log n-3\log\beta}\ll 1$, the fixed interval $[a,b]$ contains a near constant fraction of length of $T\setminus T^{\prime}$ and thus a near constant fraction of the modes.
 Figure 2: (Left) Plot of the average number of modes as a function of $\beta$ for $n=10^{3}$ (top) and $n=10^{4}$ (bottom).
@@
 $\frac{\,\mathrm{d}}{\,\mathrm{d}\tau}x_{i}(\tau)=\sum_{j=1}^{n}\frac{e^{\beta\langle x_{i}(\tau),x_{j}(\tau)\rangle}}{\displaystyle\sum_{k=1}^{n}e^{\beta\langle x_{i}(\tau),x_{j}(\tau)\rangle}}\mathsf{P}_{x_{i}(\tau)}^{\perp}(x_{j}(\tau)),$
-evolving on the unit sphere $\mathbb{S}^{d-1}$ because of $\mathsf{P}_{x}^{\perp}:=I_{d}-xx^{\top}$.
+evolving on the unit sphere $\mathbb{S}^{d-1}$ because of $\mathsf{P}_{x}^{\perp}\coloneqq I_{d}-xx^{\top}$.
 Here, $\tau\geqslant 0$ plays the role of layers, the $n$ particles $x_{i}(\tau)$ represent tokens evolving through a dynamical system.
@@
 Proposition 1.6.
-If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for arbitrarily small $c>0$, then under 4.2,
+If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for arbitrarily small $c>0$, then
   1. 1.
@@
 We bound the error due to the third order term of the expansion directly, and deal with the higher order terms by appealing to the error bounds of densities in the Edgeworth approximation similar to [BR10, Theorem 19.2].
+This strategy has been used in [BCP19], in a different context.
 We note that [KM97] employ the same theorem to justify the Gaussian process approximation over a fixed interval.
@@
   4. 4.
-    $\mathbb{P}(\upomega(\eta)>\varepsilon)=O(\eta)$ as $\eta\to 0^{+}$ for any $\varepsilon>0$, where $\upomega(\cdot)$ denotes the modulus of continuity111defined, for $f:\mathbb{R}\to\mathbb{R}$, as $\upomega(\eta)=\sup_{t,s\colon|t-s|\leq\eta}|f(t)-f(s)|$.
+    $\mathbb{P}(\upomega(\eta)>\varepsilon)=O(\eta)$ as $\eta\searrow 0^{+}$ for any $\varepsilon>0$, where $\upomega(\cdot)$ denotes the modulus of continuity111defined, for $f:\mathbb{R}\to\mathbb{R}$, as $\upomega(\eta)=\sup_{t,s\colon|t-s|\leqslant\eta}|f(t)-f(s)|$.
 of $\Psi^{\prime}(\cdot)$.
 Define the number of up-crossings in $T$ of $\Psi$ at level $u\in\mathbb{R}$ as
-$U_{u}(\Psi,T):=\mathopen{}\mathclose{{}\left|\{t\in T:\Psi(t)=u,\Psi^{\prime}(t)>0\}}\right|.$
+$U_{u}(\Psi,T)\coloneqq\mathopen{}\mathclose{{}\left|\{t\in T:\Psi(t)=u,\Psi^{\prime}(t)>0\}}\right|.$
 Then, with expectation taken over the randomness of $\Psi$,
@@
 The mean and covariance matrix of the random vector $(F_{n}(t),F_{n}^{\prime}(t))$ are given respectively by
-$\mu_{t}$ $:=\sqrt{n}\begin{bmatrix}\mathbb{E}G(t)\\ \mathbb{E}G^{\prime}(t)\end{bmatrix}\asymp n^{\frac{1}{2}}\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}t\\ -t^{2}\end{bmatrix}$ (2.5)
-$\Sigma_{t}$ $:=\begin{bmatrix}\operatorname{Var}G(t)&\operatorname{Cov}(G(t),G^{\prime}(t))\\ \operatorname{Cov}(G(t),G^{\prime}(t))&\operatorname{Var}G^{\prime}(t)\end{bmatrix}\asymp\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}1&-t\\ -t&\beta\end{bmatrix}.$
+$\mu_{t}$ $\coloneqq\sqrt{n}\begin{bmatrix}\mathbb{E}G(t)\\ \mathbb{E}G^{\prime}(t)\end{bmatrix}\asymp n^{\frac{1}{2}}\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}t\\ -t^{2}\end{bmatrix}$ (2.5)
+$\Sigma_{t}$ $\coloneqq\begin{bmatrix}\operatorname{Var}G(t)&\operatorname{Cov}(G(t),G^{\prime}(t))\\ \operatorname{Cov}(G(t),G^{\prime}(t))&\operatorname{Var}G^{\prime}(t)\end{bmatrix}\asymp\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}1&-t\\ -t&\beta\end{bmatrix}.$
 We proceed to centering and rescaling the density $p_{t}$. Let $Y_{i}(t)$, $i\in[n]$, be independent copies of
@@
 Now, let $\varphi:\mathbb{R}^{2}\to\mathbb{R}$ be the density of $N(0,I_{2})$, i.e.,
-$\varphi(x):=\frac{1}{\sqrt{2\pi}}e^{-\frac{\|x\|^{2}}{2}}.$
+$\varphi(x)\coloneqq\frac{1}{\sqrt{2\pi}}e^{-\frac{\|x\|^{2}}{2}}.$
 We aim to approximate the Kac-Rice integral 2.1 as follows:
@@
 Since $\det\Sigma_{t}\asymp\beta^{-2}e^{-t^{2}}$, we have
-$\Omega:=\Sigma_{t}^{-1}\asymp\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}\begin{bmatrix}\beta&t\\ t&1\end{bmatrix}.$
+$\Omega\coloneqq\Sigma_{t}^{-1}\asymp\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}\begin{bmatrix}\beta&t\\ t&1\end{bmatrix}.$
 Now as $t\in T$ and so $t^{2}\ll\beta$, we find
@@
 $\int_{T}e^{-A_{t}}\,\mathrm{d}t\asymp\sqrt{\log\beta}\quad\text{and}\quad\int_{T^{\prime}}e^{-A_{t}}\,\mathrm{d}t\lesssim 1.$ (2.11)
-As $A_{t}>0$, the integral is at most the length of $T$, which is $O(\sqrt{\log\beta})$ by 1.2. Let $t_{s}:=\sqrt{2\log n-s\log\beta}$. As the integrand is positive, for constants $C,C^{\prime}>0$
+As $A_{t}>0$, the integral is at most the length of $T$, which is $O(\sqrt{\log\beta})$ by 1.2.
+Let $t_{s}\coloneqq\sqrt{2\log n-s\log\beta}$.
+As the integrand is positive, for constants $C,C^{\prime}>0$
 $\int_{T}e^{-A_{t}}\,\mathrm{d}t$ $\geqslant\int_{t_{2}}^{t_{5/2}}\exp\mathopen{}\mathclose{{}\left(-C\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}}\right)\,\mathrm{d}t$
@@
 The next term in the Edgeworth series is $n^{-1/2}\psi$ with
-$\psi(\mathbf{x}):=\frac{\varphi(\mathbf{x})}{6}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}(\mathbf{x})\quad\text{where}\quad H^{\alpha}:=(-1)^{|\alpha|}\frac{\partial^{\alpha}\varphi}{\varphi}$ (3.2)
+$\psi(\mathbf{x})\coloneqq\frac{\varphi(\mathbf{x})}{6}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}(\mathbf{x})\quad\text{where}\quad H^{\alpha}\coloneqq(-1)^{|\alpha|}\frac{\partial^{\alpha}\varphi}{\varphi}$ (3.2)
 If $|\alpha|=3$, $\kappa_{t}^{\alpha}$ are the third moments of $Y$, which we bound in Section A.1.
@@
 Let $\eta_{3}$ be the largest third cumulant of $Y$. Then
-$\eta_{3}:=\max\mathopen{}\mathclose{{}\left\{\kappa_{t}^{(k,3-k)}:k\in\{0,1,2,3\}}\right\}\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}.$
+$\eta_{3}\coloneqq\max\mathopen{}\mathclose{{}\left\{\kappa_{t}^{(k,3-k)}:k\in\{0,1,2,3\}}\right\}\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}.$
 We bound $H^{(k,3-k)}\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-1/2}[(0,y)-\mu_{t}]}\right)$ by a polynomial in $\tilde{y}=\Theta(\beta^{1/2}e^{{t^{2}}/{4}}y)$.
 Now, by a similar method as Lemma 2.3, we obtain the following bound.
-It says that when we integrate the Edgeworth series $p_{t}=\varphi+n^{-1/2}\psi+\dots$ over $y$ and $t$, the contribution $\varphi$ dominates $n^{-1/2}\psi$, hinting at the validity of the approximation.
+It says that when we integrate the Edgeworth series $p_{t}=\varphi+n^{-\frac{1}{2}}\psi+\dots$ over $y$ and $t$, the contribution $\varphi$ dominates $n^{-\frac{1}{2}}\psi$, hinting at the validity of the approximation.
 Lemma 3.2.
@@
 $\mathopen{}\mathclose{{}\left(1+\|\mathbf{x}\|^{3}}\right)\mathopen{}\mathclose{{}\left|q_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right|(\mathbf{x})\lesssim e^{-\frac{\omega(\beta)}{4}}.$ (3.3)
-We defer the discussion of the assumption to 4.2 and defer the proof to Section A.3.
+We defer the proof to Section A.3.
 Here, we comment on the differences with [BR10, Theorem 19.2] for $s=3$, which we follow in spirit but modify to allow $n$ dependence in $q_{t}$ in the form of $\beta$.
-There, it is shown that the left hand side is $o(n^{-1/2})$ provided third moments $\eta_{3}=O(1)$.
-With the bound 3.1 on $\eta_{3}$ in our case which is not constant, the error from the asymptotic series decays not in powers of $n^{-1/2}$ but powers of $O(n^{-1/2}e^{{t^{2}}/{4}}\beta^{1/4})=O(\exp(-\omega(\beta)/2))$.
+There, it is shown that the left hand side is $o(n^{-\frac{1}{2}})$ provided third moments $\eta_{3}=O(1)$.
+With the bound 3.1 on $\eta_{3}$ in our case which is not constant, the error from the asymptotic series decays not in powers of $n^{-\frac{1}{2}}$ but powers of $O(n^{-1/2}e^{{t^{2}}/{4}}\beta^{1/4})=O(\exp(-\omega(\beta)/2))$.
 Corollary 3.5.
@@
 Then $\upmu_{t}$ admits a density $p_{t}\in\mathscr{C}^{0}(\mathbb{R}^{2})$ satisfying $p_{t}({\bf{x}})\to 0$ as $\|\bf{x}\|\to\infty$.
-Moreover, conditions 1, 2 in Theorem 2.1 also hold for $\Psi(t)=F_{n}(t)$.
-We defer the proof to Section A.4. We work under assumption
-Assumption 4.2.
-Consider the setting of Proposition 4.1. We assume that condition 4 of Theorem 2.1 holds for $\Psi(t)=F_{n}(t)$.
-To check conditions on moduli of continuity such as 4 in Theorem 2.1 in the Gaussian setting, one usually resorts to using results such as the Borell-TIS inequality [AT09, Theorem 2.1.1].
-Checking the validity of this assumption in the present, non-Gaussian, setting does not appear straightforward.
-With Proposition 4.1 and 4.2, we deduce
+Moreover, conditions 1, 2, 4 in Theorem 2.1 also hold for $\Psi(t)=F_{n}(t)$, thus Theorem 2.1 applies.
+We defer the proof to Section A.4. With Proposition 4.1, we deduce
 Lemma 4.2.
@@
 Note that $\widehat{P}_{n}(t)=\sum_{i=1}^{n}g_{i}(t)$ where for $i\in[n]$ we define
-$g_{i}(t):=\sqrt{\frac{\beta}{2\pi n^{2}}}\mathsf{K}_{\beta^{-1/2}}\mathopen{}\mathclose{{}\left(t-X_{i}}\right).$ (4.2)
+$g_{i}(t)\coloneqq\sqrt{\frac{\beta}{2\pi n^{2}}}\mathsf{K}_{\beta^{-1/2}}\mathopen{}\mathclose{{}\left(t-X_{i}}\right).$ (4.2)
 For $i\not\in I$, $g_{i}$ is monotonically decreasing on $[X_{i},\infty)\supset(a,\infty)$, so $\sum_{i\not\in I}g_{i}(t)$ has no modes in $(a,\infty)$.
@@
 We check that entries of $\Sigma_{t}$ are asymptotically the corresponding second moments. Indeed, surpressing the dependence on $t$, we have that
-  • $\mathbb{E}G^{2}\asymp\beta^{-\frac{5}{2}}(t^{2}+\beta)e^{\frac{t^{2}}{2}}\gg\beta^{-3}t^{2}E^{2}_{t}\asymp(\mathbb{E}G)^{2},$
-  • $\mathbb{E}GG^{\prime}\asymp-\beta^{-\frac{7}{2}}e^{\frac{t^{2}}{2}}(\beta^{2}t+\beta t^{3})\gg\beta^{-4}E^{2}_{t}t(\beta-\beta t^{2})\asymp(\mathbb{E}G)(\mathbb{E}G^{\prime}),$
-  • $\mathbb{E}G^{\prime 2}\asymp\beta^{-\frac{9}{2}}e^{\frac{t^{2}}{2}}(\beta^{4}+\beta^{2}t^{4})\gg\beta^{-5}E^{2}_{t}\beta^{2}(1+t^{4})\asymp(\mathbb{E}G^{\prime})^{2},$
-where we bound $e^{\frac{t^{2}}{2}}\leqslant 1$. From these computations, and the remark after A.1, we readily obtain the asymptotics of $\Sigma_{t}$ as indicated in 2.2.∎
+  • $\mathbb{E}G^{2}\asymp\beta^{-\frac{5}{2}}(t^{2}+\beta)e^{-\frac{t^{2}}{2}}\gg\beta^{-3}t^{2}e^{-t^{2}}\asymp(\mathbb{E}G)^{2},$
+  • $\mathbb{E}GG^{\prime}\asymp-\beta^{-\frac{7}{2}}e^{-\frac{t^{2}}{2}}(\beta^{2}t+\beta t^{3})\gg\beta^{-4}e^{-t^{2}}t(\beta-\beta t^{2})\asymp(\mathbb{E}G)(\mathbb{E}G^{\prime}),$
+  • $\mathbb{E}G^{\prime 2}\asymp\beta^{-\frac{9}{2}}e^{-\frac{t^{2}}{2}}(\beta^{4}+\beta^{2}t^{4})\gg\beta^{-5}e^{-t^{2}}\beta^{2}(1+t^{4})\asymp(\mathbb{E}G^{\prime})^{2},$
+where we bound $e^{-\frac{t^{2}}{2}}\leqslant 1$. From these computations, and the remark after A.1, we readily obtain the asymptotics of $\Sigma_{t}$ as indicated in 2.2.∎
 A.2 Proof of 3.1
@@
 As $p_{t}$ is bounded, so is $f$, and so [BR10, Theorem 19.1] gives $\mathscr{F}f\in L^{1}(\mathbb{R}^{2})$ and
-$\delta:=\sup_{\|\mathbf{z}\|>a}\mathopen{}\mathclose{{}\left|\mathscr{F}{f}(\mathbf{z})}\right|<1.$
+$\delta\coloneqq\sup_{\|\mathbf{z}\|>a}\mathopen{}\mathclose{{}\left|\mathscr{F}{f}(\mathbf{z})}\right|<1.$
 By properties of the Fourier transform and the product rule,
@@
 Point 1 in Theorem 2.1 can readily be seen to hold because of the explicit form of both of the fields.
+Point 4 also readily holds, since $F_{n}^{\prime\prime}$ is a Lipschitz function for every realization of $X_{i}$, as a sum of Lipschitz functions.
 We focus on showing Point 3, the proof of which can be repeated essentially verbatim to deduce Point 2.
+Proof of Point 3
 […]
@@
 […]
-$=\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\int_{\mathbb{R}}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|^{n}\,\mathrm{d}\xi.$
+$=\frac{1}{(2\pi)^{\frac{n}{2}}}\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\int_{\mathbb{R}}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|^{n}\,\mathrm{d}\xi.$
 Recalling that $n\to\infty$ in our regime, we can suppose $n\geqslant 4$, and for the above integral to be finite, it suffices to show that
@@
 $p_{t}(\mathbf{x})=\frac{1}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}e^{\mathrm{i}\langle\mathbf{x},\mathbf{z}\rangle}\int_{\mathbb{R}^{n}}\exp\mathopen{}\mathclose{{}\left(-\mathrm{i}\mathopen{}\mathclose{{}\left\langle\mathbf{z},\begin{bmatrix}\sum_{j=1}^{n}g(\xi_{j})\\ \sum_{j=1}^{n}g^{\prime}(\xi_{j})\end{bmatrix}}\right\rangle}\right)\gamma_{t}(\xi_{1})\cdots\gamma_{t}(\xi_{n})\,\mathrm{d}\xi\,\mathrm{d}\mathbf{z}.$
-We can conclude by the Lebesgue dominated convergence theorem.∎
+We can conclude by the Lebesgue dominated convergence theorem.

### v3

@@
 In this paper we are interested in determining the expected number of modes (local maxima) of $\widehat{P}_{n}$ over $\mathbb{R}$.
-While this is a classical question, addressed in even more general settings than (1.1)—such as non-Gaussian kernels, compactly supported samples, and higher dimensions [MMF92, Mam95, KM97]—a definite answer has not been given in the literature.
+While this is a classical question, addressed in even more general settings than (1.1)—such as non-Gaussian kernels, compactly supported samples, and higher dimensions [mammen91, mammen95, mammen97]—a definite answer has not been given in the literature.
 Indeed, the best-known results fall into one of two settings: either considering samples drawn from a compactly supported density (instead of $N(0,1)$ as done here), or counting the modes within a fixed compact interval.
 In the special case of the Gaussian KDE (1.1), one has in the latter setting for instance
-Theorem 1.1 ([Mam95, Thm. 1]).
-Let $\widehat{P}_{n}$ be the Gaussian KDE defined in (1.1), with bandwidth $h>0$, of $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$.
-Asymptotically as $n\to\infty$, the expected number of modes of $\widehat{P}_{n}$ in a fixed interval $[a,b]$ is $1+o(1)$ if $1\ll\beta\ll n^{\frac{2}{3}}$, and $\widetilde{\Theta}\mathopen{}\mathclose{{}\left(\sqrt{\beta}}\right)$ if $n^{\frac{2}{3}}\lesssim\beta\ll n^{2}/\log^{6}n$.
-In [MMF92, Mam95, KM97], the authors additionally conduct more refined casework on the bandwidth to provide more precise estimates, such as pinpointing the leading constants.
-In fact, [MMF92] does count modes in $\mathbb{R}$, but the underlying distribution of the samples $X_{i}$ is supported on a closed interval (thus excluding $N(0,1)$), so there are no modes outside the interval anyway.
-Our main result provides the answer to the case of counting modes of (1.1) over $\mathbb{R}$, and reads as follows. In particular, it generalizes Theorem 1.1.
+Theorem 1.1 ([mammen95, Thm. 1]).
+Let $\widehat{P}_{n}$ be the Gaussian KDE defined in (1.1), with bandwidth $h\coloneqq\beta^{-\frac{1}{2}}>0$, of $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$.
+Asymptotically as $n\to\infty$, the expected number $N$ of modes of $\widehat{P}_{n}$ in a fixed interval $[a,b]$ is
+  • $\mathbf{1}\{0\in[a,b]\}+o(1)$ if $\beta\ll n^{\frac{2}{5}}$,
+  • $\Theta(1)$ if $\beta\asymp n^{\frac{2}{5}}$
+  • $\Theta\mathopen{}\mathclose{{\left(n^{-\frac{1}{2}}\beta^{\frac{5}{4}}}}\right)=o\mathopen{}\mathclose{{\left(\sqrt{\beta}}}\right)$ if $n^{\frac{2}{5}}\ll\beta\ll n^{\frac{2}{3}}$,
+  • and $\Theta\mathopen{}\mathclose{{\left(\sqrt{\beta}}}\right)$ if $n^{\frac{2}{3}}\lesssim\beta\ll n^{2}/\log^{6}n$.
+In [mammen91, mammen95, mammen97], the authors additionally conduct more refined casework on the bandwidth to provide more precise estimates, such as pinpointing the leading constants.
+In fact, [mammen91] does count modes in $\mathbb{R}$, but the underlying distribution of the samples $X_{i}$ is supported on a closed interval (thus excluding $N(0,1)$), so there are no modes outside the interval anyway.
+In the case of counting modes of (1.1) over $\mathbb{R}$, [carreira2003number] provides an upper bound of $n$ using scale-space theory by showing adding components one-by-one to a Gaussian mixture increases the mode-count by at most one each time.
+Our main result velow provides a precise answer in this case.
+For the sake of clarity, we stick to the regime where
+$2\log n-\log\beta\asymp\log\beta\asymp\log n.$
+Through refined computations, one can determine the modes in the regime $1\ll\beta\lesssim n^{2}/\log^{\Theta(1)}(n)$ and also pinpoint the leading constant.
+We state our main theorem, and will comment on how to do expand the regime in appropriate places.
 Theorem 1.2.
@@
     Almost all modes lie in two intervals of length $\Theta\mathopen{}\mathclose{{\left(\sqrt{\log\beta}}}\right)$—namely, the expected number of modes $t\in\mathbb{R}$, such that $t^{2}\not\in\mathopen{}\mathclose{{\left[2\log n-3\log\beta,2\log n-\log\beta}}\right]$, is $o\mathopen{}\mathclose{{\left(\sqrt{\beta\log\beta}}}\right)$.
-Figure 1: A realization of the random function (1.1) for $n=10^{4}$, with $\beta=100$ (left) and $\beta=300$ (right).
-Several comments are in order.
+Figure 1: A realization of the kernel density estimator $\widehat{P}_{n}$ in (1.1) for $n=10^{4}$, with $\beta=100$ (left) and $\beta=300$ (right).
+Larger $\beta$ narrows the Gaussian kernel, which sharpens $\widehat{P}_{n}$ and reveals more small peaks on the shoulders, while the central peak remains single.
+Theorem˜1.2 later quantifies where and how many such peaks appear.
+In fact, we can bound on the rate of convergence of the little-$o$ in Point 2. This is spelled-out in Propositions˜1.7 and 1.8. Several comments are in order.
 Remark 1.3.
   • To better appreciate the range of values for $\beta$ in this theorem as well as subsequent ones, we use minimax theory as a benchmark; see, e.g., [Tsy09].
-The reparametrization $h=\beta^{-\frac{1}{2}}$ is motivated by the connection to the Transformer model described in Section 1.2.
-Using an optimal bias-variance tradeoff [Tsy09, Chapter 1], we see that the optimal scaling of the bandwidth parameter $h$ depends on the smoothness of the underlying density of interest: if the underlying density has $s$ bounded (fractional) derivatives, then the optimal choice of $h$ is given by $h\asymp n^{-\frac{1}{2s+1}}$.
+The reparametrization $h=\beta^{-\frac{1}{2}}$ is motivated by the connection to the Transformer model described in Section˜1.2.
+Using an optimal bias-variance tradeoff [Tsy09, Chapter 1] [^1] With $h=\beta^{-\frac{1}{2}}$, the usual bias–variance calculus for $s$-smooth densities gives $h\asymp n^{-\frac{1}{2s+1}}$ and hence $\beta\asymp n^{\frac{2}{2s+1}}$  [Tsy09, Ch.
+1]., we see that the optimal scaling of the bandwidth parameter $h$ depends on the smoothness of the underlying density of interest: if the underlying density has $s$ bounded (fractional) derivatives, then the optimal choice of $h$ is given by $h\asymp n^{-\frac{1}{2s+1}}$.
 This gives $\beta\asymp n^{\frac{2}{2s+1}}$.
 For $s>0$, we get $\beta\in[n^{c},n^{2-c}]$ for some $c>0$.
-In particular, the transition of the number of modes from 1 to $\sqrt{\beta}$ in Theorem 1.1 is achieved for $\beta\approx n^{\frac{2}{3}}$, which is the optimal choice for Lipschitz densities.
-The message of our main Theorem 1.2 below is that this scaling in $\sqrt{\beta}$ is the prevailing one for the whole range $\beta\in[n^{c},n^{2-c}]$ if one does not restrict counting modes in a bounded interval $[a,b]$.
+In particular, the transition of the number of modes from 1 to $\sqrt{\beta}$ in Theorem˜1.1 is achieved for $\beta\approx n^{\frac{2}{3}}$, which is the optimal choice for Lipschitz densities.
+The message of our main Theorem 1.2 above is that this scaling in $\sqrt{\beta}$ is the prevailing one for the whole range $\beta\in[n^{c},n^{2-c}]$ if one does not restrict counting modes in a bounded interval $[a,b]$.
   • Point 2.
-in Theorem 1.2 shows that most of the modes are at distance at least $C\log n$ from the origin provided $\beta>n^{\frac{2-C}{3}}$ for $C>0$ small.
+in Theorem˜1.2 shows that most of the modes are at distance at least $C\log n$ from the origin provided $\beta>n^{\frac{2-C}{3}}$ for $C>0$ small.
 This corresponds to a choice of a bandwidth adapted to smoothness $s<1$.
-This result is in agreement with and completes the picture drawn by Theorem 1.1.
+This result is in agreement with and completes the picture drawn by Theorem˜1.1.
 Remark 1.4.
-Through refined computations, one can determine the modes in the regime $1\ll\beta\ll n^{2}$ and also pinpoint the leading constant. For the sake of clarity, we stick to the regime where
-$2\log n-\log\beta\asymp\log\beta\asymp\log n,$
-and comment on how to do expand the regime in appropriate places.
+We further motivate Point 2. in Theorem˜1.2 by considering a qualitative picture of the distribution of the modes displayed in Figure˜4.
+  • Near the origin, we find most of the samples $X_{i}$ and they are densely packed in the shape of a Gaussian.
+The corresponding Gaussian summands in ˜1.1 cancel to create one mode, as shown already in Theorem˜1.1.
+  • In the two intervals of length $\Theta\mathopen{}\mathclose{{\left(\sqrt{\log\beta}}}\right)$, the samples $X_{i}$ are separated enough that the corresponding Gaussian summands do not cancel, but rather form $\Omega\mathopen{}\mathclose{{\left(\sqrt{\log\beta}}}\right)$ isolated bumps, as discussed in more generality in [devgyo, Section 9.3].
+  • Further away at the tails, the phenomena of isolated bumps occur, but there are so few samples $X_{i}$ that the number of modes created is a negligible fraction.
+Write $\widehat{P}_{n}^{\prime}(t)=\mathbb{E}\,\widehat{P}_{n}^{\prime}(t)+(\widehat{P}_{n}^{\prime}(t)-\mathbb{E}\widehat{P}_{n}^{\prime}(t))$.
+The first term (“bias”) reflects the deterministic drift toward a single broad mode, while the second (“variance”) creates random sign changes that generate extra modes.
+For the rescaled field $F_{n}(t)=-c\,\widehat{P}_{n}^{\prime}(t)$ in (2.2), Lemma˜2.2 yields
+$\mathrm{SNR}(t)^{2}\coloneqq\frac{|\mathbb{E}F_{n}(t)|^{2}}{\mathrm{Var}F_{n}(t)}\asymp nt^{2}\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}.$
+When $\mathrm{SNR}(t)\gg 1$ the bias dominates and no additional modes appear; when $\mathrm{SNR}(t)\ll 1$ the variance takes over and modes proliferate.
+The crossover $\mathrm{SNR}(t)\approx 1$ gives the inner edge $t^{2}\approx 2\log n-3\log\beta$ (up to $\log t$ terms).
+On the other hand, to form isolated bumps we also need at least one point in a kernel window of width $h=\beta^{-\frac{1}{2}}$, i.e.
+$n\varphi(t)h\approx 1$, which gives the outer edge $t^{2}\approx 2\log n-\log\beta$.
+Together these two thresholds select the belt $t^{2}\in[2\log n-3\log\beta,2\log n-\log\beta]$ and, in particular, center the localization at $|t|\approx\sqrt{2\log n-\log\beta}$.
+We revisit this discussion and Figure˜4 in Remark˜3.4.
+Remark 1.5 (Belt width).
+From Section˜2.3 the Kac–Rice density is proportional to $\sqrt{\beta}e^{-A_{t}}$ with $A_{t}\asymp\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}$, so the mass concentrates where $A_{t}=O(1)$.
+This pins down $t^{2}=2\log n-c\log\beta+O(\log\log\beta)$ with $c\in\{1,3\}$—precisely the endpoints of Theorem˜1.2—and converting from $t^{2}$ to $t$ turns the $2\log\beta$ gap into a belt of length $\Delta t\approx(2t_{*})^{-1}\cdot 2\log\beta\asymp\sqrt{\log\beta}$ around $t_{*}\approx\sqrt{2\log n-\log\beta}$.
 Remark 1.6.
-We further motivate Point 2. in Theorem 1.2 by considering a qualitative picture of the distribution of the modes displayed in Figure 4.
-  • Near the origin, we find most of the samples $X_{i}$ and they are densely packed in the shape of a Gaussian.
-The corresponding Gaussian summands in 1.1 cancel to create one mode, as shown already in Theorem 1.1.
-  • In the two intervals of length $\Theta\mathopen{}\mathclose{{}\left(\sqrt{\log\beta}}\right)$, the samples $X_{i}$ are separated enough that the corresponding Gaussian summands do not cancel, but rather form $\Omega\mathopen{}\mathclose{{}\left(\sqrt{\log\beta}}\right)$ isolated bumps, as discussed in more generality in [DG85, Section 9.3].
-  • Further away at the tails, the phenomena of isolated bumps occur, but there are so few samples $X_{i}$ that the number of modes created is a negligible fraction.
-We revisit this discussion and Figure 4 in Remark 3.3.
-To prove Theorem 1.2, we truncate the real line to the interval
-$T\coloneqq\mathopen{}\mathclose{{}\left[-\sqrt{2\log n-\log\beta-\omega(\beta)},\sqrt{2\log n-\log\beta-\omega(\beta)}}\right]$ (1.2)
-where $\omega$ is a fixed, slow growing function such that
-$1\ll\omega(\beta)\ll\log\log\beta,$
-and so $T$ is well-defined for large enough $\beta$. Motivated by Theorem 1.2, we also define the interval
-$T^{\prime}\coloneqq\mathopen{}\mathclose{{}\left[-\sqrt{2\log n-3\log\beta},\sqrt{2\log n-3\log\beta}}\right]$ (1.3)
-if $\beta\leqslant n^{\frac{2}{3}}$ and define $T^{\prime}=\varnothing$ if $\beta>n^{\frac{2}{3}}$.
-We see how Theorem 1.2 implies Theorem 1.1 with $h=\beta^{-\frac{1}{2}}$.
-From the former, we see that almost all the modes lie in $T\setminus T^{\prime}$.
-If $h\gg n^{-\frac{1}{3}}$ so that $\sqrt{2\log n-3\log\beta}\gg 1$, then $T\setminus T^{\prime}$ is disjoint from $[a,b]$, so there are few modes in $[a,b]$; if $h\ll n^{-\frac{1}{3}}$ so $\sqrt{2\log n-3\log\beta}\ll 1$, the fixed interval $[a,b]$ contains a near constant fraction of length of $T\setminus T^{\prime}$ and thus a near constant fraction of the modes.
+We compare our result with Theorem˜1.1.
+Let $J$ be the union of the symmetric intervals of length $\Theta(\sqrt{\log\beta})$ in Point 2, i.e.
+the “two belts” region where almost all modes lie.
+Our proof will show that the density of modes is $\Theta(\sqrt{\beta})$ whenever $t\in J$, and $o(\sqrt{\beta})$ whenever $t\not\in J$.
+This gives the Theorem˜1.2 upon integrating over $t$, and explains the threshold $\beta\asymp n^{\frac{2}{3}}$ in Theorem˜1.1:
+  • If $\beta\gtrsim n^{\frac{2}{3}}$, then $a,b]\subset J$ for sufficiently large $n$ and $\beta$, so the density of modes is $\Theta(\sqrt{\beta})$ everywhere on $[a,b]$, giving $\Theta(\sqrt{\beta})$ modes in total.
+  • If $\beta\ll n^{\frac{2}{3}}$, then $[a,b]$ is between (and outside of) the two belts of $J$ for sufficiently large $n$ and $\beta$, so the density of modes is $o(\sqrt{\beta})$ everywhere on $[a,b]$, giving $o(\sqrt{\beta})$ modes in total.
+In fact, Proposition˜1.7 shows this symmetric interval $T^{\prime}$ between the two belts has $O(\sqrt{\beta})$ modes in total.
+We can see that $T^{\prime}$ has length $\omega_{n\to\infty}(1)$ and the mode density is increasing as we move away from 0, so the number of modes in $[a,b]$ must be a $o(1)$-fraction of the modes in $T^{\prime}$, i.e.
+it is $o(\sqrt{\beta})$.
+Hence, this corollary of our result implies the last two bullet points of Theorem˜1.1.
+Similarly, by truncating to more refined intervals separated by $t^{2}=2\log n-5\log\beta$, we can hope to recover the threshold $\beta\asymp n^{\frac{2}{5}}$ given in the first three bullet points of Theorem˜1.1, but we do not pursue this here.
 Figure 2: (Left) Plot of the average number of modes as a function of $\beta$ for $n=10^{3}$ (top) and $n=10^{4}$ (bottom).
-(Right) Log-log plot for $n=10^{3}$ (top) and $n=10^{4}$ (bottom); the predicted linear regression line (red) corroborates a power-law of the form average # of $\text{average \# of modes}\approx 0.179\cdot\beta^{0.504}$, in line with Theorem 1.2.
+(Right) Log-log plot for $n=10^{3}$ (top) and $n=10^{4}$ (bottom); the predicted linear regression line (red) corroborates a power-law of the form $\text{average \# of modes}\approx 0.179\cdot\beta^{0.504}$, in line with Theorem˜1.2.
 1.2 Motivation
-The question of estimating the number of modes as a function of the bandwidth has a plethora of applications in statistical inference and multimodality tests—see [MMF92, Mam95, KM97] and the references therein.
+The question of estimating the number of modes as a function of the bandwidth has a plethora of applications in statistical inference and multimodality tests—see [mammen91, mammen95, mammen97] and the references therein.
 Another application which has stimulated some of the recent progress on the topic is data clustering.
-The latter can be achieved nonparametrically using a KDE, whose modes, and hence clusters, can be detected using the mean-shift algorithm [FH75, Che95, CM02, CP00, CPW03, CP07, RL14, CP15], which can essentially be seen as iterative local averaging.
+The latter can be achieved nonparametrically using a KDE, whose modes, and hence clusters, can be detected using the mean-shift algorithm [fukunaga1975estimation, cheng1995mean, comaniciu2002mean, carreira2000mode, carreira2003number, carreira2007gaussian, rodriguez2014clustering, carreira2015review], which can essentially be seen as iterative local averaging.
 The main idea in mean-shift clustering is to perform a mean-shift iteration starting from each data point and then define each mode as a cluster, with all points converging to the same mode grouped into the same cluster.
-The analysis of this algorithm has led to upper bounds on the number of modes of (1.1) [CPW03].
+The analysis of this algorithm has led to upper bounds on the number of modes of (1.1) [carreira2003number].
 Figure 3: Metastability of self-attention dynamics at temperature $\beta=81$ initialized with $n$ iid uniform points on the circle, with $n=200$ (top) and $n=1000$ (bottom).
@@
 (Code available at github.com/borjanG/2023-transformers-rotf.)
-We were instead brought to this problem from another perspective, motivated by the study of self-attention dynamics [SABP22, GLPR23, GLPR24, GRRB24]—a toy model for Transformers, the deep neural network architecture that has driven the success of large language models [VSP+17].
+We were instead brought to this problem from another perspective, motivated by the study of self-attention dynamics [sander2022sinkformers, geshkovski2023mathematical, geshkovski2024emergence, geshkovski2024measure]—a toy model for Transformers, the deep neural network architecture that has driven the success of large language models [vaswani2017attention].
 These dynamics form a mean-field interacting particle system
@@
 One sees that all particles move in time by following the field $\nabla\log(\mathsf{K}_{\beta^{-1/2}}\ast\mu_{\tau})$; here, $\mu_{\tau}$ is the empirical measure of the particles $x_{1}(\tau),\ldots,x_{n}(\tau)$ at time $\tau$.
-It is shown that for almost every initial configuration $x_{1}(0),\ldots,x_{n}(0)$, and for $\beta\geqslant 0$ in dimension $\geqslant 3$, or $\beta\leqslant 1\vee\beta\gtrsim n^{2}$ in dimension $2$, all particles converge to a single cluster in infinite time [GLPR23].
-Rather than converging quickly, the authors in [GKPR24] prove that the dynamics instead manifest metastability: particles quickly approach a few clusters, remain in the vicinity of these clusters for a very long period, and eventually coalesce into a single cluster in infinite time.
-Concurrently, and using different methods, the authors in [BPA24] show a similar result: starting from a perturbation of the uniform distribution, beyond a certain time, the empirical measure of the $n$ particles approaches an empirical measure of $O(\sqrt{\beta})$-equidistributed points on the circle in the mean-field limit, and stays near it for long time.
-This is done by a study of the linearized system and leveraging nonlinear stability results from [Gre00].
-Our interest lies in counting the number of clusters during the first metastable phase in dimension $d=2$.
-At time $\tau=0$, particles are initialized at $n$ iid points from the uniform distribution on the circle.
-In turn, the stationary points of $\mathsf{K}_{\beta^{-1/2}}\ast\mu_{0}$ partition the circle into intervals, with points clustering within their respective interval.
-This highlights the importance of counting the number of stationary points.
+It is shown that for almost every initial configuration $x_{1}(0),\ldots,x_{n}(0)$, and for $\beta\geqslant 0$, all particles converge to a single cluster in infinite time [geshkovski2023mathematical, criscitiello2024synchronization, polyanskiy2025synchronization].
+Rather than converging quickly, [geshkovski2024dynamic] prove that the dynamics instead manifest metastability: particles quickly approach a few clusters, remain in the vicinity of these clusters for a very long period, and eventually coalesce into a single cluster in infinite time.
+Concurrently, and using different methods, [bruno2024emergence] show a similar result: starting from a perturbation of the uniform distribution, beyond a certain time, the empirical measure of the $n$ particles approaches an empirical measure of $O(\sqrt{\beta})$-equidistributed points on the circle in the mean-field limit, and stays near it for long time.
+This is done by a study of the linearized system and leveraging nonlinear stability results from [grenier2000nonlinear].
+See also [bruno2025multiscale, karagodin2024clustering, alcalde2025attention].
+Our interest lies in the number of clusters during the first metastable phase in dimension $d=2$.
+At time $\tau=0$, we initialize $n$ tokens at iid uniform points on the circle.
+Under the self-attention flow, tokens follow the vector field $\nabla\log\big(K_{\beta^{-1/2}}*\mu_{\tau}\big)$, so metastable clusters coincide with local maxima of the smoothed empirical measure $K_{\beta^{-1/2}}*\mu_{\tau}$.
+In particular, at early times the circle is partitioned by the stationary points of $K_{\beta^{-1/2}}*\mu_{0}$, and each arc contracts toward its nearest maximum, making the number of clusters equal to the number of these maxima.
+Our $1d$ analysis shows that, for iid Gaussian data, the maxima concentrate in two symmetric belts where $t^{2}\in[2\log n-3\log\beta,2\log n-\log\beta]$; on the circle this corresponds to angular locations where the local inter-token spacing is $\simeq\beta^{-\frac{1}{2}}$, explaining both the $\sqrt{\beta}$ scaling of the metastable cluster count and their preferred positions.
 Here, we focus on a simplified setting by working on the real line instead of the circle (or higher-dimensional spheres), but we believe the analysis could be extended to these cases pending technical adaptations.
-Notwithstanding, Theorem 1.2 reflects what is seen in simulations (Figure 3).
+Notwithstanding, Theorem˜1.2 reflects what is seen in simulations (Figure˜3).
 1.3 Sketch of the proof
-The spirit of the proof of results such as Theorem 1.1 and others presented in [MMF92, Mam95, KM97] is similar to ours—one applies the Kac-Rice formula (Theorem 2.1) to a Gaussian approximation of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{{}^{\prime\prime}}(t))$ and argues its validity.
-However, the main limitation of these works is that modes are counted in a fixed and finite interval $[a,b]$ (and $[0,1]^{d}$ in the higher dimensional cases).
+The spirit of the proof of results such as Theorem˜1.1 and others presented in [mammen91, mammen95, mammen97] is similar to ours—one applies the Kac-Rice formula (Theorem˜2.1) to a Gaussian approximation of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{{}^{\prime\prime}}(t))$ and argues its validity.
+However, the main limitation of these works is that modes are counted in a fixed and finite interval $[a,b]$ (and $0,1]^{d}$ in the higher dimensional cases).
 Extending these techniques to the whole real line demands for different, significantly stronger, approximation results using Edgeworth expansions.
-We sketch the key ideas that allow us to count modes over $\mathbb{R}$.
-We use the Kac-Rice formula to compute the expected number of modes of $\widehat{P}_{n}$ in the symmetric interval $T$ and $T^{\prime}$ defined in 1.2 and 1.3.
+We sketch the key ideas that allow us to count modes over the real line. We truncate $\mathbb{R}$ to the interval
+$T\coloneqq\mathopen{}\mathclose{{\left[-\sqrt{2\log n-\log\beta-\omega(\beta)},\sqrt{2\log n-\log\beta-\omega(\beta)}}}\right]$ (1.2)
+where $\omega$ is a fixed, slow growing function such that
+$1\ll\omega(\beta)\ll\log\log\beta,$
+and so $T$ is well-defined for large $\beta$. Motivated by Theorems˜1.1 and 1.2, we also define the interval
+$T^{\prime}\coloneqq\mathopen{}\mathclose{{\left[-\sqrt{2\log n-3\log\beta},\sqrt{2\log n-3\log\beta}}}\right]$ (1.3)
+if $\beta\leqslant n^{\frac{2}{3}}$ and define $T^{\prime}=\varnothing$ if $\beta>n^{\frac{2}{3}}$.
+We use the Kac-Rice formula to compute the expected number of modes of $\widehat{P}_{n}$ in the symmetric intervals $T$ and $T^{\prime}$.
 All asymptotics are as $n,\beta\to\infty$.
@@
   2. 2.
-    In expectation over $X_{i}$, the number of modes of $\widehat{P}_{n}$ in $T^{\prime}$ is $O\mathopen{}\mathclose{{}\left(e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}}\right)$.
+    In expectation over $X_{i}$, the number of modes of $\widehat{P}_{n}$ in $T^{\prime}$ is $O\mathopen{}\mathclose{{\left(\sqrt{\beta}}}\right)$.
 The Kac-Rice computation appears tractable only when the joint distribution of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{\prime\prime}(t))$ is Gaussian, which it is not.
-To overcome this obstacle, we apply the Kac-Rice formula over a Gaussian approximation of the joint distribution in Section 2.
-For the specific underlying density and KDE in 1.1, we are able to justify in Section 3 the approximation for all $t$ in the growing interval $T$ instead of a fixed interval.
-This is why Theorem 1.1 only counts modes in a fixed interval.
+To overcome this obstacle, we apply the Kac-Rice formula over a Gaussian approximation of the joint distribution in Section˜2.
+For the specific underlying density and KDE in ˜1.1, we are able to justify in Section˜3 the approximation for all $t$ in the growing interval $T$ instead of a fixed interval.
+This is why Theorem˜1.1 only counts modes in a fixed interval.
 To show the validity of the Gaussian approximation, we use the Edgeworth expansion of the joint distribution of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{\prime\prime}(t))$ around the Gaussian distribution with matching first two moments.
-We bound the error due to the third order term of the expansion directly, and deal with the higher order terms by appealing to the error bounds of densities in the Edgeworth approximation similar to [BR10, Theorem 19.2].
-This strategy has been used in [BCP19], in a different context.
-We note that [KM97] employ the same theorem to justify the Gaussian process approximation over a fixed interval.
-In doing so, we will see that the normal approximation is invalid outside of $T$ (see Remark 3.3), but crucially $T$ is sufficiently large to cover almost all modes, as observed empirically in Remark 1.5 and Figure 4 and given below.
+We bound the error due to the third order term of the expansion directly, and deal with the higher order terms by appealing to the error bounds of densities in the Edgeworth approximation similar to [bhattacharya2010normal, Theorem 19.2 and 19.3].
+This strategy has been used in [bally2019non], but in a completely different context.
+We note that [mammen97] employ the same theorem to justify the Gaussian process approximation over a fixed interval.
+Indeed, as $T,\beta$ grow with $n$, the error decay rate of the Edgeworth approximation is quite delicate near the boundary of $T$.
+Instead of the usual case of powers of $n^{-\frac{1}{2}}$, it is powers of $e^{-\frac{\omega(\beta)}{4}}$ (see Lemma˜3.5 and ˜3.6).
+This is exactly why we need to introduce the $\omega(\beta)$ term in $T$.
+In doing so, we will see that the normal approximation is invalid outside of $T$ (see Remark˜3.4), but crucially $T$ is sufficiently large to cover almost all modes, as observed empirically in Remark˜1.4 and Figure˜4 and given below.
 Proposition 1.8.
-If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for arbitrarily small $c>0$, then the expectation over $X_{i}$ of the number of modes of $\widehat{P}_{n}$ that lie outside of $T$ is $O\mathopen{}\mathclose{{}\left(\sqrt{\beta\exp(\omega(\beta))}}\right)$.
-We prove this in Section 4.2 with an argument from scale-space theory: we bound the number of modes outside $T$ by the number of samples $X_{i}$ outside $T$, which we then bound naively.
-This is precisely the argument used by [CPW03, Theorem 2] to show Gaussian mixtures over $\mathbb{R}$ with $n$ components must have at most $n$ modes.
-This argument crucially relies on the kernel density estimator being Gaussian (see Remark 4.4).
-Now, Theorem 1.2 follows from Propositions 1.6 and 1.7 upon recalling that $1\ll\omega(\beta)\ll\log\log\beta$.
+If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for arbitrarily small $c>0$, then the expectation over $X_{i}$ of the number of modes of $\widehat{P}_{n}$ that lie outside of $T$ is $O\mathopen{}\mathclose{{\left(e^{\frac{\omega(\beta)}{2}}\sqrt{\beta}}}\right)$.
+We prove this in Section˜4.2 with an argument from scale-space theory: we bound the number of modes outside $T$ by the number of samples $X_{i}$ outside $T$, which we then bound naively.
+This is precisely the argument used by [carreira2003number, Theorem 2] to show Gaussian mixtures over $\mathbb{R}$ with $n$ components must have at most $n$ modes.
+This argument crucially relies on the kernel density estimator being Gaussian (see Remark˜4.4).
+Now, Theorem˜1.2 follows from Propositions˜1.7 and 1.8 provided $1\ll\omega(\beta)\ll\log\log\beta$.
+Indeed, the bounds on $\omega(\beta)$ are chosen to balance these error terms.
+In fact, all error terms other than the Kac-Rice integral over $T\setminus T^{\prime}$ of the Gaussian approximation of the density are $O\mathopen{}\mathclose{{\left(e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}}}\right)$.
 1.4 Notation
 We adopt standard notation from asymptotic analysis: we write $f(x)\ll g(x)$ or $f(x)=o(g(x))$ if $f(x)/g(x)\to 0$ as $x\to\infty$; $f(x)\lesssim g(x)$ or $f(x)=O(g(x))$ if there exists a finite, positive constant $C$ such that $f(x)\leqslant Cg(x)$; and we write $f(x)\asymp g(x)$ or $f(x)=\Theta(g(x))$ if $f(x)\lesssim g(x)$ and $g(x)\lesssim f(x)$.
+We also write $f(x)\sim g(x)$ if $f(x)/g(x)\to 1$ as $x\to\infty$.
+Similarly, for vector and matrix-valued functions $\mathbf{f}(x)\lesssim\mathbf{g}(x)$ if $f_{i}(x)\lesssim g_{i}(x)$ for every entry, indexed by $i$.
+We use the analogous notation for $\mathbf{f}(x)\asymp\mathbf{g}(x)$ and $\mathbf{f}(x)\sim\mathbf{g}(x)$.
+Note that all asymptotic constants are absolute.
 2 Kac-Rice for the normal approximation
@@
 The Kac-Rice formula allows us to compute the expected number of up-crossings when $F$ is a random field (i.e., a stochastic process).
-Theorem 2.1 (Kac-Rice, [AW09, pp. 62], [AT09, Section 11.1]).
+Theorem 2.1 (Kac-Rice, [azais2009level, pp. 62], [adler2009random, Section 11.1]).
 Consider a random $\Psi:\mathbb{R}\to\mathbb{R}$, some fixed $u\in\mathbb{R}$ and a compact $T\subset\mathbb{R}$. Suppose
@@
   4. 4.
-    $\mathbb{P}(\upomega(\eta)>\varepsilon)=O(\eta)$ as $\eta\searrow 0^{+}$ for any $\varepsilon>0$, where $\upomega(\cdot)$ denotes the modulus of continuity111defined, for $f:\mathbb{R}\to\mathbb{R}$, as $\upomega(\eta)=\sup_{t,s\colon|t-s|\leqslant\eta}|f(t)-f(s)|$.
+    $\mathbb{P}(\upomega(\eta)>\varepsilon)=O(\eta)$ as $\eta\searrow 0^{+}$ for any $\varepsilon>0$, where $\upomega(\cdot)$ denotes the modulus of continuity222defined, for $f:\mathbb{R}\to\mathbb{R}$, as $\upomega(\eta)=\sup_{t,s\colon|t-s|\leqslant\eta}|f(t)-f(s)|$.
 of $\Psi^{\prime}(\cdot)$.
@@
 $\mathbb{E}U_{u}(\Psi,T)=\int_{T}\int_{0}^{\infty}yp_{t}(u,y)\,\mathrm{d}y\,\mathrm{d}t.$ (2.1)
-The Kac-Rice formula extends to any dimension $d\geqslant 1$, and also on manifolds other than $\mathbb{R}^{d}$—see [AT09, Section 11.1].
-It is the classical tool for computing the expected number of critical points of random fields, with many recent applications including spin glasses [AAČ13, FMM21] and landscapes of loss functions arising in machine learning [MAB20].
+The Kac-Rice formula extends to any dimension $d\geqslant 1$, and also on manifolds other than $\mathbb{R}^{d}$—see [adler2009random, Section 11.1].
+It is the classical tool for computing the expected number of critical points of random fields, with many recent applications including spin glasses [auffinger2013random, fan2021tap] and landscapes of loss functions arising in machine learning [maillard2020landscape].
 While the method applies to general densities, the conditional expectation appears infeasible to compute or estimate beyond the Gaussian case.
-For the KDE $\widehat{P}_{n}$ defined in 1.1, define the random function $F_{n}:\mathbb{R}\to\mathbb{R}$ by
+Moreover, we remark that our reliance on the Kac-Rice formula precludes us from any “with high probability” anlogs of Theorem˜1.2, though we do expect such statements to hold.
+For the KDE $\widehat{P}_{n}$ defined in ˜1.1, define the random function $F_{n}:\mathbb{R}\to\mathbb{R}$ by
 $F_{n}(t)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathopen{}\mathclose{{\left(t-X_{i}}}\right)e^{-\frac{\beta}{2}\mathopen{}\mathclose{{\left(t-X_{i}}}\right)^{2}}=-\sqrt{\frac{2\pi n}{\beta^{3}}}\widehat{P}_{n}^{\prime}(t).$ (2.2)
-Then $t\in\mathbb{R}$ is an upcrossing of $F_{n}$ at level 00 if and only if $F_{n}(t)=0$ and $F^{\prime}_{n}(t)>0$.
+Then $t\in\mathbb{R}$ is an upcrossing of $F_{n}$ at level 0 if and only if $F_{n}(t)=0$ and $F^{\prime}_{n}(t)>0$.
 This is equivalent to $\widehat{P}^{\prime}_{n}(t)=0$ and $\widehat{P}_{n}^{\prime\prime}(t)<0$, i.e.
@@
 Thus, the number of modes of $\widehat{P}_{n}$ in $T$ is given by $U_{0}(F_{n},T)$.
-For $T,T^{\prime}$ defined in (1.2)–(1.3), Propositions 1.6 and 1.7 yield
+For $T,T^{\prime}$ defined in (1.2)–(1.3), Propositions˜1.7 and 1.8 are equivalent to
 $\mathbb{E}U_{0}\mathopen{}\mathclose{{\left(F_{n},T}}\right)$ $\asymp\sqrt{\beta\log\beta},$ (2.3)
@@
 2.2 Computing the Gaussian approximation
-Without loss of generality, fix $t\in T$ with $t\geqslant 0$. We can rewrite $F_{n}(t)$ from 2.2 and compute its derivative: for independent copies $(G_{i},G_{i}^{\prime})$ of
+Without loss of generality, fix $t\in T$ with $t\geqslant 0$. We can rewrite $F_{n}(t)$ from ˜2.2 and compute its derivative: for independent copies $(G_{i},G_{i}^{\prime})$ of
 $\begin{bmatrix}G(t)\\ G^{\prime}(t)\end{bmatrix}=e^{-\frac{\beta}{2}(t-X)^{2}}\begin{bmatrix}t-X\\ 1-\beta(t-X)^{2}\end{bmatrix}$, (2.4)
@@
 $\begin{bmatrix}F_{n}(t)\\ F_{n}^{\prime}(t)\end{bmatrix}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\begin{bmatrix}G_{i}(t)\\ G_{i}^{\prime}(t)\end{bmatrix}\sim p_{t}.$
-We prove that $p_{t}$ is a well-defined density in Proposition 4.1, and defer the following computation to Sections A.1 and A.1.
-Fact 2.2.
+We prove that $p_{t}$ is a well-defined density in Proposition˜4.1, and defer the following computation to Sections˜5.2 and 5.2.
+Lemma 2.2.
 The mean and covariance matrix of the random vector $(F_{n}(t),F_{n}^{\prime}(t))$ are given respectively by
-$\mu_{t}$ $\coloneqq\sqrt{n}\begin{bmatrix}\mathbb{E}G(t)\\ \mathbb{E}G^{\prime}(t)\end{bmatrix}\asymp n^{\frac{1}{2}}\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}t\\ -t^{2}\end{bmatrix}$ (2.5)
-$\Sigma_{t}$ $\coloneqq\begin{bmatrix}\operatorname{Var}G(t)&\operatorname{Cov}(G(t),G^{\prime}(t))\\ \operatorname{Cov}(G(t),G^{\prime}(t))&\operatorname{Var}G^{\prime}(t)\end{bmatrix}\asymp\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}1&-t\\ -t&\beta\end{bmatrix}.$
+$\mu_{t}$ $\coloneqq\sqrt{n}\begin{bmatrix}\mathbb{E}G(t)\\ \mathbb{E}G^{\prime}(t)\end{bmatrix}\sim n^{\frac{1}{2}}\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}t\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1}-t^{2}\end{bmatrix}$ (2.5)
+$\Sigma_{t}$ $\coloneqq\begin{bmatrix}\operatorname{Var}G(t)&\operatorname{Cov}(G(t),G^{\prime}(t))\\ \operatorname{Cov}(G(t),G^{\prime}(t))&\operatorname{Var}G^{\prime}(t)\end{bmatrix}\sim 2^{-\frac{5}{2}}\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}2&-t\\ -t&3\beta\end{bmatrix}$.
 We proceed to centering and rescaling the density $p_{t}$. Let $Y_{i}(t)$, $i\in[n]$, be independent copies of
 $Y(t)=\Sigma_{t}^{-\frac{1}{2}}\begin{bmatrix}G(t)-\mathbb{E}G(t)\\ G^{\prime}(t)-\mathbb{E}G^{\prime}(t)\end{bmatrix}$ (2.6)
-Let $q_{t}$ denote the density of $\sqrt{n}\sum_{i=1}^{n}Y_{i}(t)$. By construction $q_{t}$ has mean 00 and covariance $I_{2}$. Moreover, by the change-of-variables formula, it holds
+Let $q_{t}$ denote the density of $n^{-\frac{1}{2}}\sum_{i=1}^{n}Y_{i}(t)$. By construction $q_{t}$ has mean 0 and covariance $I_{2}$. Moreover, by the change-of-variables formula, it holds
 $p_{t}(x,y)=\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}q_{t}\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(x,y)-\mu_{t}]}}\right).$ (2.7)
 Now, let $\varphi:\mathbb{R}^{2}\to\mathbb{R}$ be the density of $N(0,I_{2})$, i.e.,
-$\varphi(x)\coloneqq\frac{1}{\sqrt{2\pi}}e^{-\frac{\|x\|^{2}}{2}}.$
-We aim to approximate the Kac-Rice integral 2.1 as follows:
-$\int_{T}\int_{0}^{\infty}yp_{t}(0,y)\,\mathrm{d}y\approx\int_{T}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t.$ (2.8)
-The validity of this approximation is deferred to Section 3. In the remainder of this section, we solely focus on computing the right hand side integral.
+$\varphi(x)\coloneqq\frac{1}{{2\pi}}e^{-\frac{\|x\|^{2}}{2}}.$
+We aim to approximate the Kac-Rice integral ˜2.1 as follows:
+$\int_{T}\int_{0}^{\infty}yp_{t}(0,y)\,\mathrm{d}y{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,\mathrm{d}t}\approx\int_{T}\int_{0}^{\infty}y\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\varphi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t.$ (2.8)
+The validity of this approximation is deferred to Section˜3. In the remainder of this section, we solely focus on computing the right hand side integral.
 Lemma 2.3.
-There exists some $A_{t}\asymp\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}$ such that
-$\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)$ $\asymp\exp\mathopen{}\mathclose{{}\left(-A_{t}-\Theta\mathopen{}\mathclose{{}\left(\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}}\right)y^{2}}\right),$ (2.9)
-$\int_{0}^{\infty}y\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y$ $\asymp\beta^{-\frac{1}{2}}e^{-\frac{t^{2}}{2}}e^{-A_{t}}.$
-Proof of Lemma 2.3.
-Since $\det\Sigma_{t}\asymp\beta^{-2}e^{-t^{2}}$, we have
-$\Omega\coloneqq\Sigma_{t}^{-1}\asymp\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}\begin{bmatrix}\beta&t\\ t&1\end{bmatrix}.$
-Now as $t\in T$ and so $t^{2}\ll\beta$, we find
+There exists $A_{t}\asymp\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}},\delta_{t}\asymp n^{\frac{1}{2}}\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}(1-t^{2}/2)$, and $\alpha_{t}\asymp\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}$ such that
+$\mathopen{}\mathclose{{\left\|\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right\|^{2}$ $\sim A_{t}+\alpha_{t}(y-\delta_{t})^{2},$ (2.9)
+$\int_{0}^{\infty}y\varphi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y$ $\asymp\alpha_{t}^{-1}e^{-A_{t}}.$
+Proof of Lemma˜2.3.
+We recall ˜2.5 to compute
+$\Omega\coloneqq\Sigma_{t}^{-1}\sim 3^{-1}2^{\frac{3}{2}}\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}\begin{bmatrix}3\beta&t\\ t&2\end{bmatrix}$
+We let the prefactor to be $\alpha_{t}/2$.
+Since we kept leading coefficients of entries of $\Omega$ and $\mu_{t}$ up to a global absolute constant that we absorb in $\alpha_{t}$ and $\delta_{t}$, it is safe to verify leading coefficients do not cancel and compute asymptotically:
 $\mathopen{}\mathclose{{\left\|\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right\|^{2}$ $=\mathopen{}\mathclose{{\left\langle(-\mu_{t,1},y-\mu_{t,2}),\Sigma_{t}^{-1}(-\mu_{t,1},y-\mu_{t,2})}}\right\rangle$
-$\asymp\Omega_{11}\mu_{t,1}^{2}-2\Omega_{12}\mu_{t,1}(y-\mu_{t,2})+\Omega_{22}(y-\mu_{t,2})^{2}$
-$\asymp\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}\mu_{t,1}^{2}\mathopen{}\mathclose{{}\left[\beta-2t\mathopen{}\mathclose{{}\left(\frac{y}{\mu_{t,1}}+t}\right)+\mathopen{}\mathclose{{}\left(\frac{y}{\mu_{t,1}}+t}\right)^{2}}\right]$
-$\asymp\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}\mu_{t,1}^{2}(\beta-t^{2})+\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}y^{2}$
-$\asymp\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}+\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}y^{2}.$
-This shows the first statement in 2.9. For the second, we have
-$\int_{0}^{\infty}y\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y$ $\asymp e^{-A_{t}}\int_{0}^{\infty}ye^{-\Theta\mathopen{}\mathclose{{}\left(\beta^{1/2}e^{t^{2}/{2}}}\right)y^{2}}\,\mathrm{d}y$
-$\asymp\beta^{-\frac{1}{2}}e^{-\frac{t^{2}}{2}}e^{-A_{t}}$
-by Gaussian integral computations (see A.1). ∎
+$=\Omega_{11}\mu_{t,1}^{2}-2\Omega_{12}\mu_{t,1}(y-\mu_{t,2})+\Omega_{22}(y-\mu_{t,2})^{2}$
+$\sim\alpha_{t}\mu_{t,1}^{2}\mathopen{}\mathclose{{\left[\frac{3\beta}{2}-t\mathopen{}\mathclose{{\left(\frac{y}{\mu_{t,1}}+\frac{t^{2}-1}{t}}}\right)+\mathopen{}\mathclose{{\left(\frac{y}{\mu_{t,1}}+\frac{t^{2}-1}{t}}}\right)^{2}}}\right]$
+$\sim\alpha_{t}\mu_{t,1}^{2}\mathopen{}\mathclose{{\left(\frac{3\beta}{2}-\frac{t^{2}}{4}}}\right)+\alpha_{t}\mu_{t,1}^{2}\mathopen{}\mathclose{{\left(\frac{y}{\mu_{t,1}}+\frac{t^{2}-1}{t}-\frac{t}{2}}}\right)^{2}$
+$\sim\frac{3}{2}\beta\alpha_{t}\mu_{t,1}^{2}+\alpha_{t}\mathopen{}\mathclose{{\left(y-\frac{\mu_{t,1}}{t}\mathopen{}\mathclose{{\left(1-\frac{t^{2}}{2}}}\right)}}\right)^{2}$
+Now, we let $A_{t}$ be the first term and let $\delta_{t}$ be the term subtracting $y$. Verifying the asymptotics of both, we obtain the first statement in ˜2.9. For the second statement, we have
+$\int_{0}^{\infty}y\varphi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\sim e^{-A_{t}}\int_{0}^{\infty}ye^{-2\alpha_{t}(y-\delta_{t})^{2}}\,\mathrm{d}y\asymp\alpha_{t}^{-1}e^{-A_{t}}$
+by a standard fact (Lemma˜5.1) in Gaussian integrals, upon checking $\alpha_{t}^{-1}\delta_{t}^{2}\ll 1$ in our parameter regime of $t\in T$ and $n\lesssim\beta^{2-c}$.
+The statement and proof of the fact is in Section˜5.1.
+∎
 2.3 The Kac-Rice integral over $\varphi$
@@
 (Code available at github.com/KimiSun18/2024-gauss-kde-attention.)
-We compute 2.1 under the approximation 2.8. By 2.9 and 2.5, we have that
+We compute ˜2.1 under the approximation ˜2.8. By ˜2.9 and ˜2.5, we have that
 $\int_{S}\int_{0}^{\infty}y\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\varphi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t\asymp\sqrt{\beta}\int_{S}e^{-A_{t}}\,\mathrm{d}t$ (2.10)
 for any measurable $S\subset\mathbb{R}$.
-Assuming validity of the Gaussian approximation (see Section 3), it follows from the Kac-Rice formula that the density of modes at $t\in\mathbb{R}$ is proportional to $\sqrt{\beta}e^{-A_{t}}$.
-We plot this density in Figure 4 with the same choice of $n$ and $\beta$ as in the empirical distribution.
-We see that they match on the highlighted interval $T$, but not outside of $T$ where the Gaussian approximation—see Remark 3.3.
-We compute 2.10 explicitly for $S=T$ and $S=T^{\prime}$.
+Assuming validity of the Gaussian approximation (see Section˜3), it follows from the Kac-Rice formula that the density of modes at $t\in\mathbb{R}$ is proportional to $\sqrt{\beta}e^{-A_{t}}$.
+We plot this density in Figure˜4 with the same choice of $n$ and $\beta$ as in the empirical distribution.
+We see that they match on the highlighted interval $T$, but not outside of $T$ where the Gaussian approximation is no longer valid—see Remark˜3.4.
+We compute ˜2.10 explicitly for $S=T$ and $S=T^{\prime}$.
 Lemma 2.4.
@@
 $\int_{T^{\prime}}\int_{0}^{\infty}y\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\varphi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$ $\lesssim\sqrt{\beta}.$
-Proof of Lemma 2.4.
-Recall $A_{t}$ from Lemma 2.3. By 2.10, it suffices to show that
+Proof of Lemma˜2.4.
+Recall $A_{t}$ from Lemma˜2.3. By ˜2.10, it suffices to show that
 $\int_{T}e^{-A_{t}}\,\mathrm{d}t\asymp\sqrt{\log\beta}\quad\text{and}\quad\int_{T^{\prime}}e^{-A_{t}}\,\mathrm{d}t\lesssim 1.$ (2.11)
-As $A_{t}>0$, the integral is at most the length of $T$, which is $O(\sqrt{\log\beta})$ by 1.2.
-Let $t_{s}\coloneqq\sqrt{2\log n-s\log\beta}$.
-As the integrand is positive, for constants $C,C^{\prime}>0$
-$\int_{T}e^{-A_{t}}\,\mathrm{d}t$ $\geqslant\int_{t_{2}}^{t_{5/2}}\exp\mathopen{}\mathclose{{}\left(-C\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}}\right)\,\mathrm{d}t$
-$\geqslant\mathopen{}\mathclose{{}\left(t_{5/2}-t_{2}}\right)\exp\mathopen{}\mathclose{{}\left(-C\beta^{-\frac{3}{2}}nt_{5/2}^{2}e^{-\frac{t_{2}^{2}}{2}}}\right)$
-$\gtrsim\sqrt{\log\beta}\exp\mathopen{}\mathclose{{}\left(-C^{\prime}\beta^{-\frac{1}{2}}\log n}\right)$
+As $A_{t}>0$, the integral is at most the length of $T$, which is $O(\sqrt{\log\beta})$ by ˜1.2. Recall that $n\lesssim\beta^{2-c}$. For this constant $c>0$ and $k=1,2$, define
+$t_{k}\coloneqq\sqrt{2\log n-\mathopen{}\mathclose{{\left(1+\frac{c}{10^{k}}}}\right)\log\beta}$
+Then, $t_{k}>0$ and $t_{k}\in T$ for both $k$. Now, as the integrand is positive, for constants $C,C^{\prime}>0$
+$\int_{T}e^{-A_{t}}\,\mathrm{d}t$ $\geqslant\int_{t_{1}}^{t_{2}}\exp\mathopen{}\mathclose{{\left(-C\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}}}\right)\,\mathrm{d}t$
+$\geqslant\mathopen{}\mathclose{{\left(t_{2}-t_{1}}}\right)\exp\mathopen{}\mathclose{{\left(-C\beta^{-\frac{3}{2}}nt_{2}^{2}e^{-\frac{t_{1}^{2}}{2}}}}\right)$
+$\gtrsim\sqrt{\log\beta}\exp\mathopen{}\mathclose{{\left(-C^{\prime}\beta^{-\frac{1}{2}+\frac{c}{10}}\log n}}\right)$
 $\gtrsim\sqrt{\log\beta}$
-as $n,\beta\to\infty$ with $\log n\asymp\log\beta$. Now if $t\in T^{\prime}=[-t_{3},t_{3}]$, we have
-$e^{-\frac{t^{2}}{2}}\geqslant e^{-\frac{t_{3}^{2}}{2}}=\beta^{\frac{3}{2}}n^{-1}.$
+as $n,\beta\to\infty$ with $\log n\asymp\log\beta$, and $c\in(0,2]$. Now if $t\in T^{\prime}$, we have
+$e^{-\frac{t^{2}}{2}}\geqslant\exp\mathopen{}\mathclose{{\left(\log n-\frac{3}{2}\log\beta}}\right)=\beta^{\frac{3}{2}}n^{-1}.$
 Hence
-$\int_{T^{\prime}}e^{-A_{t}}\,\mathrm{d}t\lesssim\int_{0}^{t_{3}}\exp\mathopen{}\mathclose{{}\left(-C\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}}\right)\,\mathrm{d}t\leqslant\int_{0}^{t_{3}}e^{-Ct^{2}}\,\mathrm{d}t\lesssim 1.\qed$
+$\int_{T^{\prime}}e^{-A_{t}}\,\mathrm{d}t\lesssim\int_{T^{\prime}}\exp\mathopen{}\mathclose{{\left(-C\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}}}\right)\,\mathrm{d}t\leqslant\int_{-\infty}^{\infty}e^{-Ct^{2}}\,\mathrm{d}t\lesssim 1.\qed$
 3 Leveraging the Edgeworth expansion
-In this section, we show the approximation of $p_{t}$ by $\varphi$ is valid in $T$ by showing
+In this section, we show the approximation of $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q_{t}$} by $\varphi$ is valid in $T$ by showing
 $\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left|q_{t}-\varphi}}\right|\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t\ll\sqrt{\beta\log\beta}.$ (3.1)
-One natural idea is to use some asymptotic series to expand $p_{t}$ around $\varphi$, e.g.
-the Edgeworth expansion, to argue that $|p_{t}-\varphi|\ll\varphi$ in the sense of the integral over $y$.
-We discuss the two major obstacles we have to overcome in order to implement this approach.
-  • Firstly, all known results on validity of asymptotic series such as Edgeworth expansions treat densities $q_{t}$ and $\varphi$ that are independent of $n$.
+One natural idea is to use some asymptotic series to expand $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q_{t}$} around $\varphi$, e.g.
+the Edgeworth expansion, to argue that $|q_{t}-\varphi|\ll\varphi$ in the sense of the integral over $y$.
+To this end, we cite a standard result in normal approximation theory in the case of identity covariance matrix, that we will follow closely,
+Theorem 3.1 ([bhattacharya2010normal, Theorem 19.2 and 19.3]).
+Let $X_{n}$ be a sequence of i.i.d.
+random vectors in $\mathbb{R}^{k}$ with mean zero and identity covariance matrix.
+Suppose $\mathbb{E}\|X_{1}\|^{s+1}<\infty$ for some integer $s\geqslant 2$, then under suitable conditions, the density $q_{n}$ of $n^{-1/2}(X_{1}+\dots+X_{n})$ admits the asymptotic expansion
+$\sup_{\mathbf{x}\in\mathbb{R}^{k}}\mathopen{}\mathclose{{\left(1+\|\mathbf{x}\|^{s}}}\right)\mathopen{}\mathclose{{\left|q_{n}(\mathbf{x})-\sum_{j=0}^{s-2}n^{-\frac{j}{2}}Q_{j}(\mathbf{x})}}\right|\lesssim n^{-\frac{s-1}{2}}$
+as $n\to\infty$, where $Q_{j}$ is the $j$-th term of the Edgeworth expansion. In particular, $Q_{0}=\varphi$.
+It is tempting to directly apply this theorem with $s=2$ to control $|q_{t}-\varphi|$ in ˜3.1. We discuss the two major obstacles we have to overcome in order to implement this approach.
+  • First, Theorem˜3.1 and similar results on validity of asymptotic series such as Edgeworth expansions treat densities $q_{t}$ and $\varphi$ that are independent of $n$.
 As $\beta$ grows in $n$, we will need to re-derive these results and carefully track the dependence on $\beta$.
@@
 Fortunately, this will be satisfied precisely when $t\in T$.
-  • Secondly, even without the $n$-dependence of $\beta$, expanding $q_{t}$ to the order $\varphi$ plus error in for example [BR10, Theorem 19.2] gives error roughly
-[…]
-    but then the integral over $y$ in 3.1 fails to converge.
-Therefore, we will need to go to the next term $\psi$ in the Edgeworth series.
-We control its integral over $y$ and $t$ in 3.1 manually in Section 3.1.
-Then, we control the higher order terms in Section 3.2 using the approach motivated above, so our error now converges upon integrating over $y$, i.e.
-roughly
-    $\mathopen{}\mathclose{{}\left|p_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right|(\mathbf{x})\lesssim\frac{1}{1+\|\mathbf{x}\|^{3}}\hskip 28.45274pt\text{ for all }\mathbf{x}\in\mathbb{R}^{2}.$
-3.1 Bounding the third order error
-Recall $Y$ from 2.6.
-For a multi-index $\alpha\in\mathbb{Z}_{\geqslant 0}^{2}$, let $\kappa^{\alpha}_{t}$ be the cumulant of $Y$ indexed by $\alpha$, which depends on $t$.
-Let $H^{\alpha}_{t}$ denote the standard Hermite polynomials of with index $\alpha$.
-The next term in the Edgeworth series is $n^{-1/2}\psi$ with
-$\psi(\mathbf{x})\coloneqq\frac{\varphi(\mathbf{x})}{6}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}(\mathbf{x})\quad\text{where}\quad H^{\alpha}\coloneqq(-1)^{|\alpha|}\frac{\partial^{\alpha}\varphi}{\varphi}$ (3.2)
-If $|\alpha|=3$, $\kappa_{t}^{\alpha}$ are the third moments of $Y$, which we bound in Section A.1.
-Fact 3.1.
-Let $\eta_{3}$ be the largest third cumulant of $Y$. Then
-$\eta_{3}\coloneqq\max\mathopen{}\mathclose{{}\left\{\kappa_{t}^{(k,3-k)}:k\in\{0,1,2,3\}}\right\}\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}.$
-We bound $H^{(k,3-k)}\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-1/2}[(0,y)-\mu_{t}]}\right)$ by a polynomial in $\tilde{y}=\Theta(\beta^{1/2}e^{{t^{2}}/{4}}y)$.
-Now, by a similar method as Lemma 2.3, we obtain the following bound.
-It says that when we integrate the Edgeworth series $p_{t}=\varphi+n^{-\frac{1}{2}}\psi+\dots$ over $y$ and $t$, the contribution $\varphi$ dominates $n^{-\frac{1}{2}}\psi$, hinting at the validity of the approximation.
+The analog of Theorem˜3.1 in our setting for $s=2,3$ is given as Lemma˜3.5.
+  • Second, if we directly apply Theorem˜3.1 with $s=2$ to control $|q_{t}-\varphi|$, then the inner integral over $y\in[0,\infty)$ in ˜3.1 fails to converge as it is of the form
+    $\int_{0}^{\infty}\frac{\tilde{y}}{1+\tilde{y}^{2}}\,\mathrm{d}\tilde{y}\quad\text{\quad}\text{where}\quad\tilde{y}\coloneqq\alpha_{t}^{\frac{1}{2}}(y-\delta_{t})$ (3.2)
+To overcome this obstacle, we use the above definition of $\tilde{y}$ and Lemma˜2.3 to obtain that
+$\mathopen{}\mathclose{{\left\|\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right\|^{2}\sim A_{t}+\tilde{y}^{2}$
+This naturally suggests casework on which term has the dominant contribution: for $|\tilde{y}|\leqslant\sqrt{A_{t}}$, we follow Theorem˜3.1 for the $s=3$ case to bound the error of $q_{t}-\varphi$, whereby ˜3.2 integrated from $y=0$ up to $\tilde{y}=\sqrt{A_{t}}$ converges; for $\tilde{y}\geqslant\sqrt{A_{t}}$, we go to the next term $n^{-1/2}\psi$ in the Edgeworth series, and manually control this third order error in Section˜3.1.
+Finally, we control the higher order terms following Theorem˜3.1 for the $s=3$ case: there, the analog of ˜3.2 for controlling
+$\int_{0}^{\infty}y\mathopen{}\mathclose{{\left|q_{t}-\varphi-n^{-\frac{1}{2}}\psi}}\right|\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y$
+converges as $s=3$. This is done in Section˜3.2.
+3.1 Bounding the third order error for small $y$
+Recall $Y$ from ˜2.6. Let $\varphi$ denote the density of $N(0,I_{2})$ and
+$H^{\alpha}(x)\coloneqq(-1)^{|\alpha|}\,\varphi(x)^{-1}\partial^{\alpha}\varphi(x)$
+the standard multivariate Hermite polynomials for a multi-index $\alpha\in\mathbb{Z}_{\geqslant 0}^{2}$.
+Writing $q_{t}$ for the density of $n^{-\frac{1}{2}}\sum_{i=1}^{n}Y_{i}(t)$, the third‑order Edgeworth expansion is the multivariate Hermite expansion of $q_{t}/\varphi$:
+$\frac{q_{t}(\mathbf{x})}{\varphi(\mathbf{x})}=1+\frac{1}{\sqrt{n}}\sum_{|\alpha|=3}\frac{\kappa_{t}^{\alpha}}{\alpha!}H^{\alpha}(\mathbf{x})+r_{n,t}(\mathbf{x}),$
+so the next term is $n^{-1/2}\psi$ with
+$\psi(\mathbf{x})=\varphi(\mathbf{x})\sum_{k=0}^{3}\frac{\kappa_{t}^{(k,3-k)}}{k!(3-k)!}H^{(k,\,3-k)}(\mathbf{x}).$ (3.3)
+Here $\kappa_{t}^{\alpha}$ denotes the (order-$|\alpha|$) cumulant of the single‑sample vector $Y(t)$, i.e.
+the $\alpha$‑th mixed derivative at 0 of the cumulant generating function $\log\mathbb{E}e^{\langle u,Y(t)\rangle}$.
+For $|\alpha|=3$ and our normalization, one has the equivalent moment identity
+$\kappa_{t}^{\alpha}=\mathbb{E}\mathopen{}\mathclose{{\left[H^{\alpha}(Y(t))}}\right]=\mathbb{E}_{Z\sim N(0,I_{2})}\mathopen{}\mathclose{{\left[\frac{q_{t}(Z)}{\varphi(Z)}H^{\alpha}(Z)}}\right].$
+If $|\alpha|\coloneqq\alpha_{1}+\alpha_{2}=s$, then $\kappa_{t}^{\alpha}$ can be bounded above asymptotically by the $s$-th moments of $\|Y\|$, which we bound in Section˜5.2.
 Lemma 3.2.
-Recall $T$ from 1.2 and $A_{t}$ from Lemma 2.3. Then
-$\int_{T}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left(n\det\Sigma_{t}}\right)^{-\frac{1}{2}}\psi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}.$
-Proof of Lemma 3.2.
-By the proof of Lemma 2.3, 3.2, and 3.1
-$\int_{T}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left(n\det\Sigma_{t}}\right)^{-\frac{1}{2}}\psi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$
-$=\frac{1}{6}\sum_{k=0}^{3}\int_{T}\mathopen{}\mathclose{{}\left(n\det\Sigma_{t}}\right)^{-\frac{1}{2}}\kappa_{t}^{(k,3-k)}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left[\varphi H^{(k,3-k)}}\right]\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$
-$\asymp\int_{T}\mathopen{}\mathclose{{}\left(n\det\Sigma_{t}}\right)^{-\frac{1}{2}}\eta_{3}e^{-A_{t}}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left(\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}y}\right)^{O(1)}e^{-\Theta\mathopen{}\mathclose{{}\left(\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}}\right)y^{2}}\,\mathrm{d}y\,\mathrm{d}t$
-[…]
+For $s\geqslant 3$, the cumulants of $Y$ with order $|\alpha|=s$ satisfy
+$\kappa_{t}^{\alpha}\lesssim\eta_{s}\quad\text{where}\quad\eta_{s}\coloneqq\mathbb{E}[\|Y\|^{s}]\lesssim\mathopen{}\mathclose{{\left(\beta e^{t^{2}}}}\right)^{\frac{s-2}{4}}.$
+Trivially, $|H^{(k,3-k)}(\mathbf{x})|\lesssim\|\mathbf{x}\|^{3}$. By Lemma˜2.3, we know for $y\geqslant\Delta_{t}\coloneqq{\delta_{t}+\sqrt{A_{t}/\alpha_{t}}}$ that
+$\tilde{y}\coloneqq\alpha_{t}^{\frac{1}{2}}(y-\delta_{t})\asymp\mathopen{}\mathclose{{\left\|\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right\|$
+Therefore, for any $t\in T$ and $y\geqslant\Delta_{t}$, we can bound
+$\mathopen{}\mathclose{{\left|H^{(k,3-k)}\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)}}\right|\lesssim\|\tilde{y}\|^{3}.$ (3.4)
+Now, by a similar method as Lemma˜2.3, we obtain the following bound.
+It says that when we integrate the Edgeworth series ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q_{t}}=\varphi+n^{-\frac{1}{2}}\psi+\dots$ over $y\geqslant\Delta_{t}$ and $t\in T$, the contribution $\varphi$ dominates $n^{-\frac{1}{2}}\psi$, hinting at the validity of the approximation.
+Lemma $3.$
+Recall $T,T^{\prime}$ from ˜1.2 and 1.3, and $A_{t}$ from Lemma˜2.3. Let $\Delta_{t}\coloneqq\delta_{t}+\sqrt{A_{t}/\alpha_{t}}$. Then
+$\int_{T}\int_{\Delta_{t}}^{\infty}y\mathopen{}\mathclose{{\left(n\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\psi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$ $\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta},$
+$\int_{T^{\prime}}\int_{\Delta_{t}}^{\infty}y\mathopen{}\mathclose{{\left(n\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\psi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$ $\lesssim\sqrt{\beta}.$
+Proof of Lemma˜3.3.
+Note that $y\geqslant\Delta_{t}$ if and only if $\tilde{y}\coloneqq\alpha_{t}^{\frac{1}{2}}(y-\delta_{t})\geqslant\sqrt{A_{t}}$. By Lemmas˜2.3 and 3.2, we can combine bounds ˜3.3 and 3.4 to obtain
+$\int_{T}\int_{\Delta_{t}}^{\infty}y\mathopen{}\mathclose{{\left(n\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\psi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$
+$\lesssim\sum_{k=0}^{3}\int_{T}\mathopen{}\mathclose{{\left(n\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\kappa_{t}^{(k,3-k)}\int_{\Delta_{t}}^{\infty}y\mathopen{}\mathclose{{\left[\varphi H^{(k,3-k)}}}\right]\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$
+$\lesssim\int_{T}\mathopen{}\mathclose{{\left(n\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\eta_{3}e^{-A_{t}}\int_{\Delta_{t}}^{\infty}ye^{-\frac{\tilde{y}^{2}}{2}}|\tilde{y}|^{3}\,\mathrm{d}y\,\mathrm{d}t$
+$\asymp\int_{T}\mathopen{}\mathclose{{\left(n\det\Sigma_{t}}}\right)^{-\frac{1}{2}}e^{-A_{t}}\eta_{3}\alpha_{t}^{-1}\mathopen{}\mathclose{{\left(\int_{\sqrt{A_{t}}}^{\infty}\tilde{y}^{4}e^{-\frac{\tilde{y}^{2}}{2}}d\tilde{y}}}\right)\,\mathrm{d}t$
 $\lesssim n^{-\frac{1}{2}}\beta^{\frac{3}{4}}\sup_{t\in T}\mathopen{}\mathclose{{\left(e^{\frac{t^{2}}{4}}}}\right)\int_{T}e^{-A_{t}}\,\mathrm{d}t$
 $\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}$
-where the last step follows 2.11 and the definition 1.2 of $T$. ∎
+where we apply ˜2.11. The second statement for $T^{\prime}$ holds similarly by considering $\sup_{t\in T^{\prime}}\mathopen{}\mathclose{{\left(e^{\frac{t^{2}}{4}}}}\right)$. ∎
 Remark 3.4.
@@
 Hence, the decay is only a factor of $e^{-\omega(\beta)/4}$.
-For $t\not\in T$, even $t=\sqrt{2\log n-0.99\log\beta}$, the last inequality in Lemma 3.2 fails and we get a bound of polynomially larger that $\sqrt{\beta}$.
+For $t\not\in T$, even $t=\sqrt{2\log n-0.99\log\beta}$, the last inequality in Lemma˜3.3 fails and we get a bound of polynomially larger that $\sqrt{\beta}$.
 Then, as the third order error is asymptotically larger than the contribution of the Gaussian approximation, so the normal approximation is no longer valid.
-This can be seen by comparing the plots in Figure 4.
+This can be seen by comparing the plots in Figure˜4.
 3.2 Bounding higher order errors
-We follow the classical proof about the validity of the Edgeworth expansion as an asymptotic series to show bound the higher order pointwise error of density function as follows.
+We follow the classical proof of Theorem˜3.1 about the validity of the Edgeworth expansion as an asymptotic series to show bound the higher order pointwise error of density function as follows.
 Lemma 3.5.
-Assume that $p_{t}$ is bounded almost everywhere for any fixed $t\in T$. If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for some $c>0$, then for any $t\in T$ and $\mathbf{x}\in\mathbb{R}^{2}$, we have that
-$\mathopen{}\mathclose{{}\left(1+\|\mathbf{x}\|^{3}}\right)\mathopen{}\mathclose{{}\left|q_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right|(\mathbf{x})\lesssim e^{-\frac{\omega(\beta)}{4}}.$ (3.3)
-We defer the proof to Section A.3.
-Here, we comment on the differences with [BR10, Theorem 19.2] for $s=3$, which we follow in spirit but modify to allow $n$ dependence in $q_{t}$ in the form of $\beta$.
-There, it is shown that the left hand side is $o(n^{-\frac{1}{2}})$ provided third moments $\eta_{3}=O(1)$.
-With the bound 3.1 on $\eta_{3}$ in our case which is not constant, the error from the asymptotic series decays not in powers of $n^{-\frac{1}{2}}$ but powers of $O(n^{-1/2}e^{{t^{2}}/{4}}\beta^{1/4})=O(\exp(-\omega(\beta)/2))$.
+Suppose $n^{c}\lesssim\beta\lesssim n^{2-c}$ for some $c>0$. Let $g_{2}\coloneqq q_{t}-\varphi$ and $g_{3}\coloneqq q_{t}-\varphi-n^{-\frac{1}{2}}\psi$. Then
+$\sup_{\mathbf{x}\in\mathbb{R}^{2}}\mathopen{}\mathclose{{\left(1+\|\mathbf{x}\|^{s}}}\right)\mathopen{}\mathclose{{\left|q_{t}-\varphi}}\right|(\mathbf{x})\lesssim n^{-\frac{s-1}{2}}\eta_{s+1}$ (3.5)
+as $n,\beta\to\infty$, for both $s=2$ and $s=3$, where we recall $\eta_{s}$ from Lemma˜3.2.
+We defer the proof to Section˜5.4.
+Here, we comment on the differences with Theorem˜3.1: there, it is shown that order $s$ error is at most $n^{-\frac{s-1}{2}}$ provided $\eta_{s+1}$ is bounded.
+In our case, we pick up an extra $\eta_{s+1}$ factor as it is dependent on $\beta$.
+By Lemma˜3.2, this bound is at most
+$n^{-\frac{s-1}{2}}\eta_{s+1}\lesssim\mathopen{}\mathclose{{\left(n^{-1}\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}}}\right)^{\frac{s-1}{2}}\lesssim\begin{cases}e^{-\frac{(s-1)\omega(\beta)}{4}}&\textup{if }t\in T\\ \beta^{-(s-1)/2}&\textup{if }t\in T^{\prime}\end{cases}$ (3.6)
+by definition of $T$ and $T^{\prime}$.
+This is the key reason for the extra $\omega(\beta)$ term in $T$: we get a small by non-negligible decay rate that matches Lemma˜3.3.
+Using Lemma˜3.5, we control the Kac-Rice integrals.
 Corollary 3.6.
-If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for $c>0$, then asymptotically in $n,\beta\to\infty$
-$\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{}\left|q_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right|\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}.$
-Proof of Corollary 3.5.
-By Lemma 3.4 and computations in the proof of Lemma 2.3
+Let $\Delta_{t}\coloneqq\delta_{t}+\sqrt{A_{t}/\alpha_{t}}$. If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for $c>0$, then asymptotically in $n,\beta\to\infty$
+$\int_{T}\int_{0}^{\Delta_{t}}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left|q_{t}-\varphi}}\right|\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$ $\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}$
+$\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left|q_{t}-\varphi-n^{-\frac{1}{2}}\psi}}\right|\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$ $\lesssim e^{-\frac{\omega(\beta)}{2}}\sqrt{\beta\log\beta}.$
+Moreover, the left hand sides of both displays above with $T$ replaced by $T^{\prime}$ are $O(\sqrt{\beta})$.
+Proof of Corollary˜3.6.
+By Lemmas˜3.5 and 2.3, we let $\tilde{y}\coloneqq\alpha_{t}^{\frac{1}{2}}(y-\delta_{t})$ as before to obtain
+$\int_{T}\int_{0}^{\Delta_{t}}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left|q_{t}-\varphi}}\right|\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$
+$\lesssim e^{-\frac{\omega(\beta)}{4}}\int_{T}\int_{0}^{\Delta_{t}}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left(1+\mathopen{}\mathclose{{\left\|\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right\|^{2}}}\right)^{-1}\,\mathrm{d}y\,\mathrm{d}t$
+$\lesssim e^{-\frac{\omega(\beta)}{4}}\int_{T}\int_{0}^{\Delta_{t}}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\mathopen{}\mathclose{{\left(\frac{y}{1+A_{t}+\tilde{y}^{2}}}}\right)\,\mathrm{d}y\,\mathrm{d}t$
+$\asymp e^{-\frac{\omega(\beta)}{4}}\int_{T}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\alpha_{t}^{-1}\mathopen{}\mathclose{{\left(\int_{-\sqrt{A_{t}}}^{\sqrt{A_{t}}}\frac{|\tilde{y}|}{1+A_{t}+\tilde{y}^{2}}\,\mathrm{d}\tilde{y}}}\right)\,\mathrm{d}t$
+$\asymp e^{-\frac{\omega(\beta)}{4}}\int_{T}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\alpha_{t}^{-1}\log\mathopen{}\mathclose{{\left(\frac{1+2A_{t}}{1+A_{t}}}}\right)\,\mathrm{d}t$
+$\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta}\int_{T}\log(2)dt$
+$\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}$
+where we check that if $y=0$, then $\tilde{y}=-\alpha_{t}^{\frac{1}{2}}\delta_{t}$ and $\alpha_{t}^{\frac{1}{2}}\delta_{t}\ll A_{t}^{\frac{1}{2}}$, so we may symmetrize the integral over $\tilde{y}$ up to a constant factor.
+Now similarly for the second display
 $\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left|q_{t}-\varphi-n^{-\frac{1}{2}}\psi}}\right|\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$
-$\lesssim e^{-\frac{\omega(\beta)}{4}}\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{}\left(1+\mathopen{}\mathclose{{}\left\|\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right\|^{3}}\right)^{-1}\,\mathrm{d}y\,\mathrm{d}t$
-$\lesssim e^{-\frac{\omega(\beta)}{4}}\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{}\left(1+\Theta\mathopen{}\mathclose{{}\left(\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}+\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}y^{2}}\right)^{\frac{3}{2}}}\right)^{-1}\,\mathrm{d}y\,\mathrm{d}t$
-$\lesssim e^{-\frac{\omega(\beta)}{4}}\int_{T}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}\mathopen{}\mathclose{{}\left(\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}}\right)^{-2}\int_{0}^{\infty}\frac{\tilde{y}}{1+\tilde{y}^{3}}\,\mathrm{d}\tilde{y}$
-$\asymp e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}$
-for $\tilde{y}\asymp\beta^{1/4}e^{{t^{2}}/{4}}y$, where we note that $T$ has length $\Theta(\sqrt{\log\beta})$ by 1.2 and
-$\int_{0}^{\infty}\frac{\tilde{y}}{1+\tilde{y}^{3}}\,\mathrm{d}\tilde{y}=\frac{2\pi}{3\sqrt{3}}=O(1).\qed$
-4 Proof of Theorem 1.2
-We prove Propositions 1.6 and 1.7 by checking 2.3, thereby proving Theorem 1.2.
-4.1 Proof of Proposition 1.6
-To prove Proposition 1.6 we seek to apply Theorem 2.1 to $\mathbb{E}U_{0}(F_{n},T)$. This in turn requires checking all the assumptions of Theorem 2.1. We have
+$\lesssim e^{-\frac{\omega(\beta)}{2}}\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left(1+\mathopen{}\mathclose{{\left\|\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right\|^{3}}}\right)^{-1}\,\mathrm{d}y\,\mathrm{d}t$
+$\lesssim e^{-\frac{\omega(\beta)}{2}}\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left(1+\mathopen{}\mathclose{{\left(A_{t}+\tilde{y}^{2}}}\right)^{\frac{3}{2}}}}\right)^{-1}\,\mathrm{d}y\,\mathrm{d}t$
+$\lesssim e^{-\frac{\omega(\beta)}{2}}\int_{T}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\alpha_{t}^{-1}\mathopen{}\mathclose{{\left(\int_{-\infty}^{\infty}\frac{|\tilde{y}|}{1+|\tilde{y}|^{3}}\,\mathrm{d}\tilde{y}}}\right)\,\mathrm{d}t$
+$\asymp e^{-\frac{\omega(\beta)}{2}}\sqrt{\beta}\int_{T}\frac{2\pi}{3\sqrt{3}}dt$
+$\asymp e^{-\frac{\omega(\beta)}{2}}\sqrt{\beta\log\beta}$
+Now, the exact same computation but with $T$ replaced by $T^{\prime}$ gives bounds $O(\sqrt{\beta})$ upon replacing exponential in $\omega(\beta)$ decay rates with those in ˜3.6 for Lemma˜3.5.
+In particular, by non-negativity of the integrand, the second equation holds upon replacing the bounds of integration of $y$ from $y\geqslant 0$ to $y\geqslant\Delta_{t}$.
+This is the form we will use.
+4 Proof of Theorem˜1.2
+We prove Propositions˜1.7 and 1.8 by checking ˜2.3, thereby proving Theorem˜1.2.
+4.1 Proof of Proposition˜1.7
+To prove Proposition˜1.7 we seek to apply Theorem˜2.1 to $\mathbb{E}U_{0}(F_{n},T)$. This in turn requires checking all the assumptions of Theorem˜2.1. We have
 Proposition 4.1.
-Let $\beta,n$ be as in Theorem 1.2, and fix $t\in T$.
+Fix any $\beta>0$, an integer $n\geqslant 5$, and $t\in T$.
 Let $\upmu_{t}$ denote the law of $(F_{n}(t),F_{n}^{\prime}(t))$ defined in (2.2).
 Then $\upmu_{t}$ admits a density $p_{t}\in\mathscr{C}^{0}(\mathbb{R}^{2})$ satisfying $p_{t}({\bf{x}})\to 0$ as $\|\bf{x}\|\to\infty$.
-Moreover, conditions 1, 2, 4 in Theorem 2.1 also hold for $\Psi(t)=F_{n}(t)$, thus Theorem 2.1 applies.
-We defer the proof to Section A.4. With Proposition 4.1, we deduce
+Moreover, conditions 1, 2, 4 in Theorem˜2.1 also hold for $\Psi(t)=F_{n}(t)$, thus Theorem˜2.1 applies.
+We defer the proof to Section˜5.5. With Proposition˜4.1, we deduce
 Lemma 4.2.
-With the notation as in Theorem 2.1,
+With the notation as in Theorem˜2.1,
 $\mathbb{E}U_{0}\mathopen{}\mathclose{{\left(F_{n},T}}\right)=\int_{T}\int_{0}^{\infty}yp_{t}(0,y)\,\mathrm{d}y\,\mathrm{d}t.$ (4.1)
-Proof of Proposition 1.6.
-Combining Lemmas 2.4, 3.2 and 3.5 gives
+Proof of Proposition˜1.7.
+We decompose the Kac-Rice integral as follows:
 $\mathbb{E}U_{0}\mathopen{}\mathclose{{\left(F_{n},T}}\right)$ $=\int_{T}\int_{0}^{\infty}yp_{t}(0,y)\,\mathrm{d}y\,\mathrm{d}t$
@@
 $=\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\varphi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$
-$\quad+\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(n\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\psi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$
-$\quad+\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{}\left[q_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right]\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$
+$\quad+\int_{T}\int_{0}^{\Delta_{t}}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left[q_{t}-\varphi}}\right]\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$
+$\quad+\int_{T}\int_{\Delta_{t}}^{\infty}\mathopen{}\mathclose{{\left(n\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\psi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$
+$\quad+\int_{T}\int_{\Delta_{t}}^{\infty}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left[q_{t}-\varphi-n^{-\frac{1}{2}}\psi}}\right]\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$
 $\asymp\sqrt{\beta\log\beta}$
-for $t\in T$, as the first summand is $\Theta\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$ while the last two are $o\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$.
-Replacing $T$ by $T^{\prime}=[-\sqrt{2\log n-3\log\beta},\sqrt{2\log n-3\log\beta}]$, the first summand is $o\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$ by Lemma 2.4.
-By positivity of the integrand, we bound the last two integrals over $T^{\prime}$ by those over $T$, which are themselves $o\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$.
-Now, all three summands are $o\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$, as desired in Proposition 1.6.
-∎
+Now, by Lemma˜2.4, the first summand is $\Theta\mathopen{}\mathclose{{\left(\sqrt{\beta\log\beta}}}\right)$, while the last three are $O\mathopen{}\mathclose{{\left(e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}}}\right)$ by Lemmas˜3.3 and 3.6.
+Replacing $T$ by $T^{\prime}$, all four summands are $O(\sqrt{\beta})$, as desired.
 Figure 5: An estimate of the density $p_{t}=p_{t}(x,y)$ of $(F_{n}(t),F_{n}^{\prime}(t))$ for $t=0,1,2,3$ (clockwise from top left), where $\beta=81$ and $n=6500$, so that $\sqrt{2\log n-\log\beta}\approx 3$.
 (Code available at github.com/KimiSun18/2024-gauss-kde-attention.)
-4.2 Proof of Proposition 1.7
-In this section, we prove Proposition 1.7.
+4.2 Proof of Proposition˜1.8
+In this section, we prove Proposition˜1.8.
 Lemma 4.3.
 For any $a>0$ and $X_{1},\dots,X_{n}\in\mathbb{R}$, the number of modes of $\widehat{P}_{n}$ in $(a,\infty)$ is at most $|I|$ where $I=\{i\in[n]:X_{i}\geqslant a\}$.
-Proof of Lemma 4.3.
+By symmetry, the same estimate holds for modes in $(-\infty,-a)$.
+Proof of Lemma˜4.3.
 Note that $\widehat{P}_{n}(t)=\sum_{i=1}^{n}g_{i}(t)$ where for $i\in[n]$ we define
 $g_{i}(t)\coloneqq\sqrt{\frac{\beta}{2\pi n^{2}}}\mathsf{K}_{\beta^{-1/2}}\mathopen{}\mathclose{{\left(t-X_{i}}}\right).$ (4.2)
-For $i\not\in I$, $g_{i}$ is monotonically decreasing on $[X_{i},\infty)\supset(a,\infty)$, so $\sum_{i\not\in I}g_{i}(t)$ has no modes in $(a,\infty)$.
+For $i\not\in I$, $g_{i}$ is monotonically decreasing on $X_{i},\infty)\supset(a,\infty)$, so $\sum_{i\not\in I}g_{i}(t)$ has no modes in $(a,\infty)$.
 To this Gaussian mixture, we add in $g_{i}(t)$ for $i\in I$ one-by-one.
-By [CPW03, Theorem 2], each time the number of modes in $(a,\infty)$ increases by at most one.
+By [carreira2003number, Theorem 2], each time the number of modes in $(a,\infty)$ increases by at most one.
 In $|I|$-many steps, there are at most $|I|$ such modes.
@@
 Remark 4.4.
-This argument crucially relies on the KDE being Gaussian.
-As discussed in [CPW03], the Gaussian kernel is the only kernel where for any fixed samples the number of modes of the KDE is non-increasing in the bandwidth $h$.
-For other kernels, we do suspect the analog of Lemma 4.3 to hold, but a different argument is needed.
-In particular, [MMF92, Mam95, KM97] avoids this problem by counting modes on compact sets.
-Proof of Proposition 1.7.
-By Lemma 4.3, symmetry of $T$ in 1.2 around $t=0$, linearity of expectations, and the tail bound $\mathbb{P}(|X|\geqslant a)\leqslant 2e^{-a^{2}/2}$ for $X\sim N(0,1)$
+Two remarks are in order:
+  • As discussed in [carreira2003number], the scale-space property of the Gaussians allow us to view adding a component to the Gaussian mixture as adding a delta distribution to the mixture, which adds one mode, and applying a Gaussian blurring that does not create new modes.
+  • This argument crucially relies on the KDE being Gaussian: as discussed in [carreira2003number], the Gaussian kernel is the only kernel where for any fixed samples the number of modes of the KDE is non-increasing in the bandwidth $h$, which enables the blurring step.
+For other kernels, we do suspect the analog of Lemma˜4.3 to hold, but a different argument is needed.
+In particular, [mammen91, mammen95, mammen97] avoids this problem by counting modes on compact sets.
+Proof of Proposition˜1.8.
+By Lemma˜4.3, symmetry of $T$ in ˜1.2 around $t=0$, linearity of expectations, and the tail bound $\mathbb{P}(|X|\geqslant a)\leqslant 2e^{-a^{2}/2}$ for $X\sim N(0,1)$
 $\mathbb{E}U_{0}\mathopen{}\mathclose{{\left(F_{n},\mathbb{R}\setminus T}}\right)$ $\leqslant\mathbb{E}\mathopen{}\mathclose{{\left|\{i:X_{i}\not\in T\}}}\right|$ (4.3)
@@
 $\ll\sqrt{\beta\log\beta}$
-by the definition of $\omega(\beta)$, proving Proposition 1.7. ∎
-This concludes the proof of Theorem 1.2.
-5 Concluding remarks
-We showed that the expected number of modes of a Gaussian KDE with bandwidth $\beta^{-\frac{1}{2}}$ of $n\geqslant 1$ samples drawn iid from $N(0,1)$ is of order $\Theta(\sqrt{\beta\log\beta})$ for $n^{c}\lesssim\beta\lesssim n^{2-c}$, where $c>0$ is arbitrarily small.
-We also provide a precise picture of where the modes are located.
-The question in the higher-dimensional case, as well as on the unit sphere $\mathbb{S}^{d-1}$ with uniformly distributed samples, remains open.
-Acknowledgments
-The authors would like to thank Enno Mammen for useful discussion and sharing important references.
-We also thank Dan Mikulincer for discussions on Gaussian approximation using Edgeworth expansions, Valeria Banica for comments on the method of stationary phase, and Alexander Zimin for providing Figure 2.
-Funding. P.R. was supported by NSF grants DMS-2022448, CCF-2106377, and a gift from Apple. Y.S. was supported by the MIT UROP and MISTI France Programs.
-Appendix A Additional proofs
-A.1 Proof of 2.2
-In this section we compute the first two moments of $(G,G^{\prime})$ to prove 2.2.
+by the definition of $\omega(\beta)$, proving Proposition˜1.8. ∎
+Having proven Propositions˜1.7 and 1.8, we conclude Theorem˜1.2.
+5 Additional proofs
+5.1 Proof of Lemma˜5.1
+We frequently make use of the following standard exercise on Gaussian integrals and dominated convergence.
+Lemma 5.1.
+Let $\Gamma$ denote the Gamma function. For any $\alpha>0$ and integer $n\geqslant 0$,
+$\int_{0}^{\infty}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u}^{n}e^{-\alpha u^{2}}\,\mathrm{d}u=\frac{1}{2}\Gamma\mathopen{}\mathclose{{\left(\frac{n+1}{2}}}\right)\alpha^{-\frac{n+1}{2}}.$
+Moreover, for a fixed $n$, as $\epsilon\to 0$
+$\int_{0}^{\infty}v^{n}e^{-(v-\epsilon)^{2}}dv\to\int_{0}^{\infty}v^{n}e^{-v^{2}}dv\quad\text{and}\quad\int_{-\infty}^{\infty}|v|^{n}e^{-(v-\epsilon)^{2}}dv\to 2\int_{0}^{\infty}v^{n}e^{-v^{2}}dv$
+Proof.
+For the first display, by a change of variables $v=\alpha u^{2}$, we get
+$\int_{0}^{\infty}{u}^{n}e^{-\alpha u^{2}}\,\mathrm{d}u=\frac{1}{2}\alpha^{-\frac{n+1}{2}}\int_{0}^{\infty}v^{\frac{n-1}{2}}e^{-v}\,\mathrm{d}v$
+and we recognize the integral as the definition of the Gamma function.
+For the second display, we apply the dominated convergence theorem: clearly, the integrands converges pointwise as $\varepsilon\to 0$, and it is dominated by an integrable function that is $e2^{n}$ for $v\in[-2,2]$ and $v^{n}e^{-v^{2}/4}$ for $v\geqslant 2$ as we may take $\varepsilon\leqslant 1\leqslant|v|/2$.
+5.2 Proof of Lemma˜2.2
+In this section we compute the first two moments of $(G,G^{\prime})$ to prove Lemma˜2.2.
 Note that if $n^{c}\lesssim\beta\lesssim n^{2-c}$ for some $c>0$, and for $t\in T$, then we have $\exp\Theta(t^{2}/\beta)\to 1$.
 This implies that exponentials in the moments are asymptotically $e^{-{t^{2}}/{2}}$.
-We frequently use the following fact about Gaussian integrals both in exact and asymptotic forms.
-Fact A.1.
-Let $\Gamma$ denote the Gamma function. For any $\alpha>0$ and integer $n\geqslant 0$,
-$\int_{0}^{\infty}z^{n}e^{-\alpha u^{2}}\,\mathrm{d}u=\frac{1}{2}\Gamma\mathopen{}\mathclose{{}\left(\frac{n+1}{2}}\right)\alpha^{-\frac{n+1}{2}}.$
 We first compute $\mu_{t}$. Completing the square gives
 $\frac{\beta}{2}z^{2}+\frac{1}{2}(z-t)^{2}=\frac{\beta+1}{2}u^{2}+\frac{\beta t^{2}}{2(\beta+1)}\quad\text{where}\quad u=z-\frac{t}{\beta+1}.$
-Hence, using A.1 we compute
-$\mathbb{E}G(t)=\int_{-\infty}^{\infty}ze^{-\frac{\beta}{2}z^{2}}\,\mathrm{d}\gamma_{t,1}(z)$ $=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}ze^{-\frac{\beta}{2}z^{2}-\frac{1}{2}(z-t)^{2}}\,\mathrm{d}z$
+Hence, using Lemma˜5.1 we compute
+$\mathbb{E}G(t)=\int_{-\infty}^{\infty}ze^{-\frac{\beta}{2}z^{2}}\,\mathrm{d}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi(z-t)}$ $=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}ze^{-\frac{\beta}{2}z^{2}-\frac{1}{2}(z-t)^{2}}\,\mathrm{d}z$
 $=\frac{e^{-\frac{\beta}{2(\beta+1)}t^{2}}}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathopen{}\mathclose{{\left(u+\frac{t}{\beta+1}}}\right)e^{-\frac{\beta+1}{2}u^{2}}\,\mathrm{d}u$
@@
 as well as
-$\mathbb{E}G^{\prime}(t)$ $=\int_{-\infty}^{\infty}(1-\beta z^{2})z^{-\frac{\beta}{2}z^{2}}\,\mathrm{d}\gamma_{t,1}(z)$
+$\mathbb{E}G^{\prime}(t)$ $=\int_{-\infty}^{\infty}(1-\beta z^{2})z^{-\frac{\beta}{2}z^{2}}\,\mathrm{d}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi(z-t)}$
 $=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathopen{}\mathclose{{\left(1-\beta z^{2}}}\right)e^{-\frac{\beta}{2}z^{2}-\frac{1}{2}(z-t)^{2}}\,\mathrm{d}z$
@@
 $=\frac{e^{-\frac{\beta}{2(\beta+1)}t^{2}}}{(\beta+1)^{\frac{5}{2}}}\mathopen{}\mathclose{{\left(1+\beta-\beta t^{2}}}\right).$
-From these computations, and the remark after A.1, we readily obtain the asymptotics of $\mu_{t}$ as in 2.2 upon multiplying by $\sqrt{n}$.
+From these computations, and the remark after Lemma˜5.1, we readily obtain the asymptotics of $\mu_{t}$ as in Lemma˜2.2 upon multiplying by $\sqrt{n}$.
 We now compute $\Sigma_{t}$. Completing the square gives
-$z^{2}+\frac{1}{2}(z-t)^{2}=\frac{2\beta+1}{2}u^{2}+\frac{\beta t^{2}}{2\beta+1}\quad\text{where}\quad u=z-\frac{t}{2\beta+1}.$
-Hence using A.1 we compute
+${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}z^{2}+\frac{1}{2}(z-t)^{2}=\frac{2\beta+1}{2}u^{2}+\frac{\beta t^{2}}{2\beta+1}\quad\text{where}\quad u=z-\frac{t}{2\beta+1}.$
+Hence using Lemma˜5.1 we compute
 $\mathbb{E}G^{2}(t)$ $=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}z^{2}e^{-\beta z^{2}-\frac{1}{2}(z-t)^{2}}\,\mathrm{d}z$
@@
 $=\frac{e^{-\frac{\beta}{2\beta+1}t^{2}}}{(2\beta+1)^{\frac{9}{2}}}\mathopen{}\mathclose{{\left(12\beta^{4}+4\beta^{3}(t^{2}+5)+\beta^{2}(t^{4}-2t^{2}+15)-2\beta(t^{2}-3)+1}}\right).$
-We check that entries of $\Sigma_{t}$ are asymptotically the corresponding second moments. Indeed, surpressing the dependence on $t$, we have that
-  • $\mathbb{E}G^{2}\asymp\beta^{-\frac{5}{2}}(t^{2}+\beta)e^{-\frac{t^{2}}{2}}\gg\beta^{-3}t^{2}e^{-t^{2}}\asymp(\mathbb{E}G)^{2},$
-  • $\mathbb{E}GG^{\prime}\asymp-\beta^{-\frac{7}{2}}e^{-\frac{t^{2}}{2}}(\beta^{2}t+\beta t^{3})\gg\beta^{-4}e^{-t^{2}}t(\beta-\beta t^{2})\asymp(\mathbb{E}G)(\mathbb{E}G^{\prime}),$
-  • $\mathbb{E}G^{\prime 2}\asymp\beta^{-\frac{9}{2}}e^{-\frac{t^{2}}{2}}(\beta^{4}+\beta^{2}t^{4})\gg\beta^{-5}e^{-t^{2}}\beta^{2}(1+t^{4})\asymp(\mathbb{E}G^{\prime})^{2},$
-where we bound $e^{-\frac{t^{2}}{2}}\leqslant 1$. From these computations, and the remark after A.1, we readily obtain the asymptotics of $\Sigma_{t}$ as indicated in 2.2.∎
-A.2 Proof of 3.1
-In this section, we prove 3.1 on third moments of $Y=\Sigma_{t}^{-1/2}\mathopen{}\mathclose{{}\left(G-\mathbb{E}G,G^{\prime}-\mathbb{E}G^{\prime}}\right)$.
-To upper bound, we do not need to track the leading coefficients to ensure that they do not vanish when we combine applications of A.1.
-Recalling $\Sigma_{t}^{-1}$ from the proof of Lemma 2.3, we upper bound asymptotically via Hölder’s inequality:
-$\eta_{3}$ $=\max_{k}\mathopen{}\mathclose{{}\left|Y^{(k,3-k)}}\right|$
-$\leqslant\mathbb{E}\|Y\|^{3}$
-$\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left\|\mathopen{}\mathclose{{}\left(G-\mathbb{E}G,G^{\prime}-\mathbb{E}G^{\prime}}\right)^{\intercal}\Sigma_{t}^{-1}\mathopen{}\mathclose{{}\left(G-\mathbb{E}G,G^{\prime}-\mathbb{E}G^{\prime}}\right)}\right\|^{\frac{3}{2}}$
-$\lesssim\beta^{\frac{3}{4}}e^{\frac{3t^{2}}{4}}\mathbb{E}\mathopen{}\mathclose{{}\left|\beta(G-\mathbb{E}G)^{2}+2t(G-\mathbb{E}G)(G^{\prime}-\mathbb{E}G^{\prime})+(G^{\prime}-\mathbb{E}G^{\prime})^{2}}\right|^{\frac{3}{2}}$
-$\lesssim\beta^{\frac{3}{4}}e^{\frac{3t^{2}}{4}}\mathopen{}\mathclose{{}\left(\beta^{\frac{3}{2}}\mathbb{E}|G|^{3}+\mathbb{E}|G^{\prime}|^{3}}\right)$
-$\lesssim\beta^{\frac{3}{4}}e^{\frac{3t^{2}}{4}}\int_{-\infty}^{\infty}\mathopen{}\mathclose{{}\left(\beta^{\frac{3}{2}}|z|^{3}+|1-\beta z^{2}|^{3}}\right)e^{-\frac{3\beta}{2}z^{2}-\frac{1}{2}(z-t)^{2}}\,\mathrm{d}z$
-$\lesssim\beta^{\frac{3}{4}}e^{\frac{t^{2}}{4}}\int_{0}^{\infty}h\mathopen{}\mathclose{{}\left(u+\frac{t}{3\beta+1}}\right)e^{-\frac{3\beta+1}{2}u^{2}}\,\mathrm{d}u,$
-where we factor out $e^{-\frac{3\beta}{2(3\beta+1)}t^{2}}\asymp e^{-\frac{t^{2}}{2}}$ of the integral by substituting
-$h(z)=\beta^{\frac{3}{2}}z^{3}+(1+\beta z^{2})^{3}\quad\text{and}\quad u=z-\frac{t}{3\beta+1}.$
-By linearity of integration and A.1, each monomial $u^{n}$ integrates to $O(\beta^{-(n+1)/2})$. By monotonicity of $h$ on $\mathbb{R}_{\geqslant 0}$ and since $t\in T$, for some constant $C>0$,
-$\eta_{3}\lesssim\beta^{\frac{3}{4}}e^{\frac{t^{2}}{4}}\int_{0}^{\infty}h\mathopen{}\mathclose{{}\left(u+\frac{t}{3\beta+1}}\right)e^{-\frac{3\beta+1}{2}u^{2}}\,\mathrm{d}u$ $\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}h\mathopen{}\mathclose{{}\left(C\beta^{-\frac{1}{2}}+\frac{t}{3\beta+1}}\right)$
-$\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}h\mathopen{}\mathclose{{}\left(2C\beta^{-\frac{1}{2}}}\right)$
-$\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}$
-upon noting $w\mapsto h(w/\sqrt{\beta})$ has constant coefficients. This proves 3.1.∎
-A.3 Proof of Lemma 3.4
-Fix $n,\beta$ sufficiently large. For a multi-index $\alpha$, define
-$h(\mathbf{x})$ $=\mathbf{x}^{\alpha}\mathopen{}\mathclose{{}\left(q_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right)(\mathbf{x}),$
-$\mathscr{F}h(\mathbf{z})$ $=\partial^{\alpha}\mathopen{}\mathclose{{}\left(\mathscr{F}(q_{t})-\frac{\varphi}{6\sqrt{n}}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}}\right)(\mathbf{z}),$
-where
-$(\mathscr{F}h)(\mathbf{z})=\int_{\mathbb{R}^{2}}e^{-\mathrm{i}\langle\mathbf{x},\mathbf{z}\rangle}h(\mathbf{x})\,\mathrm{d}\mathbf{x}$
-denotes the Fourier transform of $h$. By Fourier inversion, it suffices to show that for any multi-index $\alpha$ with order $|\alpha|\leqslant 3$,
-$\mathopen{}\mathclose{{}\left|h(\mathbf{x})}\right|=\mathopen{}\mathclose{{}\left|\frac{1}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}e^{-\mathrm{i}\langle\mathbf{z},\mathbf{x}\rangle}\mathscr{F}{h}(\mathbf{z})\,\mathrm{d}\mathbf{z}}\right|\lesssim\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\mathscr{F}{h}(\mathbf{z})}\right|\,\mathrm{d}\mathbf{z}$ (A.1)
-is $O(\exp(-\omega(\beta)/2))$.
-We apply [BR10, Theorem 9.12]—which is not asymptotic and has explicit constants—so we may use it even though $q_{t}$ depends on $n$.
-Upon verifying the conditions on $Y$ via 2.2 and 3.1, we have that
-$\mathopen{}\mathclose{{}\left|\mathscr{F}{h}(\mathbf{z})}\right|\lesssim n^{-\frac{1}{2}}\eta_{3}^{\frac{1}{2}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{4}}$
-provided $\|\mathbf{z}\|\leqslant a\sqrt{n}$ for some $a\asymp\eta_{3}^{-1/2}$. By 3.1, we have that
-$\int_{\|\mathbf{z}\|\leqslant a\sqrt{n}}\mathopen{}\mathclose{{}\left|\mathscr{F}{h}(\mathbf{z})}\right|\,\mathrm{d}\mathbf{z}\lesssim n^{-\frac{1}{2}}\eta_{3}^{\frac{1}{2}}\int_{\mathbb{R}^{2}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{4}}\,\mathrm{d}\mathbf{z}\lesssim e^{-\frac{\omega(\beta)}{2}}.$ (A.2)
-Recall that $q_{t}$ is the density of $n^{-1/2}\sum_{i=1}^{n}Y_{i}$, and let $f$ denote the density of $Y$.
-Now, we proceed as the proof of [BR10, Theorem 19.2].
-As $p_{t}$ is bounded, so is $f$, and so [BR10, Theorem 19.1] gives $\mathscr{F}f\in L^{1}(\mathbb{R}^{2})$ and
-$\delta\coloneqq\sup_{\|\mathbf{z}\|>a}\mathopen{}\mathclose{{}\left|\mathscr{F}{f}(\mathbf{z})}\right|<1.$
-By properties of the Fourier transform and the product rule,
-$\int_{\|\mathbf{z}\|>a\sqrt{n}}\mathopen{}\mathclose{{}\left|\partial^{\alpha}\mathscr{F}{q_{t}}(\mathbf{z})}\right|\,\mathrm{d}\mathbf{z}$ $\lesssim\eta_{|\alpha|}n^{\frac{|\alpha|}{2}}\delta^{n-|\alpha|-1}\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\mathscr{F}{f}\mathopen{}\mathclose{{}\left(\frac{\mathbf{z}}{\sqrt{n}}}\right)}\right|\,\mathrm{d}\mathbf{z}$ (A.3)
-$\lesssim\mathopen{}\mathclose{{}\left(n\beta e^{\frac{t^{2}}{2}}}\right)^{O(1)}\delta^{n-O(1)}$
-$\lesssim e^{-\frac{\omega(\beta)}{4}}$
-for sufficiently large $n$. Finally, we bound similar to Lemma 3.2:
-$\int_{\|\mathbf{z}\|>a\sqrt{n}}$ $\mathopen{}\mathclose{{}\left|\partial^{\alpha}\frac{\varphi}{6\sqrt{n}}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}}\right|(\mathbf{z})\,\mathrm{d}\mathbf{z}$ (A.4)
-$\lesssim n^{-\frac{1}{2}}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\partial^{\alpha}H^{(k,3-k)}\varphi}\right|(\mathbf{z})\,\mathrm{d}\mathbf{z}$
-$\lesssim n^{-\frac{1}{2}}\eta_{3}\int_{\mathbb{R}^{2}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{2}}\,\mathrm{d}\mathbf{z}$
-$\lesssim e^{-\frac{\omega(\beta)}{4}}.$
-Combining A.2, A.3 and A.4 proves A.1 and hence Lemma 3.4.∎
-A.4 Proof of Proposition 4.1
-Point 1 in Theorem 2.1 can readily be seen to hold because of the explicit form of both of the fields.
+It is easy to check that entries of $\Sigma_{t}$ are asymptotically the corresponding second moments. Together, we readily obtain the asymptotics of $\Sigma_{t}$ as indicated in Lemma˜2.2.∎
+5.3 Proof of Lemma˜3.2
+In this section, we prove Lemma˜3.2 on cumulants of $Y=\Sigma_{t}^{-\frac{1}{2}}\mathopen{}\mathclose{{\left(G-\mathbb{E}G,G^{\prime}-\mathbb{E}G^{\prime}}}\right)$.
+To upper bound, we do not need to track the leading coefficients to ensure that they do not vanish when we combine applications of Lemma˜5.1.
+First, as $s$ is a constant (we only apply $s=2,3$), cumulants of order $s$ are clearly $O(\eta_{s})$.
+To bound $\eta_{s}$, we recall $\Sigma_{t}^{-1}$ from the proof of Lemma˜2.3 and apply Hölder’s inequality:
+$\eta_{s}$ $=\mathbb{E}\mathopen{}\mathclose{{\left[\|Y\|^{s}}}\right]$
+$\leqslant\mathbb{E}\mathopen{}\mathclose{{\left\|\mathopen{}\mathclose{{\left(G-\mathbb{E}G,G^{\prime}-\mathbb{E}G^{\prime}}}\right)^{\intercal}\Sigma_{t}^{-1}\mathopen{}\mathclose{{\left(G-\mathbb{E}G,G^{\prime}-\mathbb{E}G^{\prime}}}\right)}}\right\|^{\frac{s}{2}}$
+$\lesssim\beta^{\frac{s}{4}}e^{\frac{st^{2}}{4}}\mathbb{E}\mathopen{}\mathclose{{\left|\beta(G-\mathbb{E}G)^{2}+2t(G-\mathbb{E}G)(G^{\prime}-\mathbb{E}G^{\prime})+(G^{\prime}-\mathbb{E}G^{\prime})^{2}}}\right|^{\frac{s}{2}}$
+$\lesssim\beta^{\frac{s}{4}}e^{\frac{st^{2}}{4}}\mathopen{}\mathclose{{\left(\beta^{\frac{s}{2}}\mathbb{E}|G|^{s}+\mathbb{E}|G^{\prime}|^{s}}}\right)$
+$\lesssim\beta^{\frac{s}{4}}e^{\frac{(s-2)t^{2}}{4}}\int_{0}^{\infty}h_{s}\mathopen{}\mathclose{{\left(|z|}}\right)e^{-\frac{3\beta+1}{2}\mathopen{}\mathclose{{\left(z-\frac{t}{3\beta+1}}}\right)^{2}}\,\mathrm{d}z,$
+where $h_{s}(u)\coloneqq\beta^{\frac{s}{2}}u^{s}+(1+\beta u^{2})^{s}$.
+Note that the shift $t/(3\beta+1)\ll((3\beta+1)/2)^{-1/2}$, so by linearity of integration, we may apply Lemma˜5.1 to bound the integral of each monomial $|z|^{\ell}$ by $O(\beta^{-(\ell+1)/2})$.
+By monotonicity of $h$ on $\mathbb{R}_{\geqslant 0}$ and since $t\in T$, for some constant $C>0$,
+$\eta_{s}\lesssim\beta^{\frac{s-2}{4}}e^{\frac{(s-2)t^{2}}{4}}h\mathopen{}\mathclose{{\left(C\beta^{-\frac{1}{2}}}}\right)\lesssim\beta^{\frac{s-2}{4}}e^{\frac{(s-2)t^{2}}{4}}$
+upon noting $u\mapsto h(u/\sqrt{\beta})$ has constant coefficients. This proves Lemma˜3.2.
+5.4 Proof of Lemma˜3.5
+Fix $n,\beta$ sufficiently large as well as $t$. Define for a multi-index $\alpha$ that $h(\mathbf{x})=\mathbf{x}^{\alpha}g_{s}(\mathbf{x})$. Then, for $s=3$
+$\mathscr{F}h(\mathbf{z})=\partial^{\alpha}\mathopen{}\mathclose{{\left(\mathscr{F}(q_{t})-2\pi\varphi-\sum_{k=0}^{3}\frac{\varphi}{k!(3-k)!\sqrt{n}}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}}}\right)(\mathbf{z}),$
+and for $s=2$ we have the same statement with the last summand omitted. Note that we use
+$\mathscr{F}h(\mathbf{z})\coloneqq\int_{\mathbb{R}^{2}}e^{-\mathrm{i}\langle\mathbf{x},\mathbf{z}\rangle}h(\mathbf{x})\,\mathrm{d}\mathbf{x}$
+to denote the Fourier transform of $h$.
+We also omit the dependence of $h$ on $s$ and $\alpha$ for brevity.
+By Fourier inversion, it suffices to show that for any multi-index $\alpha$ with order $|\alpha|\leqslant s$ that
+$\mathopen{}\mathclose{{\left|h(\mathbf{x})}}\right|=\mathopen{}\mathclose{{\left|\frac{1}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}e^{-\mathrm{i}\langle\mathbf{z},\mathbf{x}\rangle}\mathscr{F}{h}(\mathbf{z})\,\mathrm{d}\mathbf{z}}}\right|\lesssim\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{\left|\mathscr{F}{h}(\mathbf{z})}}\right|\,\mathrm{d}\mathbf{z}\lesssim n^{-\frac{s-1}{2}}\eta_{s+1}$ (5.1)
+We apply [bhattacharya2010normal, Theorem 9.10]—which is not asymptotic and has explicit constants in $s$ only—so we may use it even though $q_{t}$ depends on $\beta$ to obtain that
+$\mathopen{}\mathclose{{\left|\mathscr{F}{h}(\mathbf{z})}}\right|\lesssim n^{-\frac{s-1}{2}}\eta_{s+1}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{4}}$ (5.2)
+provided $\|\mathbf{z}\|\leqslant a\sqrt{n}$ for some $a\asymp\eta_{s+1}^{-\frac{1}{s-1}}$. By Lemma˜3.2, we have that
+$\int_{\|\mathbf{z}\|\leqslant a\sqrt{n}}\mathopen{}\mathclose{{\left|\mathscr{F}{h}(\mathbf{z})}}\right|\,\mathrm{d}\mathbf{z}\lesssim n^{-\frac{s-1}{2}}\eta_{s+1}\int_{\mathbb{R}^{2}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{4}}\,\mathrm{d}\mathbf{z}\lesssim n^{-\frac{s-1}{2}}\eta_{s+1}.$ (5.3)
+Recall that $q_{t}$ is the density of $n^{-\frac{1}{2}}\sum_{i=1}^{n}Y_{i}$.
+Let $f$ denote the density of $W\coloneqq\sqrt{5}(Y_{1}+\dots+Y_{5})$ which exists and is bounded by Proposition˜4.1.
+Hence, $\mathscr{F}f\in L^{1}(\mathbb{R}^{2})$ and
+$\varepsilon\coloneqq\sup_{\|\mathbf{z}\|>a}\mathopen{}\mathclose{{\left|\mathscr{F}{f}(\mathbf{z})}}\right|<1.$
+Now, $q_{t}$ is the density of $\sqrt{n/5}$ times the sum of $n/5$ many i.i.d. copies of $W$, so by properties of the Fourier transform and the product rule,
+$\int_{\|\mathbf{z}\|>a\sqrt{n}}\mathopen{}\mathclose{{\left|\partial^{\alpha}\mathscr{F}{q_{t}}(\mathbf{z})}}\right|\,\mathrm{d}\mathbf{z}$ $\lesssim\eta_{|\alpha|}n^{\frac{|\alpha|}{2}}\varepsilon^{n/5-|\alpha|-1}\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{\left|\mathscr{F}{f}\mathopen{}\mathclose{{\left(\frac{\mathbf{z}}{\sqrt{n}}}}\right)}}\right|\,\mathrm{d}\mathbf{z}$ (5.4)
+$\lesssim\mathopen{}\mathclose{{\left(n\beta e^{\frac{t^{2}}{2}}}}\right)^{O(1)}\varepsilon^{n/5-s-1}$
+$\ll n^{-\frac{s-1}{2}}\eta_{s+1}$
+for sufficiently large $n$. Finally, we bound similar to Lemma˜3.3:
+$\int_{\|\mathbf{z}\|>a\sqrt{n}}\mathopen{}\mathclose{{\left|\partial^{\alpha}}}\right|(\mathbf{z})\,\mathrm{d}\mathbf{z}\lesssim\int_{{\|\mathbf{z}\|>a\sqrt{n}}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{2}}\,\mathrm{d}\mathbf{z}\ll n^{-\frac{s-1}{2}}\eta_{s+1}$ (5.5)
+and for the $s=3$ we also have the additional term
+$\int_{\|\mathbf{z}\|>a\sqrt{n}}$ $\mathopen{}\mathclose{{\left|\partial^{\alpha}\sum_{k=0}^{3}\frac{\varphi}{k!(3-k)!\sqrt{n}}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}}}\right|(\mathbf{z})\,\mathrm{d}\mathbf{z}$ (5.6)
+$\lesssim n^{-\frac{1}{2}}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}\int_{{\|\mathbf{z}\|>a\sqrt{n}}}\mathopen{}\mathclose{{\left|\partial^{\alpha}H^{(k,3-k)}\varphi}}\right|(\mathbf{z})\,\mathrm{d}\mathbf{z}$
+$\lesssim n^{-\frac{1}{2}}\eta_{3}\int_{{\|\mathbf{z}\|>a\sqrt{n}}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{2}}\,\mathrm{d}\mathbf{z}$
+$\ll n^{-\frac{s-1}{2}}\eta_{s+1}$
+Now, in the last step of both ˜5.5 and 5.6, we use that the standard Gaussian integral outside the ball at the origin converges to zero exponentially quickly as radius $a\sqrt{n}\to\infty$ by ˜3.6, so in particular
+$\int_{{\|\mathbf{z}\|>a\sqrt{n}}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{2}}\,\mathrm{d}\mathbf{z}\ll(a\sqrt{n})^{-(s-1)}\asymp n^{-\frac{s-1}{2}}\eta_{s+1}$
+Combining ˜5.3, 5.4, 5.5 and 5.6 proves ˜5.1 and hence Lemma˜3.5 for both $s=2$ and $s=3$ cases. ∎
+5.5 Proof of Proposition˜4.1
+Point 1 in Theorem˜2.1 can readily be seen to hold because of the explicit form of both of the fields.
 Point 4 also readily holds, since $F_{n}^{\prime\prime}$ is a Lipschitz function for every realization of $X_{i}$, as a sum of Lipschitz functions.
@@
 […]
+We write $\xi=(\xi_{1},\xi_{2})$ with $\|\xi\|=\rho$, and $\omega=(\cos\theta,\sin\theta)$ so that $\xi=\rho\omega$. We then have
+[…]
+with phase $\phi_{\theta}(x)=\cos\theta g(x)+\sin\theta g^{\prime}(x)$. We will show the uniform bound
+$\mathopen{}\mathclose{{\left|\mathscr{F}\nu_{t}(\xi)}}\right|\lesssim\frac{1}{\sqrt{1+\|\xi\|}},$ (5.7)
+for all $\xi\in\mathbb{R}^{2}$, where the implicit constant depends only on $\beta$. To this end, we compute
+$g^{\prime}(x)$ $=(1-\beta x^{2})e^{-\beta\frac{x^{2}}{2}},$
+$g^{\prime\prime}(x)$ $=\beta x(\beta x^{2}-3)e^{-\beta\frac{x^{2}}{2}},$
+$g^{\prime\prime\prime}(x)$ $=\beta(-\beta^{2}x^{4}+6\beta x^{2}-3)e^{-\beta\frac{x^{2}}{2}}.$
+Set $\psi(x)=(g(x),g^{\prime}(x))$. We have
+$\det(\psi^{\prime}(x),\psi^{\prime\prime}(x))=g^{\prime}(x)g^{\prime\prime\prime}(x)-(g^{\prime\prime}(x))^{2}=-\beta(\beta^{2}x^{4}+3)e^{-\beta x^{2}}<0$
+for all $x.$ In particular the determinant never vanishes.
+Observe that $\phi_{\theta}^{\prime}(x)=\cos\theta g^{\prime}(x)+\sin\theta g^{\prime\prime}(x)$, and any stationary point $x_{0}$ of $\phi_{\theta}$ satisfies $g^{\prime}(x_{0})=0$.
+If also $\phi^{\prime\prime}(x_{0})=\cos\theta g^{\prime\prime}(x_{0})+\sin\theta g^{\prime\prime\prime}(x_{0})=0$, then $(\cos\theta,\sin\theta)$ would be a nontrivial vector orthogonal to both $(g^{\prime}(x_{0}),g^{\prime\prime}(x_{0}))$ and $(g^{\prime\prime}(x_{0}),g^{\prime\prime\prime}(x_{0}))$, forcing $g^{\prime}(x_{0})g^{\prime\prime\prime}(x_{0})-(g^{\prime\prime}(x_{0}))^{2}=0$, a contradiction.
+Hence all stationary points of $\phi_{\theta}$ are non-degenerate, uniformly in $\theta$.
+Fix $R>0$ and a smooth cutoff function $\chi\in C_{c}^{\infty}(\mathbb{R})$ with $\chi\equiv 1$ on [1,1][-1,1], supported within [2,2][-2,2]. Then set $\chi_{R}(x)=\chi(x/R)$. We write
+$\mathscr{F}\nu_{t}(\xi)$ $=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}e^{-\mathrm{i}\rho\phi_{\theta}(x)}\chi_{R}(x)e^{-\frac{x^{2}}{2}}\,\mathrm{d}x+\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}e^{-\mathrm{i}\rho\phi_{\theta}(x)}(1-\chi_{R}(x))e^{-\frac{x^{2}}{2}}\,\mathrm{d}x$
+$=I_{1}+I_{2}.$
 We have
-[…]
-$=\frac{1}{(2\pi)^{\frac{n}{2}}}\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\int_{\mathbb{R}}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|^{n}\,\mathrm{d}\xi.$
-Recalling that $n\to\infty$ in our regime, we can suppose $n\geqslant 4$, and for the above integral to be finite, it suffices to show that
-$\mathopen{}\mathclose{{}\left|\int_{\mathbb{R}}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|\lesssim\frac{1}{\sqrt{\xi_{1}+\xi_{2}}}\hskip 28.45274pt\text{ as }\xi_{1},\xi_{2}\to\infty.$
-Observe that the critical points of $g$ are ±$\pm\sqrt{\frac{1}{\beta}}$, whereas those of $g^{\prime}$ are 00 and ±$\pm\sqrt{\frac{3}{\beta}}$.
-Since $\beta\to\infty$ as well, pick $\varepsilon>0$ sufficiently small and such that $\varepsilon>\sqrt{\frac{3}{\beta}}$.
-We first see that
-$\mathopen{}\mathclose{{}\left|\int_{|x|>\varepsilon}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
-$=\mathopen{}\mathclose{{}\left|\int_{|x|>\varepsilon}\frac{1}{\mathrm{i}}\frac{1}{\xi_{1}g^{\prime}(x)+\xi_{2}g^{\prime\prime}(x)}\frac{\,\mathrm{d}}{\,\mathrm{d}x}\mathopen{}\mathclose{{}\left(e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}}\right)e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
-$\lesssim\frac{1}{\xi_{1}+\xi_{2}},$
-where we used integration by parts to obtain the last inequality—this is in fact an elementary version of the method of non-stationary phase.
-For the integral over $\{|x|\leqslant\varepsilon\}$, we look to use the method of stationary phase as $\xi_{1},\xi_{2}\to\infty$, by distinguishing the three regimes $\xi_{1}\gg\xi_{2}$, $\xi_{2}\gg\xi_{1}$, and $\xi_{1}\sim\xi_{2}$.
-When $\xi_{1}\gg\xi_{2}$, we have
-$\mathopen{}\mathclose{{}\left|\int_{|x|\leqslant\varepsilon}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
-$=\mathopen{}\mathclose{{}\left|\int_{|x|\leqslant\varepsilon}e^{-\mathrm{i}(\xi_{1}+\xi_{2})\mathopen{}\mathclose{{}\left(\frac{\xi_{1}}{\xi_{1}+\xi_{2}}g(x)+\frac{\xi_{2}}{\xi_{1}+\xi_{2}}g^{\prime}(x)}\right)}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
-$=\mathopen{}\mathclose{{}\left|\int_{|x|\leqslant\varepsilon}e^{-\mathrm{i}(\xi_{1}+\xi_{2})(g(x)+O(\xi_{1}\varepsilon))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
-$\lesssim\frac{1}{\sqrt{\xi_{1}+\xi_{2}}}+o\mathopen{}\mathclose{{}\left(\frac{1}{\sqrt{\xi_{1}+\xi_{2}}}}\right)$
-by the method of stationary phase applied to the phase $g$, since $g^{\prime\prime}$ is non-degenerate at the critical points ±$\pm\sqrt{\frac{1}{\beta}}$. Similarly when $\xi_{2}\gg\xi_{1}$, we have
-$\mathopen{}\mathclose{{}\left|\int_{|x|\leqslant\varepsilon}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$ $=\mathopen{}\mathclose{{}\left|\int_{|x|\leqslant\varepsilon}e^{-\mathrm{i}(\xi_{1}+\xi_{2})(g^{\prime}(x)+O(\xi_{2}\varepsilon))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
-$\lesssim\frac{1}{\sqrt{\xi_{1}+\xi_{2}}}+o\mathopen{}\mathclose{{}\left(\frac{1}{\sqrt{\xi_{1}+\xi_{2}}}}\right)$
-by the method of stationary phase applied to the phase $g^{\prime}$, since $g^{\prime\prime\prime}$ is non-degenerate at the critical points 00 and ±$\pm\sqrt{\frac{3}{\beta}}$.
-The same argument then carries through when $\xi_{1}\sim\xi_{2}$, using the phase $g+g^{\prime}$.
-This yields the desired conclusion.
+$|I_{2}|\leqslant\int_{|x|\geqslant R}e^{-\frac{x^{2}}{2}}\leqslant\frac{1}{\sqrt{2\pi}}\frac{2}{R}e^{-\frac{R^{2}}{2}},$
+which, as $R$ is fixed, is smaller than $(1+\rho)^{-\frac{1}{2}}$ whenever $\rho$ is large enough, uniformly in $\theta$.
+Concerning $I_{1}$, because $\phi^{\prime}_{\theta}$ is analytic it only has finitely many zeros $x_{1}(\theta),\ldots,x_{m}(\theta)$ on [2R,2R][-2R,2R] (with $m\leqslant 3$).
+They are all non-degenerate: $\phi^{\prime\prime}_{\theta}(x_{j}(\theta))\neq 0$.
+Therefore there exist disjoint intervals $J_{j}(\theta)\subset[-2R,2R]$ around each $x_{j}(\theta)$ of radius $\delta>0$, and some $c_{*}=c_{*}(R,\beta,\delta)>0$ such that $|\phi^{\prime\prime}_{\theta}(x)|\geqslant c_{*}$ for all $x$ in the union of the intervals $J_{j}(\theta)$ and all $\theta\in\mathbb{S}^{1}$.
+On each $J_{j}(\theta)$, we may apply the method of stationary phase to find
+$\mathopen{}\mathclose{{\left|\int_{\cup_{j}J_{j}(\theta)}e^{-\mathrm{i}\rho\phi_{\theta}(x)}\chi_{R}(x)e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}}\right|$ $\leqslant\frac{C}{\sqrt{\rho}}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\|\chi_{R}(\cdot)e^{-\frac{(\cdot)^{2}}{2}}}}\right\|_{L^{\infty}(\mathbb{R})}+\mathopen{}\mathclose{{\left\|\mathopen{}\mathclose{{\left(\chi_{R}(\cdot)e^{-\frac{(\cdot)^{2}}{2}}}}\right)^{\prime}}}\right\|_{L^{\infty}(\mathbb{R})}}}\right)$
+$\leqslant\frac{C}{\sqrt{\rho}}\mathopen{}\mathclose{{\left(1+e^{-\frac{1}{2}}+\frac{\|\chi^{\prime}\|_{L^{\infty}(\mathbb{R})}}{R}}}\right),$
+where $C=C(R,\beta,\delta)>0$ is independent of $\rho,\theta$.
+Now set $W(\theta)\coloneqq[-2R,2R]\setminus\cup_{j}J_{j}(\theta)$.
+By compactness and continuity, we again have $c_{W}\coloneqq\inf_{\theta\in\mathbb{S}^{1}}\inf_{x\in W(\theta)}|\phi_{\theta}^{\prime}(x)|>0$.
+We now look to use the method of non-stationary phase: integration by parts gives
+$\mathopen{}\mathclose{{\left|\int_{W(\theta)}e^{-\mathrm{i}\rho\phi_{\theta}(x)}\chi_{R}(x)e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}}\right|\leqslant\frac{1}{\rho}\int_{W(\theta)}\mathopen{}\mathclose{{\left|\frac{\mathopen{}\mathclose{{\left(\chi_{R}e^{-\frac{(\cdot^{2})}{2}}}}\right)^{\prime}(x)}{\phi_{\theta}^{\prime}(x)}-\frac{\chi_{R}(x)e^{-\frac{x^{2}}{2}}\phi_{\theta}^{\prime\prime}(x)}{(\phi_{\theta}^{\prime}(x))^{2}}}}\right|\,\mathrm{d}x$
+$\leqslant\frac{1}{\rho}\mathopen{}\mathclose{{\left(\frac{1}{c_{W}}\mathopen{}\mathclose{{\left\|\mathopen{}\mathclose{{\left(\chi_{R}e^{-\frac{(\cdot)^{2}}{2}}}}\right)^{\prime}}}\right\|_{L^{1}(\mathbb{R})}+\frac{M}{c_{W}^{2}}\mathopen{}\mathclose{{\left\|\chi_{R}e^{-\frac{(\cdot)^{2}}{2}}}}\right\|_{L^{1}(\mathbb{R})}}}\right),$
+where $M=\sup_{x\in[-2R,2R]}\sup_{\theta\in\mathbb{S}^{1}}|\phi_{\theta}^{\prime\prime}(x)|<\infty$.
+All in all, both bounds yield $|\mathscr{F}\nu_{t}(\xi)|\lesssim\|\xi\|^{-\frac{1}{2}}$ for $\rho\coloneqq\|\xi\|\geqslant 1$ sufficiently large, which proves (5.7).
+We now have
+$\int_{\mathbb{R}^{2}}|\mathscr{F}\nu_{t}(\xi)|^{n}\,\mathrm{d}\xi\lesssim\int_{\|\xi\|\leqslant 1}1\,\mathrm{d}\xi+\int_{\|\xi\|>1}|\xi|^{-\frac{n}{2}}\,\mathrm{d}\xi,$
+which is finite as long as $n>4$. By Fourier inversion, we have the desired conclusion.
 To deduce that $(t,\mathbf{x})\mapsto p_{t}(\mathbf{x})$ is continuous on $T\times\mathbb{R}^{2}$, we note that
@@
 We can conclude by the Lebesgue dominated convergence theorem.
+6 Concluding remarks
+We showed that the expected number of modes of a Gaussian KDE with bandwidth $\beta^{-\frac{1}{2}}$ of $n\geqslant 1$ samples drawn iid from $N(0,1)$ is of order $\Theta(\sqrt{\beta\log\beta})$ for $n^{c}\lesssim\beta\lesssim n^{2-c}$, where $c>0$ is arbitrarily small.
+We also provide a precise picture of where the modes are located.
+The with high probability version of the statements and the question in the higher-dimensional case remains open: we conjecture the number of modes to be $\Theta(\sqrt{\beta^{d}\log\beta})$.
+We also raise the question for non-Gaussian densities as well as the case of the unit sphere $\mathbb{S}^{d-1}$ with uniformly distributed samples.
+Acknowledgments
+The authors would like to thank Enno Mammen for useful discussion and sharing important references.
+We also thank Dan Mikulincer for discussions on Gaussian approximation using Edgeworth expansions, Valeria Banica for comments on the method of stationary phase, and Alexander Zimin for providing Figure˜2.
+We finally thank all the reviewers for their comments, which have greatly improved the quality of the paper.
+Funding
+B.G.
+was supported by a Sorbonne Emergences grant, and a gift from Google.
+P.R.
+was supported by NSF grants DMS-2022448, CCF-2106377, and a gift from Apple.
+Y.S.
+was supported by the MIT UROP and MISTI France Programs.
+Borjan Geshkovski
+Inria & Laboratoire Jacques-Louis Lions
+Sorbonne Université
+4 Place Jussieu
+75005 Paris, France
+e-mail: borjan.geshkovski@inria.fr
+Philippe Rigollet
+Department of Mathematics
+Massachusetts Institute of Technology
+77 Massachusetts Ave
+Cambridge 02139 MA, United States
+e-mail: rigollet@math.mit.edu
+Yihang Sun
+Department of Mathematics
+Stanford University
+450 Jane Stanford Way Building 380
+Stanford, CA 94305, United States
+e-mail: kimisun@stanford.edu